### v3

# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

Abstract
Let NN be the maximal unipotent subgroup in the simple algebraic group of type $\Phi$. It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called coadjoint. Such orbits play the key role in the A.A. Kirillov’s orbit method. In this work, we classify the orbits of this action in the case of $\Phi=F_{4}$ in terms of supports of canonical forms. This means that we will present a set $S$ of linear forms from $\mathfrak{n}^{*}$ such that for any coadjoint orbit there exists a unique form from $S$ belonging to that orbit. The set of canonical forms will be explicitly described in terms of supports. The support of a form $\lambda\in\mathfrak{n}^{*}$ is the set of positive roots $\alpha\in\Phi^{+}$ such that $\lambda(e_{\alpha})\neq 0$.
Keywords: coadjoint orbit, the orbit method, root system.
AMS subject classification: 17B08, 17B10, 17B22, 17B25, 17B30, 17B35.
1. Introduction and the main result
Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra in $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space to $\mathfrak{n}$. The group NN is also a maximal unipotent subgroup in the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup in $G$. The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of the group NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$. According to the orbit method introduced by A.A. Kirillov in 1962, coadjoint orbits play the key role in the representation theory of the group NN. Almost all coadjoint orbits that have been studied well so far are so-called orbits associated with rook placements. They are discussed in detail in the works [An95], [AN06], [Ko12], [Ko13], [Pa08], [IPa09], [Ig09], [Ig11], [Ig12], [IS21] and [IS23].
Explicitly, the coadjoint action looks as follows:
$\begin{split}(\exp(x).\lambda)(y)=\lambda(\exp(-\mathrm{ad}_{x})(y))=\lambda(y)-\lambda([x,y])+\dfrac{1}{2!}\lambda([x,[x,y]])-\ldots,\\ x,y\in\mathfrak{n},~\lambda\in\mathfrak{n}^{*}.\end{split}$
Let $\{e_{\alpha},\alpha\in\Phi^{+}\}$ be a basis in $\mathfrak{n}$ consisting of root vectors, and $\{e_{\alpha}^{*},\alpha\in\Phi^{+}\}$ be the dual basis in $\mathfrak{n}^{*}$. The support of a linear form $\lambda\in\mathfrak{n}^{*}$ is defined as the subset of the set of positive roots of the form
$\textup{supp}(\lambda)=\{\alpha\in\Phi^{+}\colon\lambda(e_{\alpha})\neq 0\}.$
Our goal is to describe the coadjoint orbits in the case of $\Phi=F_{4}$. The description will be given in terms of supports of canonical forms. A list of supports will be provided with the following properties.
  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ with supports from the given list (possibly, having the same support), the orbits of these forms $N.\lambda_{1}$ and $N.\lambda_{2}$ do not coincide.
  • For any linear form $\lambda\in\mathfrak{n}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
In other words, there exists a unique linear form with a support from the given list on each coadjoint orbit.
The set of simple roots in $F_{4}$ can be identified with the following subset of $\mathbb{R}^{4}$:
$\Delta=\{\alpha_{1}=\varepsilon_{2}-\varepsilon_{3},~\alpha_{2}=\varepsilon_{3}-\varepsilon_{4},~\alpha_{3}=\varepsilon_{4},~\alpha_{4}=\frac{1}{2}(\varepsilon_{1}-\varepsilon_{2}-\varepsilon_{3}-\varepsilon_{4})\}.$
Here $\{\varepsilon_{i}\}_{i=1}^{4}$ is the standard basis of $\mathbb{R}^{4}$ (with the standard scalar product).
Recall that there exists the natural order on $\Phi^{+}$ defined as follows: $\alpha>\beta$ if $\alpha-\beta$ is a sum of positive roots. Let us fix a linear lexicographic order $\succ$ on the positive roots of $F_{4}$. We will say that $\alpha=a_{1}\alpha_{1}+a_{2}\alpha_{2}+a_{3}\alpha_{3}+a_{4}\alpha_{4}\succ\beta=b_{1}\alpha_{1}+b_{2}\alpha_{2}+b_{3}\alpha_{3}+b_{4}\alpha_{4}$ if $a_{1}>b_{1}$ or $a_{1}=b_{1},a_{2}>b_{2}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}>b_{3}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}=b_{3},a_{4}>b_{4}$. Thus, the roots will be ordered as follows:
$\begin{split}\Phi^{+}=\{&\alpha_{4},~\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},\\ &\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{1}+\alpha_{2},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},\\ &\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},\\ &\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},\\ &\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}\}.\end{split}$
Clearly, $\alpha>\beta$ implies $\alpha\succ\beta$. Further, if we talk about order on positive roots, we will always mean the lexicographic order as above, unless otherwise stated.
Let $\lambda\in\mathfrak{n}^{*}$, $\gamma\in\Phi^{+}$. Consider the following matrices.
$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
The columns of the matrices are numbered by all positive roots, and the rows are numbered by certain positive roots. The matrix $A_{\lambda,\gamma}$ is the matrix $B_{\lambda,\gamma}$ with one more row. Further, let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: for any $\gamma\in\textup{supp}(\lambda)$,
$\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}.$ (1)
Finally, we can formulate the main result of this work.
Theorem 1.1. For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
This theorem generalizes the result of the work [IS23]. This article discusses the so-called orbits associated with rook placements. By definition, a subset $D$ of $\Phi^{+}$ is called a non-singular rook placement if $\alpha-\beta\notin\Phi^{+}$ for all distinct $\alpha,\beta\in D$. From this it follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$. Given a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$, we put
$f_{D,\xi}=\sum_{\alpha\in D}\xi(\alpha)e_{\alpha}^{*}\in\mathfrak{n}^{*}.$
We say that the coadjoint orbit $\Omega_{D,\xi}$ of the linear form $f_{D,\xi}$ is associated with the rook placement $D$ and the map $\xi$. In essence, $f_{D,\xi}$ is just a linear form with support $D$.
The work [IS23] proves that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$. But in fact if $D$ is a non-singular rook placement then $\lambda=f_{D,\xi}\in S$. Because if $\gamma\in D$ then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$. Hence, in the matrix $A_{\lambda,\gamma}$ the line with the number $\gamma$ consists of zeros, and condition (1) is true for $\gamma\in D$. It follows from this that for distinct linear forms with support $D$ the corresponding coadjoint orbits do not coincide.
The result is similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is true for the root system $\Phi=G_{2}$ and field $\mathbb{C}$. In the same work [IS23] the equations of the orbits of basic subvarieties $\mathcal{O}_{D,\xi}$ were written out. By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
$\mathcal{O}_{D,\xi}=\sum_{\alpha\in D}\Omega_{\{\alpha\},\xi_{\alpha}},$
where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. The work also proved the Andrè’s stratification for $G_{2}$:
$\mathfrak{n}^{*}=\bigsqcup_{D,\xi}\mathcal{O}_{D,\xi},$
where the union is taken over all non-singular rook placements $D$ and all maps $\xi\colon D\to\mathbb{C}^{\times}$.
In this case $\Phi^{+}=\{\alpha,\beta,\alpha+\beta,2\alpha+\beta,3\alpha+\beta,3\alpha+2\beta\}$. Let us introduce the notation of structural constants as in the work [IS23]: $e_{\alpha},e_{\beta}]=c_{1}\cdot e_{\alpha+\beta}$, $e_{\alpha},e_{\alpha+\beta}]=c_{2}\cdot e_{2\alpha+\beta}$, $e_{\alpha},e_{2\alpha+\beta}]=c_{3}\cdot e_{3\alpha+\beta}$, $e_{3\alpha+\beta},e_{\beta}]=c_{4}\cdot e_{3\alpha+2\beta}$, $e_{\alpha+\beta},e_{2\alpha+\beta}]=c_{5}\cdot e_{3\alpha+2\beta}$.
For almost every rook placement $D$ it is true that $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$. This is only wrong for $D=\{\alpha+\beta,3\alpha+\beta\}$. In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon\{\beta,\alpha+\beta,3\alpha+\beta\}\to\mathbb{C}^{\times}$ and $\xi^{\prime}|_{D}=\xi$. Now, let us write out the equations of the orbits for $D=\{\alpha+\beta,3\alpha+\beta\}$. For a linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}\in\Omega_{D,\xi}$ we have the following system of equations.
$\begin{split}6c_{3}^{2}\lambda_{3\alpha+\beta}^{2}\lambda_{\beta}-6c_{1}c_{3}\lambda_{2\alpha+\beta}\lambda_{3\alpha+\beta}\lambda_{\alpha+\beta}+5c_{1}c_{2}\lambda_{2\alpha+\beta}^{3}=&0,\\ 2c_{3}\lambda_{\alpha+\beta}\lambda_{3\alpha+\beta}-c_{2}\lambda_{2\alpha+\beta}^{2}=&2c_{3}\xi(\alpha+\beta)\xi(3\alpha+\beta),\\ \lambda_{3\alpha+\beta}=&\xi(3\alpha+\beta),\\ \lambda_{3\alpha+2\beta}=&0.\end{split}$
For a linear form $\lambda\in\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ we have the following system of equations.
$\begin{split}6c_{3}^{2}\lambda_{3\alpha+\beta}^{2}\lambda_{\beta}-6c_{1}c_{3}\lambda_{2\alpha+\beta}\lambda_{3\alpha+\beta}\lambda_{\alpha+\beta}+5c_{1}c_{2}\lambda_{2\alpha+\beta}^{3}=&6c_{3}^{2}(\xi^{\prime}(3\alpha+\beta))^{2}\xi^{\prime}(\beta),\\ 2c_{3}\lambda_{\alpha+\beta}\lambda_{3\alpha+\beta}-c_{2}\lambda_{2\alpha+\beta}^{2}=&2c_{3}\xi^{\prime}(\alpha+\beta)\xi^{\prime}(3\alpha+\beta),\\ \lambda_{3\alpha+\beta}=&\xi^{\prime}(3\alpha+\beta),\\ \lambda_{3\alpha+2\beta}=&0.\end{split}$
In the case $\Phi=G_{2}$ the set $S$ consists of linear forms with supports from this list:
$\begin{split}\varnothing,\{\alpha\},\{\beta\},\{\alpha,\beta\},\{\alpha+\beta\},\{2\alpha+\beta\},\{2\alpha+\beta,\beta\},\{3\alpha+\beta\},\\ \{3\alpha+\beta,\alpha+\beta\},\{3\alpha+\beta,\alpha+\beta,\beta\},\{3\alpha+\beta,\beta\},\{3\alpha+2\beta\},\{3\alpha+2\beta,\alpha\}.\end{split}$
The orbits of these forms correspond to all orbits whose equations are written out in the article [IS23] and here.
Funding. The work is supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075–15–2022–287).
Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during writing the paper.
2. Description of the set of canonical forms
In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and its use. At first glance, it seems that it does not have a good description. In fact, this is not the case. The set $S$ is very well described in terms of supports. In the Appendix, a list of 880 supports will be presented.
Proposition 2.1. The set $S$ consists of all linear forms with supports from the list and forms with one of the three supports
[…]
whose coordinates satisfy the equation
$\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}.$
Here, we mean the coordinates of linear forms in the dual basis, $\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*}$.
This section is devoted to the proof of this proposition.
In the sequel, we will use the following table. It is a $24\times 24$ table, its rows and columns are numbered by positive roots. We write $\beta-\alpha$ in the cell $(\alpha,\beta)$ if $\beta-\alpha$ is the sum of positive roots, otherwise the cell is empty. The main diagonal contains zeros, it is easy to see that all cells below will be empty. Below we present the upper left corner of this table indexed by the positive roots whose decomposition does not contain $\alpha_{1}$. If at position $(\alpha,\beta)$ the difference $\beta-\alpha$ is also a positive root, this cell is highlighted in orange. Almost all examples will be considered specifically on this part of the table. Four columns are highlighted in blue because they will be encountered most often in the examples.
Table 1
$\ee\backslash e^{*}$ $\alpha_{4}$ $\alpha_{3}$ $\alpha_{3}+\alpha_{4}$ $\alpha_{2}$ $\alpha_{2}+\alpha_{3}$ $\alpha_{2}+\alpha_{3}+\alpha_{4}$ $\alpha_{2}+2\alpha_{3}$ $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
[…]
To describe the set $S$, we wrote a program, which is included in the Appendix. The program output supports that satisfy a sufficient condition for being included in the list, or supports that satisfy a necessary condition, these will be described below.
First, let us reformulate condition (1). The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: let $x,y\in\mathfrak{n},\lambda\in\mathfrak{n}^{*}$, then $(x.\lambda)(y):=-\lambda([x,y])$. This action is called tangent.
Let us expand $x\in\mathfrak{n}$ in the basis:
$x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}.$
Let $\lambda\in\mathfrak{n}^{*}$, $\beta\in\Phi^{+}$. Consider the coordinates of $x.\lambda$ in the dual basis:
$(x.\lambda)(e_{\beta})=-\lambda([\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},e_{\beta}])=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
We can consider these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$. By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$. Clearly, $\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$). In other words, $(x.\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
Example 2.2. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that such $\lambda$ lies in $S$. To visualize how the orbit of the tangent action looks like, let us refer to the table. The highlighted blue columns correspond to the roots in $D$. The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, this is equivalent to $\alpha+\beta\in D$. It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row numbered $\beta$ in the highlighted column. For example, the combination $(x.\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
Now let us check condition (1) for all roots from $D$. As you can see from the table, $(x.\lambda)(e_{\alpha_{2}+2\alpha_{3}})=0$. Naturally, this will always be the case for the maximal root in the support. Note that $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$. Further, $(x.\lambda)(e_{\alpha_{2}})=0$, because there are no orange cells in the blue columns in the row $\alpha_{2}$; $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})=0$ for similar reasons. Finally, $(x.\lambda)(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$, which means that it is a scalar multiple of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$.
It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x.\lambda)(e_{\alpha_{4}})$. It is just because $x_{\alpha_{3}}$ as a linear combination was expressed through $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient. The Sufficient condition below is based on this idea.
Consider the tangent action $\mathfrak{n}\curvearrowright\mathfrak{n}^{*}$. Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$. Starting from the support $D$ and the root $\gamma$, we will construct the set $\zeta_{D,\gamma}$ consisting of positive roots. We will add roots to $\zeta_{D,\gamma}$ one by one. We will add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$ then $\alpha\in\zeta_{D,\gamma}$. Using this procedure, we will add to $\zeta_{D,\gamma}$ the maximum possible number of roots. Note that $\zeta_{D,\gamma}$ does not depend on the values of the coordinates of the form $\lambda$ in the dual basis, but only on its support $D$. The set $\zeta_{D,\gamma}$ also does not depend on the order in which to add the roots. Indeed, let us define the sets $R_{n}$ for $n\in\mathbb{N}$. The root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$ then $\alpha\in R_{i}$ for $i<n$. It is clear that we must add all the roots of $R_{n},n\in\mathbb{N}$ to $\zeta_{D,\gamma}$, and we cannot add anything else.
Proposition 2.3. (Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$. Suppose that any root $\gamma\in D$ satisfy the following condition: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$ then $\alpha\in\zeta_{D,\gamma}$. Then $\lambda\in S$.
Proof. Fix a root $\gamma\in D$. Consider the orbit of the tangent action $\mathfrak{n}.\lambda$. Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$, while $\mu$ is given, and the values of $x_{\alpha}$ are not.
We will show by induction on $|\zeta_{D,\gamma}|$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ can be expressed as a linear combination of $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. Indeed, consider the step of adding $\delta$ to $\zeta_{D,\gamma}$. Suppose that we proved this for all roots that are already in $\zeta_{D,\gamma}$. There exists a root $\beta\succ\gamma$ such that $\delta$ is the unique root not from $\zeta_{D,\gamma}$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$. Consider $\mu(e_{\beta})=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}$. Let us rewrite:
$\lambda([e_{\delta},e_{\beta}])x_{\delta}=-\mu(e_{\beta})-\sum\limits_{\alpha\neq\delta}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
We claim that the right hand side is a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Indeed, if the coefficient at $x_{\alpha}$ in the right hand side is non-zero (i.e., $\lambda([e_{\alpha},e_{\beta}])\neq 0$), then $\alpha\in\zeta_{D,\gamma}$, hence, by the inductive assumption, $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Thus, since $\lambda([e_{\delta},e_{\beta}])\neq 0$, $x_{\delta}$ can also be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
[…]
Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ will also be satisfied for it. So, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none. Therefore, it is sufficient to check the conditions for one form with a given support, for example, for the form with coordinates 1 at $e_{\alpha}$, where $\alpha$ lies in the support.
Proposition 2.4. (Necessary condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$, $\gamma\in D$. Suppose that there exist positive roots $\delta_{1},\delta_{2},\dots\delta_{k},\epsilon_{1},\epsilon_{2},\dots\epsilon_{k}$ (not necessarily distinct) satisfying the conditions:
  • 1.
    $\delta_{1}=\gamma$ and $\delta_{i}\succ\gamma$ for all $i=\overline{2,k}$;
  • 2.
    $\lambda([e_{\delta_{i}},e_{\beta}])\neq 0$ and $\beta\notin\zeta_{D,\gamma}\iff\beta=\epsilon_{i-1}$ or $\beta=\epsilon_{i}$ for all $i=\overline{1,k-1}$;
  • 3.
    $\lambda([e_{\alpha},e_{\epsilon_{i}}])\neq 0$ and $\alpha\succeq\gamma\iff\alpha=\delta_{i}$ or $\alpha=\delta_{i+1}$ for all $i=\overline{1,k}$;
  • 4.
    $\lambda([e_{\delta_{k}},e_{\epsilon_{k}}])\neq 0$ and $\epsilon_{k}\notin\zeta_{D,\gamma}$.
Then $\lambda\notin S$.
Proof. It is enough to show that $(x.\lambda)(e_{\gamma})$ can not be expressed as a linear combination of $\{(x.\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. In other words, consider the orbit of the tangent action $\mathfrak{n}.\lambda$, and let $\mu\in\mathfrak{n}.\lambda$. It is enough to check that if all $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary. We will present $\mu^{\prime}\in\mathfrak{n}.\lambda$ for which $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is a given value. Let $\mu^{\prime}=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$ then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, as the corresponding values for $\mu$. We will find the values $x_{\epsilon_{i}}$ using induction on $i=\overline{1,k-1}$.
Base. Let $i=1$, $A=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$. Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$ then $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$. Hence, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ in the right hand side, except $x_{\epsilon_{1}}$, are already defined. Since that $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined, too. The only $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that can still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$. If $k=1$, then nothing else depends on $x_{\epsilon_{1}}$.
Step. Let the value of $x_{\epsilon_{i-1}}$ be found. Consider $\mu(e_{\delta_{i}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$. On the right side, there can only be $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$. The value of $x_{\epsilon_{i-1}}$ was determined in the previous step. That is, on the right side, only $x_{\epsilon_{i}}$ is undefined, and it is definitely there. From this condition, the value of $x_{\epsilon_{1}}$ is uniquely determined. The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that still depends on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
Case $i=k$. Consider $\mu(e_{\delta_{k}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$. The right side definitely has $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$. The value of $x_{\varepsilon_{k-1}}$ was determined in the previous step (or it is not there at all if $k=1$). All other values $x_{\alpha}$ are also defined. Therefore, the value of $x_{\epsilon_{k}}$ is uniquely determined. Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depend on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
Thus, all values $x_{\alpha}$ were calculated and the required $\mu^{\prime}$ was presented. $\square$
Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the coordinates of $\lambda$, but only on its support. Therefore, it is enough to check these conditions only for one form with a given support, for example, we may assume that all nonzero coordinates equal 11.
The conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked using programs. Below we present algorithms for checking the necessary and sufficient conditions for a support belonging to the set $S$.
Algorithm
An incomplete enumeration of supports is performed. For each support, the necessary (respectively, sufficient) condition is checked, and then the next support to be checked is built based on it. The enumeration is based on the following fact. Let $D$ be a support, $\alpha$ a positive root that is not in $D$ and is smaller than all roots from $D$. Then if $D$ does not satisfy the necessary (respectively, sufficient) condition, then $D\cup\{\alpha\}$ also does not satisfy the necessary (respectively, sufficient) condition. Indeed, the necessary (respectively, sufficient) condition for the support does not hold when there exists a root $\gamma$ in the support that satisfies certain conditions. These conditions depend only on $\beta\in D$ such that $\beta\succeq\gamma$. Neither the set $\zeta_{D,\gamma}$, nor the values $\lambda([e_{\alpha},e_{\beta}])$, where $\alpha$ or $\beta\succeq\gamma$, depend on the roots in the support that are smaller than $\gamma$.
The enumeration is performed according to the following rules. First, the support $\{2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}\}$ is checked for compliance with the necessary (respectively, sufficient) condition. Further, let us assume that we have checked the support $D$ for compliance with the necessary (respectively, sufficient) condition. Let $\delta$ be the minimal root in $D$. First, let us consider the case when $\delta\neq\alpha_{4}$. If $D$ passed the check, we add to $D$ the positive root closest to $\delta$ from below. If $D$ did not pass the check, we remove $\delta$ from $D$ and add the positive root closest to $\delta$ from below. Then we check the new support for compliance with the necessary (respectively, sufficient) condition. Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if exists) and add the root closest to $\beta$ from below. If $D=\{\alpha_{4}\}$, then the enumeration ends here. In the end, we include the empty set in the list of supports, because it definitely satisfies the necessary (respectively, sufficient) condition. Thus, we will enumerate all supports that potentially satisfy the necessary (respectively, sufficient) condition, and we will not check supports that definitely do not satisfy these conditions.
Let us clarify that if $D$ is the support of a form $\lambda$, then the condition $\lambda([e_{\alpha},e_{\beta}])\neq 0$ is equivalent to $\alpha+\beta\in D$. Checking this condition can be easily implemented. As it was already mentioned, the necessary (respectively, sufficient) condition needs to be checked for all roots of the support $D$. It is obvious that it is enough to check it only for the root $\gamma$ that we just added to the support, because for the rest of the roots we checked the condition before. The first thing to do is to build the set $\zeta_{D,\gamma}$. It can be built as follows. We launch a cycle and go through the rows that are numbered by roots $\alpha\succ\gamma$. If there is only one non-zero element $\lambda([e_{\alpha},e_{\delta}])$ in the row $\alpha$ such that $\delta\notin\zeta_{D,\gamma}$, then we add $\delta$ to $\zeta_{D,\gamma}$. If, after performing the cycle, the number of roots in $\zeta_{D,\gamma}$ has increased, we go through it again. The process stops when the cardinality of $\zeta_{D,\gamma}$ stabilizes. Thus, we have built $\zeta_{D,\gamma}$.
Checking the sufficient condition for the support $D$ is trivial. We go through the row numbered $\gamma$ and check the condition of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The necessary condition is checked as follows. Let us describe the $i$-th step of the check. We go through the row numbered $\delta_{i}$ (recall that $\delta_{1}=\gamma$). If we meet a non-zero element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$ and $\varepsilon\neq\varepsilon_{i-1}$, we check the column numbered $\varepsilon$, starting from $\gamma$ and higher. If there are no non-zero elements in this part of the column, except for $\lambda([e_{\delta_{i}},e_{\varepsilon}])$, then the necessary condition is not satisfied. If this is not the case, but only one non-zero element was found in the row (possibly except for $\lambda([e_{\delta_{i}},e_{\varepsilon_{i-1}}])$) of the form $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$, and at the same time, only one non-zero element was found in the corresponding part of the column, except for ours, of the form $\lambda([e_{\delta},e_{\varepsilon}])$, then $\varepsilon=\varepsilon_{i}$, and $\delta=\delta_{i+1}$. And then we check the condition already for the column numbered $\delta_{i+1}$. In fact, our program allows checking the necessary condition only up to the second step. And if the necessary condition is still not violated by that moment, it considers that the check has passed. But this hardly complicates the final calculation.
The results of the check are as follows: 878 supports satisfy the sufficient condition, while 911 satisfy the necessary condition, so 33 supports remain to be checked manually. Here they are:
  1. 1.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  2. 2.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{1}$
  3. 3.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}$
  4. 4.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  5. 5.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  6. 6.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  7. 7.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  8. 8.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  9. 9.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  10. 10.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  11. 11.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  12. 12.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3},~\alpha_{4}$
  13. 13.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  14. 14.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3},~\alpha_{4}$
  15. 15.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{4}$
  16. 16.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  17. 17.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  18. 18.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  19. 19.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  20. 20.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  21. 21.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  22. 22.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  23. 23.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  24. 24.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  25. 25.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  26. 26.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  27. 27.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  28. 28.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  29. 29.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{4}$
  30. 30.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  31. 31.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  32. 32.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  33. 33.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
Linear forms with support 5 actually don’t satisfy the necessary condition, but checking this requires three steps, while the algorithm checks only two steps, see above. It is clear that linear forms with the supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1), i.e., such forms do not belong to $S$. Indeed, it follows immediately from Table 1 as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. In other cases, it is necessary to explicitly write out the coordinates of the tangent action. It turns out that linear forms with the supports numbered 25 and 33 lie in $S$ without restrictions on the coordinates. Linear forms with supports numbered 30, 31 and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$. We obtain 883 supports whose forms belong to $S$, 3 of which require a restriction on the coordinates of forms. Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved.
Example 2.5. Consider a form $\lambda$ with support 33. We will show that $\lambda\in S$. Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}$. Let us explicitly write down the coordinates we are interested in. We will denote the structural constants in the expression $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. Then
$\begin{split}\mu_{\alpha_{1}+\alpha_{2}}&=\mu_{\alpha_{2}+2\alpha_{3}}=\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=0;\\ \mu_{\alpha_{1}}&=-\lambda([x,e_{\alpha_{1}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{1}}])=-C_{\alpha_{2},\alpha_{1}}x_{\alpha_{2}}\lambda(e_{\alpha_{1}+\alpha_{2}})=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ &=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\lambda_{\alpha_{1}+\alpha_{2}}}\mu_{\alpha_{1}};\\ \mu_{\alpha_{3}}&=-\lambda([x,e_{\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{3}+\alpha_{4}}-2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{4}}&=-\lambda([x,e_{\alpha_{4}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])\\ &=-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\\ &=\dfrac{\lambda_{\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{2}+\alpha_{3}}-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{3}}.\end{split}$
Let us also consider the support numbered 25. All roots, except $\alpha_{2}$, obviously satisfy condition (1). The coordinate $\mu_{\alpha_{2}}$ is equal to $x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}}$, so if the polynomial $x_{\alpha_{3}}$ can be expressed as a linear combination of $\{\mu_{\beta}\}_{\beta\succ\alpha_{2}}$ then $\alpha_{2}$ satisfies condition (1). That is exactly the case:
$\begin{split}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+2\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}};\\ x_{\alpha_{3}}&=\dfrac{\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}-\mu_{\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}+\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{3\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
Finally, let us consider forms with the supports 30, 31 and 32. Let the support contain roots $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$. Let us check whether condition (1) holds for the root $\alpha_{2}$. Let us note that
$\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+x_{\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{2}}&=-x_{\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}}.\end{split}$
We see that if $\beta\succ\alpha_{2}$ and $x_{\alpha_{3}+\alpha_{4}}$ enters $\mu_{\beta}$ then $\beta=\alpha_{2}+\alpha_{3}$; similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ enters $\mu_{\beta}$ then $\beta=\alpha_{2}+2\alpha_{3}$. Hence, the coordinate $\mu_{\alpha_{2}}$ should be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$. It is easy to check that this is true if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
Let us check whether condition (1) holds for the root $\alpha_{4}$. Note that
$\begin{split}\mu_{\alpha_{1}}&=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}}-x_{\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}.\end{split}$
Arguing as above we obtain that the coordinate $\mu_{\alpha_{4}}$ should be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$. It turns out that this is true if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
3. Classification of orbits
In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, which will complete a classification of coadjoint orbits for type $F_{4}$. For the next stage, we will need a concept of a singular root.
Definition 3.1. Let $\alpha,\gamma\in\Phi^{+}$. The root $\alpha$ is called singular to $\gamma$ if $\gamma-\alpha\in\Phi^{+}$. The set of all roots singular to $\gamma$ is denoted by $S_{\gamma}$.
It is obvious that the singular roots from $S_{\gamma}$ are divided into pairs of the form $\alpha+\beta=\gamma$.
Next, unless stated otherwise, everything will take place over the closure of the field $\overline{\mathbb{F}}$. We introduce two $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$-modules:
$\begin{split}M&=\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]\otimes_{\mathbb{F}}\mathfrak{n},\\ M^{*}&=\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]\otimes_{\mathbb{F}}\mathfrak{n}^{*}.\end{split}$
Here $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ denotes the ring of polynomials over $\overline{\mathbb{F}}$ in independent variables in $x_{\alpha},\alpha\in\Phi^{+}$. Next, we define the pairing $M^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ by putting
$\langle f\otimes\lambda,g\otimes x\rangle\mapsto\lambda(x)fg.$
Everywhere below we will write $\lambda(x)$ instead of $\langle\lambda,x\rangle$ for $\lambda\in M^{*},x\in M$. Notice that $\mathfrak{n}$ is a $\mathbb{F}$-subspace in $M$, and $\mathfrak{n}^{*}$ is a $\mathbb{F}$-subspace in $M^{*}$. Therefore, two other pairings $\mathfrak{n}^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $M^{*}\times\mathfrak{n}\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ are automatically defined. We introduce a Lie algebra structure on $M$ in an obvious way: for $f,g\in\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $x,y\in\mathfrak{n}$, then
$f\otimes x,g\otimes y]=fg\otimes[x,y].$
Since $\mathfrak{n}$ can be identified with upper triangular matrices with coefficients from $\mathbb{F}$, we can identify $M$ with upper triangular matrices with coefficients from $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$. Therefore, we can define $\exp(x)$ for $x\in M$. Since the characteristic of $\mathbb{F}$ is sufficiently large, $\exp(x)$ is well defined for all $x\in M$. Let $N_{M}$ be the group consisting of the matrices $\{\exp(x),x\in M\}$. The actions $N_{M}\curvearrowright M$ and $N_{M}\curvearrowright M^{*}$ are naturally defined. The explicit description of the action $N_{M}\curvearrowright M^{*}$ is as follows: if $x\in M,\lambda\in M^{*}$, then
$\exp(x).\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda\left(\sum\limits_{i=0}^{\infty}\dfrac{(-1)^{i}}{i!}\underbrace{[x,[x,\dots,[x,e_{\alpha}]\dots]]}_{\text{$i$ commutators}}\right)\otimes e_{\alpha}^{*}.$
Denote the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$. For forms $\lambda\in\mathfrak{n}^{*}$, we will consider the elements $\exp(\chi).\lambda\in M^{*}$. Let us designate $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$. Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ for certain $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
Now we formulate one of the most important theorems of this work. Let $y_{\beta},\beta\in\Phi^{+}$, be independent variables.
Theorem 3.2. Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, $\gamma\in\Phi^{+}$ satisfies condition (1). Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
$\mu(e_{\gamma})=\lambda(e_{\gamma})+F_{D,\gamma}(\mu(e_{\beta}),\beta\succ\gamma),$ (2)
where $\mu=\exp(\chi).\lambda\in M^{*}$.
Example 3.3. Consider the same support as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that for $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds. Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}$. Let us explicitly write down the coordinates we are interested in. As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we will denote the structural constants in the expression $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. We have
$\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=0;\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}}&=\lambda(e_{\alpha_{2}})+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{2}}]])=\lambda_{\alpha_{2}}+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}}]])\\ &=\lambda_{\alpha_{2}}+\dfrac{1}{2}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}C_{\alpha_{3},\alpha_{2}}x_{\alpha_{3}}^{2}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}^{2}}{4\mu_{\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{3}+\alpha_{4}}])\\ &=\lambda_{\alpha_{4}}-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
In fact, the validity of formula (2) in this example could have been understood without explicitly writing out the coordinates. It is enough to look at the Table 1. Due to the fact that $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, $x_{\alpha_{3}}$ can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Further, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$. Since $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
Now let us move on to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. First, we will prove it in a particular case, for forms with a support consisting of one root. Obviously, any such a form lies in $S$.
Lemma 3.4. Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for any $\lambda\in\mathfrak{n}^{*}$ with $|\textup{supp}(\lambda)|=1$.
Proof. Denote $D=\textup{supp}(\lambda)=\{\delta\}$. Let us note that in this case, the equality (1) only depends on $D$ and not on $\lambda$, and holds for those $\gamma$ that are not singular to $\delta$. Let $\mu=\exp(\chi).\lambda$. We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}. Note that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$, so for these cases, the condition is obviously satisfied. It remains to prove it for $\gamma\prec\delta$.
For any $\beta\in S_{\delta}$, the following equality holds:
$\begin{split}\mu(e_{\beta})=&-\lambda([x,e_{\beta}])+\dfrac{1}{2!}\lambda([x,[x,e_{\beta}]])-\dfrac{1}{3!}\lambda([x,[x,[x,e_{\beta}]]])+\ldots\\ =&-\lambda([x_{\delta-\beta}e_{\delta-\beta},e_{\beta}])+F(x)=-\lambda([e_{\delta-\beta},e_{\beta}])x_{\delta-\beta}+F(x)=C\lambda(e_{\delta})x_{\delta-\beta}+F(x),\end{split}$
where $C\neq 0$ and $F(x)$ is a polynomial in $x_{\alpha}$, where $\alpha<\delta-\beta$. Thus,
$x_{\delta-\beta}=\dfrac{\mu(e_{\beta})-F(x)}{C\lambda(e_{\delta})}$ (3)
Let $\beta_{1},\beta_{2},\ldots,\beta_{2k}$ be the singular roots to $\delta$, listed in decreasing order with respect to lexicographical order. They are paired such that $\beta_{i}+\beta_{2k-i+1}=\delta$. Let $\gamma\notin S_{\delta}$ with $\delta\succ\gamma$. Consider $\mu(e_{\gamma})$. Suppose that there is no rational function $H$ over $\overline{\mathbb{F}}$ of independent variables $t_{\beta},\beta\succ\gamma$ such that $\mu(e_{\gamma})=H(\mu(e_{\beta}),\beta\succ\gamma)$.
The coordinate $\mu(e_{\gamma})$ is a polynomial in $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$. It does not contain $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$. Let $\beta_{2k-j+1}$ be the largest root from $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ is in $\mu(e_{\gamma})$. We express $x_{\beta_{2k-j+1}}$ using formula (3) through $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}. Thus, now $\mu(e_{\gamma})$ is a polynomial that does not include $x_{\beta_{2k-j+1}}$. We perform a similar operation for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$. We note that at each step, $x_{\beta_{i}}$, which was initially absent and has already been removed, cannot reappear, as we are proceeding in decreasing order of singular roots. As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}(t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$. Note that all $\beta_{i}\succ\gamma,i=1,\ldots,j$.
According to our assumption, $\mu(e_{\gamma})$ cannot be expressed in terms of the subsequent coordinates, which means that this polynomial is not identically equal to some number and contains at least one variable $x_{\alpha}$. Then, there exists an open Zariski subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is not a constant. Fix any such set $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$. Choose $\varepsilon\notin S_{\delta}$ and $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$ such that the polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is not constant. The field is algebraically closed, therefore for any $c_{0}\in\overline{\mathbb{F}}$, there exists such $c_{\varepsilon}\in\overline{\mathbb{F}}$ that $g(c_{\varepsilon})=c_{0}$. We fix $c_{0}$ and $c_{\varepsilon}$. We can choose any $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$. By substituting $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into the formula (3), we will sequentially assign values to $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ so that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$. For these values, $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
Thus, there exists an open Zariski subset $V\subseteq\overline{\mathbb{F}}^{2k}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ there exist $\overline{\mu}\in\overline{N}.\lambda$ for which the coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values. The orbit $\overline{N}.\lambda$ is an algebraic variety. From our construction, it follows that there is a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, therefore $\dim\overline{N}.\lambda>2k$. But $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$. A variety of dimension $2k$ cannot contain $\overline{\mathbb{F}}^{2k+1}$. This is a contradiction. $\square$
We introduce more notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
$T_{\alpha,\beta}(x):=(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\in\overline{\mathbb{F}}[x_{\gamma},\gamma\in\Phi^{+}]$. It is clear that if $\lambda\in\mathfrak{n}^{*},\alpha\in\Phi^{+}$, then
$\begin{split}(\exp(\chi).\lambda)(e_{\alpha})=\left(\exp(\chi).\left(\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})e_{\beta}^{*}\right)\right)(e_{\alpha})&=\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\\ &=\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})T_{\alpha,\beta}(x).\end{split}$
Now consider the following construction (cf. the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$). Let $\lambda\in S,~D=\textup{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1). Starting from $D$ and $\gamma$, we will construct the set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
Example 3.5. Table 1 helps to visualize how the polynomials $T_{\alpha,\beta}(x)$ look like. It is enough to look at the cell in the row numbered $\alpha$ and the column numbered $\beta$. For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, we denote the non-zero constant $C_{\gamma,\delta}$ by letting $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$. Each monomial in $T_{\alpha,\beta}(x)$ has the form
$(-1)^{n+1}\dfrac{1}{n!}C_{\gamma_{1},\alpha}C_{\gamma_{2},\alpha+\gamma_{1}}\ldots C_{\gamma_{n},\alpha+\gamma_{1}+\ldots+\gamma_{n-1}}x_{\gamma_{1}}x_{\gamma_{2}}\ldots x_{\gamma_{n}},$
where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 1 to $n$, $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$. Hence, it is enough to subsequently consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots. For instance, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
Note that if $\alpha\nleq\beta$ then $T_{\alpha,\beta}(x)=0$, while $T_{\alpha,\alpha}(x)=1$. In all other cases, $T_{\alpha,\beta}(x)$ has non-zero degree. Let us call a pair of roots $(\alpha,\beta)$ nontrivial if $\alpha<\beta$ (i.e., $T_{\alpha,\beta}(x)\neq const$). Obviously, each trivial pair belongs to all $\zeta^{\prime}_{D,\gamma}$.
Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ again. There $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ do not contain nontrivial pairs. Indeed, there are no $\beta$ in $D$ such that $\beta>\alpha_{2}+2\alpha_{3}$ or $\beta>\alpha_{2}+\alpha_{3}+\alpha_{4}$. Since
$(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})=2x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}(\exp(\chi).\lambda)(e_{\alpha_{2}+2\alpha_{3}})$
and $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)=-x_{\alpha_{3}}^{2}$, the pair $(\alpha_{2},\alpha_{2}+2\alpha_{3})$ enters $\zeta^{\prime}_{D,\alpha_{2}}$. This pair also belongs to $\zeta^{\prime}_{D,\alpha_{3}+\alpha_{4}}$ and $\zeta^{\prime}_{D,\alpha_{4}}$, but $\zeta^{\prime}_{D,\alpha_{4}}$ contains one more nontrivial pair, namely, $(\alpha_{4},\alpha_{3}+\alpha_{4})$, because $T_{\alpha_{4},\alpha_{3}+\alpha_{4}}(x)=x_{\alpha_{3}}$. The remaining pairs of the form $(\gamma,\alpha)$, where $\alpha\in D$ and $\gamma$ satisfies condition (1), are trivial, therefore, they lie in all $\zeta^{\prime}_{D,\gamma}$. The lemma below highlights why we consider exactly these pairs.
Lemma 3.6. Let $\lambda\in S,D=\textup{supp}(\lambda)$, $\gamma\in\Phi^{+}$ satisfy condition (1). Let for any $\alpha\in D$ the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$ and formula (2) is true for all $\gamma^{\prime}\succ\gamma$ satisfying condition (1). Then formula (2) is also true for $\lambda$, $D$ and $\gamma$.
Proof. Consider
$(\exp(\chi).\lambda)(e_{\gamma})=\sum\limits_{\alpha\in D}\lambda(e_{\alpha})T_{\gamma,\alpha}(x)=\lambda(e_{\gamma})+\sum\limits_{\alpha\in D,\alpha\succ\gamma}\lambda(e_{\alpha})T_{\gamma,\alpha}(x).$
Let $\alpha\in D$, $\alpha\succ\gamma$. We have $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, hence $T_{\gamma,\alpha}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. It follows from the construction of $S$ that $\alpha$ satisfies condition (1). Hence, by our assumption, $\lambda(e_{\alpha})$ can be also expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
$\lambda(e_{\alpha})=(\exp(\chi).\lambda)(e_{\alpha})-F_{D,\alpha}((\exp(\chi).\lambda)(e_{\beta}),\beta\succ\alpha).$
Finally, $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, so condition (2) is satisfied, as required. $\square$
Finally, we are now ready for the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
Proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. If $D=\{\alpha\}$ then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. In fact, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$. Furthermore, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$. Finally, for $\gamma\prec\alpha$ satisfying condition (1), we see that $T_{\gamma,\alpha}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
Now, we will prove formula (2) step by step for all $\lambda\in S$ by subsequent consideration of supports from the list. (Note that most of the supports can be considered similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
  • Let the largest (with respect to the lexicographic order) root in the support be less than $\alpha_{2}+\alpha_{3}+\alpha_{4}$. In these case, Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for all supports and all $\gamma\in\Phi^{+}$ satisfying condition (1), so the prove is complete. (In fact, $T_{\gamma,\alpha}(x)$ for all $\gamma\in\Phi^{+}$ satisfying condition (1) are equal up to multiplication by a certain nonzero scalar to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0 or 1.)
  • Let the largest root in the support be equal to $\alpha_{2}+\alpha_{3}+\alpha_{4}$. There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$. Due to the fact that the condition of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is satisfied for supports consisting of one root, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ or not (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$). The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, hence there is noting to prove in this case.
  • Next, let the largest root in the support be equal to $\alpha_{2}+2\alpha_{3}$. Here, the first non-trivial case arises, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ can not be proved for all such supports just by subsistent applying of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$. This happened due to the fact that in the table cells at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ the same root $\alpha_{2}+\alpha_{3}$ is written. However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$. Hence, the following situation can occur: $\alpha_{4}$ satisfies condition (1), but we can not apply Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Therefore, here we have to compute the coordinates of $\mu$ explicitly:
    $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}=&~\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=&~\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}}=&-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-\\ &-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{3}+\alpha_{4}}=&-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&-x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{3}}=&-\lambda([x,e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{3}}]])=-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{2}}e_{\alpha_{2}},[x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{4}}e_{\alpha_{4}},e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x_{\alpha_{4}}e_{\alpha_{4}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}}])\\ =&-C_{\alpha_{2}+\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\dfrac{1}{2}C_{\alpha_{2},\alpha_{3}}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\\ &+\dfrac{1}{2}(C_{\alpha_{4},\alpha_{3}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}+C_{\alpha_{2},\alpha_{3}}C_{\alpha_{4},\alpha_{2}+\alpha_{3}})x_{\alpha_{2}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&~2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}+x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}-x_{\alpha_{2}}x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{4}}=&~\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{4}}])+\frac{1}{2}\lambda([x,[x,e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})+\\ &+\dfrac{1}{2}C_{\alpha_{3},\alpha_{4}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\\ &+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\mu_{\alpha_{3}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
    This shows that formula (2) holds for root $\alpha_{4}$.
  • Now, let the largest root in the support be equal either to $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or to $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$. In these cases, everything is also almost trivial due to Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The only non-trivial case is when the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$. At first glance, it seems that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$. In fact, this is not the case. From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ is expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$. Therefore, $x_{\alpha_{3}}$ can be found from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$, thus $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ contains only $x_{\alpha_{3}}$.
  • If the largest root in the support is equal to $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the non-trivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$. This case follows immediately from the formulas in the Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
  • If the largest root in the support is equal to $\alpha_{1}+\alpha_{2}+\alpha_{3}$, then the all non-trivial cases can be reduced to the case $D=\{\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4},\alpha_{2}\}$. Let us write down the coordinates we are interested in:
    $\begin{split}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}&=\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}},\\ \mu_{\alpha_{1}+\alpha_{2}}&=-x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}},\\ \mu_{\alpha_{1}}&=-x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}},\\ \mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}&=\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+2\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=-\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}},\\ \mu_{\alpha_{2}+\alpha_{3}}&=x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}}+x_{\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}-\dfrac{3}{2}x_{\alpha_{3}}x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}}&=\lambda_{\alpha_{2}}-x_{\alpha_{1}}x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}}-x_{\alpha_{3}+\alpha_{4}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{3}}^{2}x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}\\ &=\lambda_{\alpha_{2}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}}+\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}}^{2}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}^{2}},\\ \mu_{\alpha_{3}+\alpha_{4}}&=-x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\dfrac{\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}}.\end{split}$
    We conclude that for roots $\alpha_{2}$ and $\alpha_{3}+\alpha_{4}$ formula (2) holds.
  • Next, if the largest root in the support is equal to $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those with roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$. One can easily show that
    $\mu_{\alpha_{2}}=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{1}+\alpha_{2}}^{2}\mu_{\alpha_{2}+2\alpha_{3}+2\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}.$
    First, consider the case when $\lambda_{\alpha_{2}+2\alpha_{3}}=0$. Clearly, the root $\alpha_{2}$ satisfies formula (2). Since $\alpha_{2}+2\alpha_{3}\notin D$, $D$ may contain either $\alpha_{3}$ or $\alpha_{2}+\alpha_{3}$. But
    $\begin{split}\mu_{\alpha_{3}}&=\lambda_{\alpha_{3}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}+\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}}+2\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+2\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}},\\ \mu_{\alpha_{2}+\alpha_{3}}&=\lambda_{\alpha_{2}+\alpha_{3}}+\dfrac{-\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}\mu_{\alpha_{1}+\alpha_{2}}+\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}}-\\ &-2\dfrac{\mu_{\alpha_{2}+2\alpha_{3}+2\alpha_{4}}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}.\end{split}$
    One can see that in both cases formula (2) is satisfied. Second, if $\lambda_{\alpha_{2}+2\alpha_{3}}\neq 0$ formula (2) for $\alpha_{2}$ follows from the fact that $\lambda_{\alpha_{2}+2\alpha_{3}}$ can be found from $\mu_{\alpha_{2}+2\alpha_{3}}$, while $x_{\alpha_{3}}$ can be found from $\mu_{\alpha_{2}+\alpha_{3}}$.
  • The largest root in the support is equal to $\alpha_{1}+\alpha_{2}+2\alpha_{3}$. Here everything essentially boils down to the supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$, $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.(All other cases can be reduced to this two supports arguing similarly to the above.)In the first case, we have the following equations on coordinates of $\mu$:
    $\begin{split}\mu_{\alpha_{2}+\alpha_{3}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+2\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{2}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+2\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{4}}&=\dfrac{-\mu_{\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
    For the second support, we have the following equations on coordinates of $\mu$:
    $\begin{split}\mu_{\alpha_{2}}&=-\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+\alpha_{3}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}^{2}\mu_{\alpha_{2}+2\alpha_{3}}}{4\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}^{2}};\\ \mu_{\alpha_{4}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-2\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
  • The largest root in the support is equal to $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$. Despite the fact that there are a lot of such supports, they are all amenable to simple analysis. Each such a support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{1}+\alpha_{2}+\alpha_{3}$. Therefore, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$ and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as a rational function in $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as a polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$). Besides, $x_{\alpha_{1}}$ can be expressed as a rational function in $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$. It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ enters a monomial in a coordinate of $\mu_{\alpha}$ then this monomial can be expressed as a rational function in coordinates $\mu_{\beta}$ with $\beta\succ\alpha$. Moreover, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$. Finally, those monomials that can not expressed in such a way satisfy formula (2) due to the analysis above. A similar analysis can also be done with supports where the largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
Further steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ works or the equations for the coordinates can be easily explicitly computed. Thus, Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved. $\square$
Finally, let us formulate and prove the last theorem, which allows us to obtain the main result.
Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
Proof. Assume the contrary. Let $\lambda\neq\lambda^{\prime}$ and at the same time $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$. Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$. Then $\gamma$ must be in the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$. Since both of these forms belong to $S$, the root $\gamma$ satisfies condition (1). Let $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$. It follows from Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that there exist a rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
$\mu(e_{\gamma})=\lambda(e_{\gamma})+F_{D,\gamma}(\mu(e_{\beta}),\beta\succ\gamma)$
and
$\mu^{\prime}(e_{\gamma})=\lambda^{\prime}(e_{\gamma})+F_{D^{\prime},\gamma}(\mu^{\prime}(e_{\beta}),\beta\succ\gamma).$
Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$. Hence, we may choose our functions $F_{D,\gamma}$ for various $D$ and $\gamma$ so that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$. Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary form from their common orbit, $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$. Then
$\widetilde{\mu}(e_{\gamma})=\lambda(e_{\gamma})+F_{D,\gamma}(\widetilde{\mu}(e_{\beta}),\beta\succ\gamma)$
and
$\widetilde{\mu}(e_{\gamma})=\lambda^{\prime}(e_{\gamma})+F_{D,\gamma}(\widetilde{\mu}(e_{\beta}),\beta\succ\gamma).$
[…]
Thus, distinct forms in $S$ define have distinct orbits. Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$ then $N.\lambda\neq N.\lambda^{\prime}$.
It remains to prove that all coadjoint orbits are obtained in this way. Consider the orbits over the field $\mathbb{F}_{q}$. It is easy to calculate $|S|$. Let $D$ be a support from the list, $|D|=k$. There are $(q-1)^{k}$ linear forms with this support. If they all lie in $S$, then this support defines $(q-1)^{k}$ forms from $S$. If there is one restriction on the scalars, then the support defines $(q-1)^{k-1}$ forms from $S$, since the restriction means that one coordinate is expressed through others. For example,
$\lambda_{\alpha_{1}+\alpha_{2}}=\dfrac{\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}{\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{2}}$
for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, one can easily calculate that
$\begin{split}|S|&=1+24(q-1)+140(q-1)^{2}+288(q-1)^{3}+256(q-1)^{4}\\ &+124(q-1)^{5}+40(q-1)^{6}+9(q-1)^{7}+(q-1)^{8}.\end{split}$
The sum of the coefficients is exactly 883, the number of supports in the list plus three supports with restrictions.
Since each form from $S$ defines its own orbit, the set $S$ defines $|S|$ orbits. But in work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is calculated using group-theoretic methods. And it is exactly equal to $|S|$. Hence, over the field $\mathbb{F}_{q}$ there are no other orbits. The main result now is completely proved.
Similarly, other root systems $\Phi$ can be considered. Let us fix a linear lexicographic order $\succ$ on the positive roots of $\Phi^{+}$. Let $\lambda\in\mathfrak{n}^{*}$, $\gamma\in\Phi^{+}$. Consider the following matrices.
$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: any $\gamma\in\textup{supp}(\lambda)$ satisfies condition (1). We can formulate a hypothesis.
Conjecture 3.8. For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
Matvey A. Surkov: Samara National Research University, Ak. Pavlova 1, Samara, 443011, Russia; Saint Petersburg University, 7/9 Universitetskaya nab., St. Petersburg, 199034, Russia E-mail adress: surkovmatveya@yandex.ru
Appendix
  1. 1.
    $\emptyset$
  2. 2.
    $\alpha_{4}$
  3. 3.
    $\alpha_{3}$
  4. 4.
    $\alpha_{3},~\alpha_{4}$
  5. 5.
    $\alpha_{3}+\alpha_{4}$
  6. 6.
    $\alpha_{2}$
  7. 7.
    $\alpha_{2},~\alpha_{3}+\alpha_{4}$
  8. 8.
    $\alpha_{2},~\alpha_{3}$
  9. 9.
    $\alpha_{2},~\alpha_{3},~\alpha_{4}$
  10. 10.
    $\alpha_{2},~\alpha_{4}$
  11. 11.
    $\alpha_{2}+\alpha_{3}$
  12. 12.
    $\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  13. 13.
    $\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  14. 14.
    $\alpha_{2}+\alpha_{3},~\alpha_{4}$
  15. 15.
    $\alpha_{2}+\alpha_{3}+\alpha_{4}$
  16. 16.
    $\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  17. 17.
    $\alpha_{2}+2\alpha_{3}$
  18. 18.
    $\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  19. 19.
    $\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  20. 20.
    $\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  21. 21.
    $\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  22. 22.
    $\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  23. 23.
    $\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{4}$
  24. 24.
    $\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  25. 25.
    $\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  26. 26.
    $\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  27. 27.
    $\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  28. 28.
    $\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  29. 29.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  30. 30.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  31. 31.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  32. 32.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  33. 33.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  34. 34.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}$
  35. 35.
    $\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  36. 36.
    $\alpha_{1}$
  37. 37.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  38. 38.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  39. 39.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  40. 40.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  41. 41.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  42. 42.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}$
  43. 43.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  44. 44.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  45. 45.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  46. 46.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3}$
  47. 47.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  48. 48.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  49. 49.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  50. 50.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  51. 51.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  52. 52.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{4}$
  53. 53.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  54. 54.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  55. 55.
    $\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  56. 56.
    $\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  57. 57.
    $\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  58. 58.
    $\alpha_{1},~\alpha_{2}+\alpha_{3}$
  59. 59.
    $\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  60. 60.
    $\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  61. 61.
    $\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  62. 62.
    $\alpha_{1},~\alpha_{2}$
  63. 63.
    $\alpha_{1},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  64. 64.
    $\alpha_{1},~\alpha_{2},~\alpha_{3}$
  65. 65.
    $\alpha_{1},~\alpha_{2},~\alpha_{3},~\alpha_{4}$
  66. 66.
    $\alpha_{1},~\alpha_{2},~\alpha_{4}$
  67. 67.
    $\alpha_{1},~\alpha_{3}+\alpha_{4}$
  68. 68.
    $\alpha_{1},~\alpha_{3}$
  69. 69.
    $\alpha_{1},~\alpha_{3},~\alpha_{4}$
  70. 70.
    $\alpha_{1},~\alpha_{4}$
  71. 71.
    $\alpha_{1}+\alpha_{2}$
  72. 72.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  73. 73.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  74. 74.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  75. 75.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  76. 76.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  77. 77.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  78. 78.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}$
  79. 79.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  80. 80.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  81. 81.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  82. 82.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  83. 83.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  84. 84.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  85. 85.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  86. 86.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  87. 87.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  88. 88.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  89. 89.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  90. 90.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  91. 91.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  92. 92.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  93. 93.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{3},~\alpha_{4}$
  94. 94.
    $\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  95. 95.
    $\alpha_{1}+\alpha_{2},~\alpha_{3}+\alpha_{4}$
  96. 96.
    $\alpha_{1}+\alpha_{2},~\alpha_{3}$
  97. 97.
    $\alpha_{1}+\alpha_{2},~\alpha_{3},~\alpha_{4}$
  98. 98.
    $\alpha_{1}+\alpha_{2},~\alpha_{4}$
  99. 99.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}$
  100. 100.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  101. 101.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  102. 102.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  103. 103.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  104. 104.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  105. 105.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  106. 106.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  107. 107.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  108. 108.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  109. 109.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  110. 110.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  111. 111.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  112. 112.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  113. 113.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  114. 114.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  115. 115.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{4}$
  116. 116.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  117. 117.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  118. 118.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  119. 119.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  120. 120.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  121. 121.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}$
  122. 122.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  123. 123.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  124. 124.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  125. 125.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  126. 126.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  127. 127.
    $\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{4}$
  128. 128.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  129. 129.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  130. 130.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  131. 131.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  132. 132.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  133. 133.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  134. 134.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  135. 135.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{3}$
  136. 136.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  137. 137.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  138. 138.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  139. 139.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  140. 140.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  141. 141.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}$
  142. 142.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  143. 143.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  144. 144.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  145. 145.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  146. 146.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  147. 147.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  148. 148.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{3}$
  149. 149.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  150. 150.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  151. 151.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  152. 152.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  153. 153.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  154. 154.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{3}$
  155. 155.
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  156. 156.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  157. 157.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  158. 158.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  159. 159.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  160. 160.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  161. 161.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  162. 162.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  163. 163.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  164. 164.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  165. 165.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  166. 166.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  167. 167.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  168. 168.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  169. 169.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  170. 170.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  171. 171.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  172. 172.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  173. 173.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  174. 174.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  175. 175.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  176. 176.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  177. 177.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  178. 178.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  179. 179.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  180. 180.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  181. 181.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  182. 182.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  183. 183.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  184. 184.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  185. 185.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  186. 186.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  187. 187.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  188. 188.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  189. 189.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  190. 190.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  191. 191.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  192. 192.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  193. 193.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  194. 194.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  195. 195.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{4}$
  196. 196.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{3}+\alpha_{4}$
  197. 197.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  198. 198.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  199. 199.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  200. 200.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  201. 201.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  202. 202.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  203. 203.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  204. 204.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  205. 205.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  206. 206.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  207. 207.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{4}$
  208. 208.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  209. 209.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  210. 210.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  211. 211.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  212. 212.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  213. 213.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  214. 214.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  215. 215.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  216. 216.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  217. 217.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  218. 218.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  219. 219.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  220. 220.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{4}$
  221. 221.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  222. 222.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  223. 223.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  224. 224.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  225. 225.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}$
  226. 226.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  227. 227.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  228. 228.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  229. 229.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  230. 230.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  231. 231.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  232. 232.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  233. 233.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  234. 234.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  235. 235.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  236. 236.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  237. 237.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  238. 238.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  239. 239.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  240. 240.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  241. 241.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  242. 242.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}$
  243. 243.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  244. 244.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  245. 245.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  246. 246.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  247. 247.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  248. 248.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  249. 249.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  250. 250.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  251. 251.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  252. 252.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  253. 253.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  254. 254.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  255. 255.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  256. 256.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  257. 257.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  258. 258.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  259. 259.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  260. 260.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  261. 261.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  262. 262.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  263. 263.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  264. 264.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  265. 265.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  266. 266.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  267. 267.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  268. 268.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  269. 269.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  270. 270.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  271. 271.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  272. 272.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  273. 273.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  274. 274.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  275. 275.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  276. 276.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  277. 277.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  278. 278.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  279. 279.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  280. 280.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  281. 281.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  282. 282.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  283. 283.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  284. 284.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  285. 285.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  286. 286.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  287. 287.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  288. 288.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  289. 289.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  290. 290.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  291. 291.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  292. 292.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  293. 293.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  294. 294.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  295. 295.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  296. 296.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  297. 297.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  298. 298.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  299. 299.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  300. 300.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  301. 301.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  302. 302.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  303. 303.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  304. 304.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  305. 305.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  306. 306.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  307. 307.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}$
  308. 308.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  309. 309.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  310. 310.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  311. 311.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  312. 312.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  313. 313.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  314. 314.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  315. 315.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  316. 316.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  317. 317.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  318. 318.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  319. 319.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  320. 320.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  321. 321.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  322. 322.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  323. 323.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  324. 324.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  325. 325.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  326. 326.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  327. 327.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  328. 328.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  329. 329.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  330. 330.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  331. 331.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  332. 332.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  333. 333.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  334. 334.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  335. 335.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  336. 336.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  337. 337.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  338. 338.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  339. 339.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  340. 340.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  341. 341.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  342. 342.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  343. 343.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  344. 344.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  345. 345.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  346. 346.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  347. 347.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  348. 348.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  349. 349.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  350. 350.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  351. 351.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  352. 352.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  353. 353.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}$
  354. 354.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  355. 355.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  356. 356.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  357. 357.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  358. 358.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  359. 359.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  360. 360.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  361. 361.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  362. 362.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  363. 363.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  364. 364.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  365. 365.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  366. 366.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  367. 367.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  368. 368.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  369. 369.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  370. 370.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  371. 371.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  372. 372.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  373. 373.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  374. 374.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  375. 375.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  376. 376.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  377. 377.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  378. 378.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  379. 379.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  380. 380.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  381. 381.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  382. 382.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}$
  383. 383.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  384. 384.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}$
  385. 385.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  386. 386.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  387. 387.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  388. 388.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  389. 389.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  390. 390.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  391. 391.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  392. 392.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  393. 393.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  394. 394.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}$
  395. 395.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  396. 396.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  397. 397.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  398. 398.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  399. 399.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  400. 400.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  401. 401.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{3}$
  402. 402.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  403. 403.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  404. 404.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  405. 405.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{3}$
  406. 406.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{3}$
  407. 407.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  408. 408.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  409. 409.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  410. 410.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  411. 411.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  412. 412.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  413. 413.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  414. 414.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  415. 415.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  416. 416.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  417. 417.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  418. 418.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  419. 419.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  420. 420.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  421. 421.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  422. 422.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{3}$
  423. 423.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  424. 424.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  425. 425.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  426. 426.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}$
  427. 427.
    $\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  428. 428.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}$
  429. 429.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  430. 430.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  431. 431.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  432. 432.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  433. 433.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  434. 434.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  435. 435.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  436. 436.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  437. 437.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  438. 438.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  439. 439.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  440. 440.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{4}$
  441. 441.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  442. 442.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  443. 443.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  444. 444.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  445. 445.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  446. 446.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  447. 447.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  448. 448.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  449. 449.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  450. 450.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3},~\alpha_{4}$
  451. 451.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  452. 452.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  453. 453.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  454. 454.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  455. 455.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  456. 456.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  457. 457.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  458. 458.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{4}$
  459. 459.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  460. 460.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  461. 461.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  462. 462.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3},~\alpha_{4}$
  463. 463.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  464. 464.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  465. 465.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}$
  466. 466.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3},~\alpha_{4}$
  467. 467.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{4}$
  468. 468.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  469. 469.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  470. 470.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  471. 471.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  472. 472.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  473. 473.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{4}$
  474. 474.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  475. 475.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  476. 476.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  477. 477.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3},~\alpha_{4}$
  478. 478.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  479. 479.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  480. 480.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  481. 481.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{3},~\alpha_{4}$
  482. 482.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  483. 483.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  484. 484.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  485. 485.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  486. 486.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  487. 487.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  488. 488.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  489. 489.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  490. 490.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  491. 491.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  492. 492.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  493. 493.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  494. 494.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  495. 495.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  496. 496.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  497. 497.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  498. 498.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  499. 499.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}$
  500. 500.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  501. 501.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  502. 502.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  503. 503.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  504. 504.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  505. 505.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  506. 506.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  507. 507.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  508. 508.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}$
  509. 509.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  510. 510.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  511. 511.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  512. 512.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  513. 513.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  514. 514.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{3}$
  515. 515.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  516. 516.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  517. 517.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  518. 518.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  519. 519.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  520. 520.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  521. 521.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  522. 522.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  523. 523.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  524. 524.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  525. 525.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  526. 526.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  527. 527.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  528. 528.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  529. 529.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  530. 530.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  531. 531.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}$
  532. 532.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  533. 533.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  534. 534.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  535. 535.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  536. 536.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  537. 537.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  538. 538.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  539. 539.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  540. 540.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  541. 541.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  542. 542.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  543. 543.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  544. 544.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  545. 545.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  546. 546.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  547. 547.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  548. 548.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  549. 549.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  550. 550.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  551. 551.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  552. 552.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  553. 553.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  554. 554.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  555. 555.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  556. 556.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3}$
  557. 557.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  558. 558.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  559. 559.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  560. 560.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  561. 561.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  562. 562.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}$
  563. 563.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  564. 564.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  565. 565.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  566. 566.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  567. 567.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  568. 568.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  569. 569.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  570. 570.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  571. 571.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  572. 572.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  573. 573.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  574. 574.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  575. 575.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  576. 576.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  577. 577.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  578. 578.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  579. 579.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  580. 580.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  581. 581.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  582. 582.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  583. 583.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  584. 584.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  585. 585.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  586. 586.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  587. 587.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  588. 588.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  589. 589.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  590. 590.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}$
  591. 591.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  592. 592.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  593. 593.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}$
  594. 594.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  595. 595.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  596. 596.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  597. 597.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  598. 598.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  599. 599.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4}$
  600. 600.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  601. 601.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  602. 602.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3}$
  603. 603.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  604. 604.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  605. 605.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  606. 606.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  607. 607.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  608. 608.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}$
  609. 609.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  610. 610.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  611. 611.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  612. 612.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  613. 613.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  614. 614.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}$
  615. 615.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  616. 616.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3}$
  617. 617.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  618. 618.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  619. 619.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  620. 620.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  621. 621.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  622. 622.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3}$
  623. 623.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  624. 624.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  625. 625.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}$
  626. 626.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{3}$
  627. 627.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  628. 628.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  629. 629.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  630. 630.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  631. 631.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  632. 632.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}$
  633. 633.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  634. 634.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{3}$
  635. 635.
    $\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  636. 636.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4}$
  637. 637.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  638. 638.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  639. 639.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{4}$
  640. 640.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  641. 641.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  642. 642.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  643. 643.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  644. 644.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{4}$
  645. 645.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  646. 646.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  647. 647.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{4}$
  648. 648.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  649. 649.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}$
  650. 650.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  651. 651.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  652. 652.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  653. 653.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1}$
  654. 654.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  655. 655.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  656. 656.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{4}$
  657. 657.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  658. 658.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2}$
  659. 659.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{2},~\alpha_{4}$
  660. 660.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{4}$
  661. 661.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  662. 662.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  663. 663.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{4}$
  664. 664.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  665. 665.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{2}$
  666. 666.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  667. 667.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+\alpha_{4},~\alpha_{4}$
  668. 668.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4}$
  669. 669.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}$
  670. 670.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  671. 671.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  672. 672.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  673. 673.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  674. 674.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  675. 675.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  676. 676.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}$
  677. 677.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  678. 678.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3}$
  679. 679.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  680. 680.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}$
  681. 681.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  682. 682.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  683. 683.
    $\alpha_{1}+2\alpha_{2}+3\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  684. 684.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4}$
  685. 685.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  686. 686.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}$
  687. 687.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  688. 688.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}$
  689. 689.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2}$
  690. 690.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  691. 691.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}$
  692. 692.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  693. 693.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  694. 694.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  695. 695.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3}$
  696. 696.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  697. 697.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}$
  698. 698.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  699. 699.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  700. 700.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  701. 701.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  702. 702.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}$
  703. 703.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3}$
  704. 704.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  705. 705.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}$
  706. 706.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  707. 707.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{4}$
  708. 708.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  709. 709.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  710. 710.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  711. 711.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  712. 712.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  713. 713.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  714. 714.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  715. 715.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  716. 716.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  717. 717.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  718. 718.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  719. 719.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  720. 720.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  721. 721.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2}$
  722. 722.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{2},~\alpha_{4}$
  723. 723.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{4}$
  724. 724.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  725. 725.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  726. 726.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  727. 727.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  728. 728.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  729. 729.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{4}$
  730. 730.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+2\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  731. 731.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  732. 732.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}$
  733. 733.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  734. 734.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  735. 735.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  736. 736.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}$
  737. 737.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  738. 738.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  739. 739.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}$
  740. 740.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2}$
  741. 741.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  742. 742.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}$
  743. 743.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2},~\alpha_{4}$
  744. 744.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  745. 745.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  746. 746.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  747. 747.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}$
  748. 748.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  749. 749.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2}$
  750. 750.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  751. 751.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{4}$
  752. 752.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  753. 753.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  754. 754.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3}$
  755. 755.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  756. 756.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  757. 757.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  758. 758.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  759. 759.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  760. 760.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3}$
  761. 761.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  762. 762.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  763. 763.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  764. 764.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2}$
  765. 765.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{2},~\alpha_{4}$
  766. 766.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{4}$
  767. 767.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  768. 768.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}$
  769. 769.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  770. 770.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2}$
  771. 771.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{2},~\alpha_{4}$
  772. 772.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  773. 773.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  774. 774.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{4}$
  775. 775.
    $\alpha_{1}+2\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  776. 776.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$
  777. 777.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  778. 778.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  779. 779.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  780. 780.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{1}$
  781. 781.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  782. 782.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}$
  783. 783.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}$
  784. 784.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  785. 785.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1}$
  786. 786.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{3}$
  787. 787.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{3}$
  788. 788.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  789. 789.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3}$
  790. 790.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  791. 791.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  792. 792.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2}$
  793. 793.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1}$
  794. 794.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{1}$
  795. 795.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3}$
  796. 796.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  797. 797.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}$
  798. 798.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{4}$
  799. 799.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  800. 800.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2}$
  801. 801.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{1}$
  802. 802.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  803. 803.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  804. 804.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{4}$
  805. 805.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{3}+\alpha_{4}$
  806. 806.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  807. 807.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  808. 808.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}$
  809. 809.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  810. 810.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  811. 811.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1},~\alpha_{4}$
  812. 812.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  813. 813.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  814. 814.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  815. 815.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$
  816. 816.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1}$
  817. 817.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{1},~\alpha_{3}$
  818. 818.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  819. 819.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3}$
  820. 820.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1}$
  821. 821.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  822. 822.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  823. 823.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{1},~\alpha_{4}$
  824. 824.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  825. 825.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  826. 826.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}+\alpha_{3},~\alpha_{4}$
  827. 827.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2}$
  828. 828.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1}$
  829. 829.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  830. 830.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{3}$
  831. 831.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{3},~\alpha_{4}$
  832. 832.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{1},~\alpha_{4}$
  833. 833.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{3}+\alpha_{4}$
  834. 834.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{3}$
  835. 835.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{3},~\alpha_{4}$
  836. 836.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}+\alpha_{2},~\alpha_{4}$
  837. 837.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1}$
  838. 838.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3}+\alpha_{4}$
  839. 839.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3}$
  840. 840.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{3},~\alpha_{4}$
  841. 841.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{1},~\alpha_{4}$
  842. 842.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  843. 843.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  844. 844.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{3},~\alpha_{4}$
  845. 845.
    $\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{4}$
  846. 846.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$
  847. 847.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
  848. 848.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  849. 849.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  850. 850.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  851. 851.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  852. 852.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}$
  853. 853.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  854. 854.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4}$
  855. 855.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}$
  856. 856.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3}$
  857. 857.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  858. 858.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}$
  859. 859.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2}$
  860. 860.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  861. 861.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  862. 862.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{2},~\alpha_{4}$
  863. 863.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4}$
  864. 864.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  865. 865.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+2\alpha_{3},~\alpha_{4}$
  866. 866.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$
  867. 867.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}$
  868. 868.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3}$
  869. 869.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4}$
  870. 870.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
  871. 871.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}+\alpha_{3},~\alpha_{4}$
  872. 872.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2}$
  873. 873.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}+\alpha_{4}$
  874. 874.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3}$
  875. 875.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3},~\alpha_{4}$
  876. 876.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{4}$
  877. 877.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{3}+\alpha_{4}$
  878. 878.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{3}$
  879. 879.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{3},~\alpha_{4}$
  880. 880.
    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{4}$
import java.util.ArrayList;
import java.util.Scanner;
public class Carriers {
public static void main(String[] args) {new Carr();}}
class Carr{
    Carr(){
        a = new int[25][25];
        m = new int[25][25];
        a[23][10]=24;
        a[22][4]=23; a[22][11]=24;
        a[21][2]=22; a[21][5]=23; a[21][12]=24;
        a[20][1]=21; a[20][3]=22; a[20][6]=23; a[20][13]=24;
        a[19][2]=21; a[19][7]=23; a[19][14]=24;
        a[18][1]=19; a[18][2]=20; a[18][3]=21; a[18][8]=23; a[18][15]=24;
        a[17][1]=18; a[17][3]=20; a[17][9]=23; a[17][16]=24;
        a[16][4]=19; a[16][5]=21; a[16][7]=22;
        a[15][1]=16; a[15][4]=18; a[15][5]=20; a[15][6]=21; a[15][8]=22;
        a[14][1]=15; a[14][4]=17; a[14][6]=20; a[14][9]=22;
        a[13][2]=15; a[13][3]=16; a[13][5]=18; a[13][6]=19; a[13][7]=20; a[13][8]=21;
        a[12][1]=13; a[12][2]=14; a[12][3]=15; a[12][5]=17; a[12][6]=18; a[12][8]=20; a[12][9]=21;
        a[11][2]=12; a[11][3]=13; a[11][7]=17; a[11][8]=18; a[11][9]=19;
        a[10][4]=11; a[10][5]=12; a[10][6]=13; a[10][7]=14; a[10][8]=15; a[10][9]=16;
        a[8][1]=9;
        a[7][1]=8;
        a[6][2]=8; a[6][3]=9;
        a[5][1]=6; a[5][2]=7; a[5][3]=8;
        a[4][2]=5; a[4][3]=6;
        a[2][1]=3;
$Ψ$
        m[23][10]=-1;
        m[22][4]=-1; m[22][11]=-1;
        m[21][2]=-2; m[21][5]=-2; m[21][12]=-2;
        m[20][1]=-1; m[20][3]=-2; m[20][6]=-2; m[20][13]=-2;
        m[19][2]=-1; m[19][7]=1; m[19][14]=1;
        m[18][1]=-2; m[18][2]=-1; m[18][3]=-1; m[18][8]=2; m[18][15]=2;
        m[17][1]=1; m[17][3]=1; m[17][9]=1; m[17][16]=1;
        m[16][4]=-1; m[16][5]=1; m[16][7]=1;
        m[15][1]=-2; m[15][4]=-1; m[15][5]=1; m[15][6]=1; m[15][8]=2;
        m[14][1]=1; m[14][4]=-1; m[14][6]=-1; m[14][9]=1;
        m[13][2]=-1; m[13][3]=2; m[13][5]=-1; m[13][6]=2; m[13][7]=-1; m[13][8]=1;
        m[12][1]=1; m[12][2]=-2; m[12][3]=-1; m[12][5]=-2; m[12][6]=-1; m[12][8]=1; m[12][9]=1;
        m[11][2]=1; m[11][3]=1; m[11][7]=-1; m[11][8]=-1; m[11][9]=-1;
        m[10][4]=1; m[10][5]=1; m[10][6]=1; m[10][7]=1; m[10][8]=1; m[10][9]=1;
        m[8][1]=-2;
        m[7][1]=1;
        m[6][2]=-1; m[6][3]=2;
        m[5][1]=1; m[5][2]=-2; m[5][3]=-1;
        m[4][2]=1; m[4][3]=1;
        m[2][1]=1;
$Ψ$
        b = new String[25];
        b[1]="a_4"; b[2]="a_3"; b[3]="a_3+a_4"; b[4]="a_2"; b[5]="a_2+a_3"; b[6]="a_2+a_3+a_4"; b[7]="a_2+2a_3";
        b[8]="a_2+2a_3+a_4"; b[9]="a_2+2a_3+2a_4"; b[10]="a_1"; b[11]="a_1+a_2"; b[12]="a_1+a_2+a_3";
        b[13]="a_1+a_2+a_3+a_4"; b[14]="a_1+a_2+2a_3"; b[15]="a_1+a_2+2a_3+a_4"; b[16]="a_1+a_2+2a_3+2a_4";
        b[17]="a_1+2a_2+2a_3"; b[18]="a_1+2a_2+2a_3+a_4"; b[19]="a_1+2a_2+2a_3+2a_4"; b[20]="a_1+2a_2+3a_3+a_4";
        b[21]="a_1+2a_2+3a_3+2a_4"; b[22]="a_1+2a_2+4a_3+2a_4"; b[23]="a_1+3a_2+4a_3+2a_4"; b[24]="2a_1+3a_2+4a_3+2a_4";
        c = new double[25][25];
        lambda = new double[25];
        for(int i=1; i<=24; i++) { lambda[i] = 1;}
        roots = new ArrayList<>();
        roots.add(24);
        n = 0;
        depend = new ArrayList<>();
        Scanner in = new Scanner(System.in);
[…]
        x = in.nextInt();
        in.close();
$Ψ$
        while(true) {
            last = roots.get(roots.size()-1);
            for(int i=24; i>=last; i--) {
                for(int j=1; j<=24; j++) {
                if(roots.contains(a[i][j])||roots.contains(a[j][i])) { c[i][j]=(m[i][j]-m[j][i])*lambda[a[i][j]+a[j][i]];}
                else { c[i][j]=0; }}}
            depend.clear();
            q = true;
            while(q) {
                q = false;
                for(int i=24; i>last; i--) {
                    p = true; dep = 0;
                    for(int j=1; j<=24; j++) {
                        if(c[i][j]!=0 && !depend.contains(j)) {
                            if(dep==0) { dep = j; }
                            else { p = false; break; }}}
                    if(dep!=0 && p) {
                        depend.add(dep); q = true; }}}
            rank = true;
            if(x==0) {
                for(int j=1; j<=24; j++) {
                    if(c[last][j]!=0 && !depend.contains(j)) { rank = false; break; }}}
            else {
                rowold = last;
                colold = 0;
                while(true) {
                    count = 0;
                    for(int j=1; j<=24; j++) {
                        if(c[rowold][j]!=0 && !depend.contains(j) && j!=colold) {
                            rank = false; colnew = j;
                            for(int i=24; i>last; i--) {
                                if(c[i][j]!=0 && i!=rowold) {
                                    rank = true; count++; rownew = i; }}
                            if (!rank) { break; }}}
                    if (!rank || count!=1) { break; }
                    else { rowold = rownew; colold = colnew; }}}
            if(rank) {
                n++;
                for(int h: roots) { System.out.print(b[h]+" "); }
                System.out.println();
                if(last!=1) { roots.add(last-1); }}
            else {
                if(last!=1) {
                    roots.remove((Integer) last);
                    roots.add(last-1); }}
            if(last==1) {
                if(roots.size()!=1) {
                    roots.remove((Integer) 1);
                    last = roots.get(roots.size()-1);
                    roots.remove((Integer) last);
                    roots.add(last-1);}
                else { break; }}}
        System.out.println("empty");
        n++;
        System.out.print(n); }
    int x; int[][]a; int[][]m; String[]b; double[][]c; double[]lambda; ArrayList<Integer> roots; ArrayList<Integer> depend;
    boolean q; boolean p; int dep; int last; boolean rank; int n; int count; int rowold; int colold; int rownew; int colnew;

### v4

@@
 Abstract
-Let NN be the maximal unipotent subgroup in the simple algebraic group of type $\Phi$.
-It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called coadjoint.
-Such orbits play the key role in the A.A.
+Let NN be the maximal unipotent subgroup of the simple algebraic group of type $\Phi$.
+It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called the coadjoint action.
+Such orbits play a key role in A.A.
 Kirillov’s orbit method.
@@
 1. Introduction and the main result
-Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra in $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space to $\mathfrak{n}$.
-The group NN is also a maximal unipotent subgroup in the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup in $G$.
-The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of the group NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$.
+Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra of $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space of $\mathfrak{n}$.
+The group NN is also a maximal unipotent subgroup of the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup of $G$.
+The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$.
 According to the orbit method introduced by A.A.
@@
 They are discussed in detail in the works [An95], [AN06], [Ko12], [Ko13], [Pa08], [IPa09], [Ig09], [Ig11], [Ig12], [IS21] and [IS23].
-Explicitly, the coadjoint action looks as follows:
+Explicitly, the coadjoint action is given by:
 $\begin{split}(\exp(x).\lambda)(y)=\lambda(\exp(-\mathrm{ad}_{x})(y))=\lambda(y)-\lambda([x,y])+\dfrac{1}{2!}\lambda([x,[x,y]])-\ldots,\\ x,y\in\mathfrak{n},~\lambda\in\mathfrak{n}^{*}.\end{split}$
-Let $\{e_{\alpha},\alpha\in\Phi^{+}\}$ be a basis in $\mathfrak{n}$ consisting of root vectors, and $\{e_{\alpha}^{*},\alpha\in\Phi^{+}\}$ be the dual basis in $\mathfrak{n}^{*}$.
-The support of a linear form $\lambda\in\mathfrak{n}^{*}$ is defined as the subset of the set of positive roots of the form
+Let $\{e_{\alpha},\alpha\in\Phi^{+}\}$ be a basis of $\mathfrak{n}$ consisting of root vectors, and let $\{e_{\alpha}^{*},\alpha\in\Phi^{+}\}$ be the dual basis in $\mathfrak{n}^{*}$.
+The support of a linear form $\lambda\in\mathfrak{n}^{*}$ is defined as the subset of positive roots given by
 $\textup{supp}(\lambda)=\{\alpha\in\Phi^{+}\colon\lambda(e_{\alpha})\neq 0\}.$
-Our goal is to describe the coadjoint orbits in the case of $\Phi=F_{4}$.
+Our goal is to describe the coadjoint orbits for the case of $\Phi=F_{4}$.
 The description will be given in terms of supports of canonical forms.
 A list of supports will be provided with the following properties.
-  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ with supports from the given list (possibly, having the same support), the orbits of these forms $N.\lambda_{1}$ and $N.\lambda_{2}$ do not coincide.
-  • For any linear form $\lambda\in\mathfrak{n}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
-In other words, there exists a unique linear form with a support from the given list on each coadjoint orbit.
-The set of simple roots in $F_{4}$ can be identified with the following subset of $\mathbb{R}^{4}$:
+  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ whose supports belong to the given list (even if the supports coincide), the orbits $N.\lambda_{1}$ and $N.\lambda_{2}$ are different.
+  • For any linear form $\lambda\in\mathfrak{n}^{*}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
+In other words, each coadjoint orbit contains a unique linear form whose support belongs to the given list.
+The set of simple roots of $F_{4}$ can be identified with the following subset of $\mathbb{R}^{4}$:
 $\Delta=\{\alpha_{1}=\varepsilon_{2}-\varepsilon_{3},~\alpha_{2}=\varepsilon_{3}-\varepsilon_{4},~\alpha_{3}=\varepsilon_{4},~\alpha_{4}=\frac{1}{2}(\varepsilon_{1}-\varepsilon_{2}-\varepsilon_{3}-\varepsilon_{4})\}.$
-Here $\{\varepsilon_{i}\}_{i=1}^{4}$ is the standard basis of $\mathbb{R}^{4}$ (with the standard scalar product).
-Recall that there exists the natural order on $\Phi^{+}$ defined as follows: $\alpha>\beta$ if $\alpha-\beta$ is a sum of positive roots.
+Here $\{\varepsilon_{i}\}_{i=1}^{4}$ is the standard basis of $\mathbb{R}^{4}$ (endowed with the standard scalar product).
+Recall that there exists a natural partial order on $\Phi^{+}$ defined as follows: $\alpha>\beta$ if $\alpha-\beta$ is a sum of positive roots.
 Let us fix a linear lexicographic order $\succ$ on the positive roots of $F_{4}$.
-We will say that $\alpha=a_{1}\alpha_{1}+a_{2}\alpha_{2}+a_{3}\alpha_{3}+a_{4}\alpha_{4}\succ\beta=b_{1}\alpha_{1}+b_{2}\alpha_{2}+b_{3}\alpha_{3}+b_{4}\alpha_{4}$ if $a_{1}>b_{1}$ or $a_{1}=b_{1},a_{2}>b_{2}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}>b_{3}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}=b_{3},a_{4}>b_{4}$.
+We say that $\alpha=a_{1}\alpha_{1}+a_{2}\alpha_{2}+a_{3}\alpha_{3}+a_{4}\alpha_{4}\succ\beta=b_{1}\alpha_{1}+b_{2}\alpha_{2}+b_{3}\alpha_{3}+b_{4}\alpha_{4}$ if $a_{1}>b_{1}$ or $a_{1}=b_{1},a_{2}>b_{2}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}>b_{3}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}=b_{3},a_{4}>b_{4}$.
 Thus, the roots will be ordered as follows:
@@
 By definition, a subset $D$ of $\Phi^{+}$ is called a non-singular rook placement if $\alpha-\beta\notin\Phi^{+}$ for all distinct $\alpha,\beta\in D$.
-From this it follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$.
+It follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$.
 Given a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$, we put
 $f_{D,\xi}=\sum_{\alpha\in D}\xi(\alpha)e_{\alpha}^{*}\in\mathfrak{n}^{*}.$
-We say that the coadjoint orbit $\Omega_{D,\xi}$ of the linear form $f_{D,\xi}$ is associated with the rook placement $D$ and the map $\xi$.
-In essence, $f_{D,\xi}$ is just a linear form with support $D$.
-The work [IS23] proves that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$.
-But in fact if $D$ is a non-singular rook placement then $\lambda=f_{D,\xi}\in S$.
-Because if $\gamma\in D$ then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$.
-Hence, in the matrix $A_{\lambda,\gamma}$ the line with the number $\gamma$ consists of zeros, and condition (1) is true for $\gamma\in D$.
-It follows from this that for distinct linear forms with support $D$ the corresponding coadjoint orbits do not coincide.
-The result is similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is true for the root system $\Phi=G_{2}$ and field $\mathbb{C}$.
-In the same work [IS23] the equations of the orbits of basic subvarieties $\mathcal{O}_{D,\xi}$ were written out.
+We say that the coadjoint orbit $\Omega_{D,\xi}$ of the linear form $f_{D,\xi}$ is associated with the rook placement $D$ and the map $\xi$.Thus, $f_{D,\xi}$ is a linear form with support $D$.
+In [IS23] it is proved that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$.
+Moreover, if $D$ is a non-singular rook placement, then $\lambda=f_{D,\xi}$ belongs to $S$.
+Indeed, if $\gamma\in D$, then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$.
+Hence, in the matrix $A_{\lambda,\gamma}$ the row corresponding to $\gamma$ is zero, and condition (1) is true for $\gamma\in D$.
+Therefore, distinct linear forms with the same support $D$ lie in different coadjoint orbits.
+A result similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for the root system $\Phi=G_{2}$ and field $\mathbb{C}$.
+In the same work [IS23], the equations defining the basic subvarieties $\mathcal{O}_{D,\xi}$ are given.
 By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
 $\mathcal{O}_{D,\xi}=\sum_{\alpha\in D}\Omega_{\{\alpha\},\xi_{\alpha}},$
-where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. The work also proved the Andrè’s stratification for $G_{2}$:
+where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. That work also proves AndrГ©’s stratification for $G_{2}$:
 $\mathfrak{n}^{*}=\bigsqcup_{D,\xi}\mathcal{O}_{D,\xi},$
@@
 Let us introduce the notation of structural constants as in the work [IS23]: $e_{\alpha},e_{\beta}]=c_{1}\cdot e_{\alpha+\beta}$, $e_{\alpha},e_{\alpha+\beta}]=c_{2}\cdot e_{2\alpha+\beta}$, $e_{\alpha},e_{2\alpha+\beta}]=c_{3}\cdot e_{3\alpha+\beta}$, $e_{3\alpha+\beta},e_{\beta}]=c_{4}\cdot e_{3\alpha+2\beta}$, $e_{\alpha+\beta},e_{2\alpha+\beta}]=c_{5}\cdot e_{3\alpha+2\beta}$.
-For almost every rook placement $D$ it is true that $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$.
-This is only wrong for $D=\{\alpha+\beta,3\alpha+\beta\}$.
-In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon\{\beta,\alpha+\beta,3\alpha+\beta\}\to\mathbb{C}^{\times}$ and $\xi^{\prime}|_{D}=\xi$.
-Now, let us write out the equations of the orbits for $D=\{\alpha+\beta,3\alpha+\beta\}$.
-For a linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}\in\Omega_{D,\xi}$ we have the following system of equations.
+For almost every rook placement $D$, we have $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$.
+The only exception is the case $D=\{\alpha+\beta,3\alpha+\beta\}$.
+In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon{\beta,\alpha+\beta,3\alpha+\beta}\to\mathbb{C}^{\times}$ is such that its restriction to $D$ coincides with $\xi$.
+Let us write down the equations defining the orbits for $D={\alpha+\beta,3\alpha+\beta}$.
+A linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}$ lies in $\Omega_{D,\xi}$ if and only if it satisfies the following system:
 $\begin{split}6c_{3}^{2}\lambda_{3\alpha+\beta}^{2}\lambda_{\beta}-6c_{1}c_{3}\lambda_{2\alpha+\beta}\lambda_{3\alpha+\beta}\lambda_{\alpha+\beta}+5c_{1}c_{2}\lambda_{2\alpha+\beta}^{3}=&0,\\ 2c_{3}\lambda_{\alpha+\beta}\lambda_{3\alpha+\beta}-c_{2}\lambda_{2\alpha+\beta}^{2}=&2c_{3}\xi(\alpha+\beta)\xi(3\alpha+\beta),\\ \lambda_{3\alpha+\beta}=&\xi(3\alpha+\beta),\\ \lambda_{3\alpha+2\beta}=&0.\end{split}$
@@
 $\begin{split}\varnothing,\{\alpha\},\{\beta\},\{\alpha,\beta\},\{\alpha+\beta\},\{2\alpha+\beta\},\{2\alpha+\beta,\beta\},\{3\alpha+\beta\},\\ \{3\alpha+\beta,\alpha+\beta\},\{3\alpha+\beta,\alpha+\beta,\beta\},\{3\alpha+\beta,\beta\},\{3\alpha+2\beta\},\{3\alpha+2\beta,\alpha\}.\end{split}$
-The orbits of these forms correspond to all orbits whose equations are written out in the article [IS23] and here.
+The orbits of these forms correspond precisely to all the orbits whose equations are provided in [IS23] and in the present work.
 Funding. The work is supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075–15–2022–287).
-Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during writing the paper.
+Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during the writing of this paper.
 2. Description of the set of canonical forms
-In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and its use.
-At first glance, it seems that it does not have a good description.
+In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and to demonstrate its use.
+At first glance, it might seem that this set does not admit a nice description.
 In fact, this is not the case.
-The set $S$ is very well described in terms of supports.
-In the Appendix, a list of 880 supports will be presented.
+The set $S$ can be described very neatly in terms of supports.
+A list of 880 supports is presented in the Appendix.
 Proposition 2.1. The set $S$ consists of all linear forms with supports from the list and forms with one of the three supports
 […]
-whose coordinates satisfy the equation
+and of the forms with one of the three supports listed below, provided their coordinates satisfy the equation
 $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}.$
-Here, we mean the coordinates of linear forms in the dual basis, $\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*}$.
+Here $\lambda=\sum_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*}$, and $\lambda_{\alpha}$ are the coordinates with respect to the dual basis.
 This section is devoted to the proof of this proposition.
-In the sequel, we will use the following table.
-It is a $24\times 24$ table, its rows and columns are numbered by positive roots.
-We write $\beta-\alpha$ in the cell $(\alpha,\beta)$ if $\beta-\alpha$ is the sum of positive roots, otherwise the cell is empty.
-The main diagonal contains zeros, it is easy to see that all cells below will be empty.
-Below we present the upper left corner of this table indexed by the positive roots whose decomposition does not contain $\alpha_{1}$.
-If at position $(\alpha,\beta)$ the difference $\beta-\alpha$ is also a positive root, this cell is highlighted in orange.
-Almost all examples will be considered specifically on this part of the table.
-Four columns are highlighted in blue because they will be encountered most often in the examples.
+In what follows, we will use the following table.
+It is a $24\times 24$ table whose rows and columns are indexed by the positive roots.
+In the cell at the intersection of row $\alpha$ and column $\beta$, we write $\beta-\alpha$ if $\beta-\alpha$ can be expressed as a sum of positive roots; otherwise, the cell is left blank.
+The main diagonal contains zeros.
+It is easy to see that all cells below the diagonal are empty.
+Below we show the upper-left corner of this table, corresponding to those positive roots whose simple root expansion does not contain $\alpha_{1}$.
+A cell at position $(\alpha,\beta)$ is highlighted in orange if the difference $\beta-\alpha$ is itself a positive root.
+Almost all examples will involve specifically this part of the table.
+Four columns are highlighted in blue, as they will appear most frequently in the examples.
 Table 1
@@
 To describe the set $S$, we wrote a program, which is included in the Appendix.
-The program output supports that satisfy a sufficient condition for being included in the list, or supports that satisfy a necessary condition, these will be described below.
+The program outputs supports that satisfy a sufficient condition for being included in the list, and supports that satisfy a necessary condition; these conditions will be described below.
 First, let us reformulate condition (1).
-The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: let $x,y\in\mathfrak{n},\lambda\in\mathfrak{n}^{*}$, then $(x.\lambda)(y):=-\lambda([x,y])$.
+The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: for $x,y\in\mathfrak{n}$ and $\lambda\in\mathfrak{n}^{*}$, define $(x.\lambda)(y):=-\lambda([x,y])$.
 This action is called tangent.
-Let us expand $x\in\mathfrak{n}$ in the basis:
+Let us expand $x\in\mathfrak{n}$ in the chosen basis:
 $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}.$
@@
 $(x.\lambda)(e_{\beta})=-\lambda([\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},e_{\beta}])=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
-We can consider these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$.
+We can view these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$.
 By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
@@
 Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
-We will show that such $\lambda$ lies in $S$.
-To visualize how the orbit of the tangent action looks like, let us refer to the table.
+We will show that such a $\lambda$ lies in $S$.
+To visualize what the orbit of the tangent action looks like, let us refer to the table.
 The highlighted blue columns correspond to the roots in $D$.
-The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, this is equivalent to $\alpha+\beta\in D$.
-It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row numbered $\beta$ in the highlighted column.
+The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, which is equivalent to $\alpha+\beta\in D$.
+It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row corresponding to $\beta$ in the highlighted column.
 For example, the combination $(x.\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
@@
 It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x.\lambda)(e_{\alpha_{4}})$.
-It is just because $x_{\alpha_{3}}$ as a linear combination was expressed through $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient.
+This is simply because $x_{\alpha_{3}}$ as a variable can be expressed in terms of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient.
 The Sufficient condition below is based on this idea.
@@
 Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$.
-Starting from the support $D$ and the root $\gamma$, we will construct the set $\zeta_{D,\gamma}$ consisting of positive roots.
+Starting from the support $D$ and the root $\gamma$, we will construct a set $\zeta_{D,\gamma}$ of positive roots.
 We will add roots to $\zeta_{D,\gamma}$ one by one.
-We will add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$ then $\alpha\in\zeta_{D,\gamma}$.
+We add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in\zeta_{D,\gamma}$.
 Using this procedure, we will add to $\zeta_{D,\gamma}$ the maximum possible number of roots.
-Note that $\zeta_{D,\gamma}$ does not depend on the values of the coordinates of the form $\lambda$ in the dual basis, but only on its support $D$.
-The set $\zeta_{D,\gamma}$ also does not depend on the order in which to add the roots.
-Indeed, let us define the sets $R_{n}$ for $n\in\mathbb{N}$.
-The root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$ then $\alpha\in R_{i}$ for $i<n$.
-It is clear that we must add all the roots of $R_{n},n\in\mathbb{N}$ to $\zeta_{D,\gamma}$, and we cannot add anything else.
+Note that $\zeta_{D,\gamma}$ does not depend on the specific values of the coordinates of $\lambda$ in the dual basis, but only on its support $D$.
+The set $\zeta_{D,\gamma}$ also does not depend on the order in which we add the roots.
+Indeed, define sets $R_{n}$ for $n\in\mathbb{N}$.
+A root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in R_{i}$ for some $i<n$.
+Clearly, we must add all roots from $\bigcup_{n\in\mathbb{N}}R_{n}$ to $\zeta_{D,\gamma}$, and we cannot add any other roots.
 Proposition 2.3.
-(Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$.
-Suppose that any root $\gamma\in D$ satisfy the following condition: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$ then $\alpha\in\zeta_{D,\gamma}$.
+(Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$.
+Suppose that for every root $\gamma\in D$, the following condition holds: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $\alpha\in\zeta_{D,\gamma}$.
 Then $\lambda\in S$.
@@
 Fix a root $\gamma\in D$.
-Consider the orbit of the tangent action $\mathfrak{n}.\lambda$.
-Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$, while $\mu$ is given, and the values of $x_{\alpha}$ are not.
+Consider the orbit of the infinitesimal action $\mathfrak{n}.\lambda$.
+Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, where $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$.
+Here $\mu$ is given, while the values $x_{\alpha}$ are unknown.
 We will show by induction on $|\zeta_{D,\gamma}|$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ can be expressed as a linear combination of $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$.
-Indeed, consider the step of adding $\delta$ to $\zeta_{D,\gamma}$.
-Suppose that we proved this for all roots that are already in $\zeta_{D,\gamma}$.
-There exists a root $\beta\succ\gamma$ such that $\delta$ is the unique root not from $\zeta_{D,\gamma}$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$.
-Consider $\mu(e_{\beta})=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}$.
-Let us rewrite:
+Consider the step of adding $\delta$ to $\zeta_{D,\gamma}$.
+Suppose we have already proved this for all roots that are already in $\zeta_{D,\gamma}$.
+There exists a root $\beta\succ\gamma$ such that $\delta$ is the unique root not in $\zeta_{D,\gamma}$ satisfying $\lambda([e_{\delta},e_{\beta}])\neq 0$.
+Consider $\mu(e_{\beta})=-\sum_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}$.
+This can be rewritten as:
 $\lambda([e_{\delta},e_{\beta}])x_{\delta}=-\mu(e_{\beta})-\sum\limits_{\alpha\neq\delta}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
-We claim that the right hand side is a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
-Indeed, if the coefficient at $x_{\alpha}$ in the right hand side is non-zero (i.e., $\lambda([e_{\alpha},e_{\beta}])\neq 0$), then $\alpha\in\zeta_{D,\gamma}$, hence, by the inductive assumption, $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
-Thus, since $\lambda([e_{\delta},e_{\beta}])\neq 0$, $x_{\delta}$ can also be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
+We claim that the right-hand side is a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
+Indeed, if the coefficient of $x\alpha$ on the right-hand side is non-zero (i.e., $\lambda([e_{\alpha},e_{\beta}])\neq 0$), then $\alpha\in\zeta_{D,\gamma}$.
+By the inductive hypothesis, $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
+Since $\lambda([e_{\delta},e_{\beta}])\neq 0$, it follows that $x_{\delta}$ can also be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
 […]
 Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ will also be satisfied for it.
-So, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none.
-Therefore, it is sufficient to check the conditions for one form with a given support, for example, for the form with coordinates 1 at $e_{\alpha}$, where $\alpha$ lies in the support.
+Therefore, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none do.
+Hence, it suffices to check the conditions for a single form with a given support, for example, for the form with all coordinates equal to 11 at the basis vectors $e_{\alpha}^{*}$ with $\alpha\in D$.
 Proposition 2.4.
@@
 In other words, consider the orbit of the tangent action $\mathfrak{n}.\lambda$, and let $\mu\in\mathfrak{n}.\lambda$.
-It is enough to check that if all $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary.
-We will present $\mu^{\prime}\in\mathfrak{n}.\lambda$ for which $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is a given value.
+It is enough to check that if all values $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary.
+We will construct $\mu^{\prime}\in\mathfrak{n}.\lambda$ such that $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is an arbitrary given value.
 Let $\mu^{\prime}=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$.
-It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$ then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$.
-We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, as the corresponding values for $\mu$.
-We will find the values $x_{\epsilon_{i}}$ using induction on $i=\overline{1,k-1}$.
-Base.
+It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$.
+We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, to be the same as for $\mu$.
+We will determine the values $x_{\epsilon_{i}}$ by induction on $i=\overline{1,k-1}$.
+Base case.
 Let $i=1$, $A=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$.
-Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$ then $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$.
-Hence, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ in the right hand side, except $x_{\epsilon_{1}}$, are already defined.
-Since that $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined, too.
-The only $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that can still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$.
+Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$, then either $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$.
+Hence, in the right-hand side, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ except possibly $x_{\epsilon_{1}}$, are already defined.
+Since $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined as well.
+After this, the only expression $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that may still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$.
 If $k=1$, then nothing else depends on $x_{\epsilon_{1}}$.
-Step.
-Let the value of $x_{\epsilon_{i-1}}$ be found.
+Inductive step.
+Suppose the value $x_{\epsilon_{i-1}}$ has been found.
 Consider $\mu(e_{\delta_{i}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$.
-On the right side, there can only be $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$.
-The value of $x_{\epsilon_{i-1}}$ was determined in the previous step.
-That is, on the right side, only $x_{\epsilon_{i}}$ is undefined, and it is definitely there.
-From this condition, the value of $x_{\epsilon_{1}}$ is uniquely determined.
-The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that still depends on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
+On the right-hand side, only $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$, can appear with non-zero coefficients.
+The value $x_{\epsilon_{i-1}}$ was determined in the previous step.
+Thus, on the right-hand side, only $x_{\epsilon_{i}}$ may be undefined, and it definitely appears.
+From this equation, the value $x_{\epsilon_{1}}$ is uniquely determined.
+The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that may still depend on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
 Case $i=k$.
 Consider $\mu(e_{\delta_{k}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$.
-The right side definitely has $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$.
-The value of $x_{\varepsilon_{k-1}}$ was determined in the previous step (or it is not there at all if $k=1$).
-All other values $x_{\alpha}$ are also defined.
-Therefore, the value of $x_{\epsilon_{k}}$ is uniquely determined.
-Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depend on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
-Thus, all values $x_{\alpha}$ were calculated and the required $\mu^{\prime}$ was presented. $\square$
-Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the coordinates of $\lambda$, but only on its support.
-Therefore, it is enough to check these conditions only for one form with a given support, for example, we may assume that all nonzero coordinates equal 11.
-The conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked using programs.
-Below we present algorithms for checking the necessary and sufficient conditions for a support belonging to the set $S$.
+The right-hand side definitely contains $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$.
+The value $x_{\varepsilon_{k-1}}$ was determined in the previous step (or does not appear if $k=1$).
+All other $x_{\alpha}$ are also defined.
+Therefore, the value $x_{\epsilon_{k}}$ is uniquely determined.
+Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depends on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
+Thus, we have computed all coefficients $x_{\alpha}$ and constructed the required $\mu^{\prime}$. $\square$
+Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the specific coordinates of $\lambda$, but only on its support.
+Therefore, it is sufficient to check these conditions for just one form with a given support; for example, we may assume that all its nonzero coordinates are equal to 11.
+The conditions of Propositions Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked algorithmically.
+Below we present algorithms for checking the necessary and sufficient conditions for a support to belong to the set SS."
 Algorithm
 An incomplete enumeration of supports is performed.
-For each support, the necessary (respectively, sufficient) condition is checked, and then the next support to be checked is built based on it.
-The enumeration is based on the following fact.
-Let $D$ be a support, $\alpha$ a positive root that is not in $D$ and is smaller than all roots from $D$.
-Then if $D$ does not satisfy the necessary (respectively, sufficient) condition, then $D\cup\{\alpha\}$ also does not satisfy the necessary (respectively, sufficient) condition.
-Indeed, the necessary (respectively, sufficient) condition for the support does not hold when there exists a root $\gamma$ in the support that satisfies certain conditions.
-These conditions depend only on $\beta\in D$ such that $\beta\succeq\gamma$.
-Neither the set $\zeta_{D,\gamma}$, nor the values $\lambda([e_{\alpha},e_{\beta}])$, where $\alpha$ or $\beta\succeq\gamma$, depend on the roots in the support that are smaller than $\gamma$.
-The enumeration is performed according to the following rules.
-First, the support $\{2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}\}$ is checked for compliance with the necessary (respectively, sufficient) condition.
-Further, let us assume that we have checked the support $D$ for compliance with the necessary (respectively, sufficient) condition.
+For each support, the necessary (respectively, sufficient) condition is checked, and the next support to be checked is then generated based on the result.
+The enumeration relies on the following observation.
+Let $D$ be a support, and let $\alpha$ be a positive root not in $D$ that is smaller than all roots in $D$.
+If $D$ does not satisfy the necessary (respectively, sufficient) condition, then $D\cup{\alpha}$ also fails to satisfy the necessary (respectively, sufficient) condition.
+Indeed, the necessary (respectively, sufficient) condition for a support fails when there exists a root $\gamma$ in the support satisfying certain conditions.
+These conditions depend only on those $\beta\in D$ such that $\beta\succeq\gamma$.
+Neither the set $\zeta_{D,\gamma}$ nor the values $\lambda([e_{\alpha},e_{\beta}])$ (where $\alpha$ or $\beta\succeq\gamma$) depend on the roots in the support that are smaller than $\gamma$.
+The enumeration proceeds according to the following rules.
+First, the support $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$} is checked against the necessary (respectively, sufficient) condition.
+Next, suppose we have checked a support $D$ against the necessary (respectively, sufficient) condition.
 Let $\delta$ be the minimal root in $D$.
-First, let us consider the case when $\delta\neq\alpha_{4}$.
-If $D$ passed the check, we add to $D$ the positive root closest to $\delta$ from below.
-If $D$ did not pass the check, we remove $\delta$ from $D$ and add the positive root closest to $\delta$ from below.
-Then we check the new support for compliance with the necessary (respectively, sufficient) condition.
-Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if exists) and add the root closest to $\beta$ from below.
+First, let us consider the case if $\delta\neq\alpha_{4}$.
+If $D$ passes the check, we add to $D$ the positive root immediately preceding $\delta$ (with respect to the order $\succ$).
+If $D$ fails the check, we remove $\delta$ from $D$ and add the positive root immediately preceding $\delta$.
+Then we check the new support against the necessary (respectively, sufficient) condition.
+Second, if $\delta=\alpha_{4}$, then we remove $\delta$, remove the new minimal root $\beta$ in the support (if it exists) and add the root closest to $\beta$ from below.
 If $D=\{\alpha_{4}\}$, then the enumeration ends here.
-In the end, we include the empty set in the list of supports, because it definitely satisfies the necessary (respectively, sufficient) condition.
-Thus, we will enumerate all supports that potentially satisfy the necessary (respectively, sufficient) condition, and we will not check supports that definitely do not satisfy these conditions.
-Let us clarify that if $D$ is the support of a form $\lambda$, then the condition $\lambda([e_{\alpha},e_{\beta}])\neq 0$ is equivalent to $\alpha+\beta\in D$.
-Checking this condition can be easily implemented.
-As it was already mentioned, the necessary (respectively, sufficient) condition needs to be checked for all roots of the support $D$.
-It is obvious that it is enough to check it only for the root $\gamma$ that we just added to the support, because for the rest of the roots we checked the condition before.
-The first thing to do is to build the set $\zeta_{D,\gamma}$.
-It can be built as follows.
-We launch a cycle and go through the rows that are numbered by roots $\alpha\succ\gamma$.
-If there is only one non-zero element $\lambda([e_{\alpha},e_{\delta}])$ in the row $\alpha$ such that $\delta\notin\zeta_{D,\gamma}$, then we add $\delta$ to $\zeta_{D,\gamma}$.
-If, after performing the cycle, the number of roots in $\zeta_{D,\gamma}$ has increased, we go through it again.
+Finally, we include the empty set in the list of supports, since it trivially satisfies the necessary (respectively, sufficient) condition.
+In this way, we enumerate all supports that could potentially satisfy the necessary (respectively, sufficient) condition, while avoiding checks for supports that certainly do not satisfy these conditions.
+Note that if $D$ is the support of a form $\lambda$, then the condition $\lambda([e_{\alpha},e_{\beta}])\neq 0$ is equivalent to $\alpha+\beta\in D$.
+Checking this condition is easy to implement.
+As mentioned earlier, the necessary (respectively, sufficient) condition must be checked for all roots in the support $D$.
+It is clear that it suffices to check it only for the root $\gamma$ that was most recently added to the support, because for all other roots the condition has already been verified.
+The first step is to construct the set $\zeta_{D,\gamma}$.
+This can be done as follows.
+We run a loop over rows indexed by roots $\alpha\succ\gamma$.
+If in a row $\alpha$ there is exactly one non-zero element $\lambda([e_{\alpha},e_{\delta}])$ such that $\delta\notin\zeta_{D,\gamma}$, then we add $\delta$ to $\zeta_{D,\gamma}$.
+After completing a full pass through the rows, if the size of $\zeta_{D,\gamma}$ has increased, we repeat the loop.
 The process stops when the cardinality of $\zeta_{D,\gamma}$ stabilizes.
-Thus, we have built $\zeta_{D,\gamma}$.
-Checking the sufficient condition for the support $D$ is trivial.
-We go through the row numbered $\gamma$ and check the condition of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+This yields the set $\zeta_{D,\gamma}$.
+Checking the sufficient condition for the support $D$ is straightforward.
+We iterate over the row indexed by $\gamma$ and verify the condition from Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
 The necessary condition is checked as follows.
-Let us describe the $i$-th step of the check.
-We go through the row numbered $\delta_{i}$ (recall that $\delta_{1}=\gamma$).
-If we meet a non-zero element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$ and $\varepsilon\neq\varepsilon_{i-1}$, we check the column numbered $\varepsilon$, starting from $\gamma$ and higher.
-If there are no non-zero elements in this part of the column, except for $\lambda([e_{\delta_{i}},e_{\varepsilon}])$, then the necessary condition is not satisfied.
-If this is not the case, but only one non-zero element was found in the row (possibly except for $\lambda([e_{\delta_{i}},e_{\varepsilon_{i-1}}])$) of the form $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$, and at the same time, only one non-zero element was found in the corresponding part of the column, except for ours, of the form $\lambda([e_{\delta},e_{\varepsilon}])$, then $\varepsilon=\varepsilon_{i}$, and $\delta=\delta_{i+1}$.
-And then we check the condition already for the column numbered $\delta_{i+1}$.
-In fact, our program allows checking the necessary condition only up to the second step.
-And if the necessary condition is still not violated by that moment, it considers that the check has passed.
-But this hardly complicates the final calculation.
-The results of the check are as follows: 878 supports satisfy the sufficient condition, while 911 satisfy the necessary condition, so 33 supports remain to be checked manually. Here they are:
+We describe the $i$-th step of the verification.
+We examine the row indexed by $\delta_{i}$ (recall that $\delta_{1}=\gamma$).
+If we encounter a non-zero element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ such that $\varepsilon\notin\zeta_{D,\gamma}$ and $\varepsilon\neq\varepsilon_{i-1}$ (where $\varepsilon_{0}$ is undefined), we inspect the column indexed by $\varepsilon$, starting from row $\gamma$ upwards.
+If there are no other non-zero elements in this segment of the column, the necessary condition is not satisfied.
+If this is not the case, but exactly one such element $\lambda([e_{\delta_{i}},e_{\varepsilon}])$ with $\varepsilon\notin\zeta_{D,\gamma}$ is found in the row (ignoring a possible $\lambda([e_{\delta_{i}},e_{\varepsilon_{i-1}}])$), and if, in the corresponding segment of the column, there is exactly one other non-zero element of the form $\lambda([e_{\delta},e_{\varepsilon}])$, then we set $\varepsilon_{i}=\varepsilon$ and $\delta_{i+1}=\delta$.
+We then proceed to check the condition for the next column, indexed by $\delta_{i+1}$.
+In practice, our program checks the necessary condition only up to the second step.
+If the condition has not been violated by that point, the program considers the check passed.
+This simplification does not significantly affect the final result.
+The results of the check are as follows: 878 supports satisfy the sufficient condition, while 911 satisfy the necessary condition.
+Therefore, 33 supports remain to be checked manually.
+They are listed below:
   1. 1.
@@
     $\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{3}+\alpha_{4},~\alpha_{4}$
-Linear forms with support 5 actually don’t satisfy the necessary condition, but checking this requires three steps, while the algorithm checks only two steps, see above.
-It is clear that linear forms with the supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1), i.e., such forms do not belong to $S$.
-Indeed, it follows immediately from Table 1 as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-In other cases, it is necessary to explicitly write out the coordinates of the tangent action.
-It turns out that linear forms with the supports numbered 25 and 33 lie in $S$ without restrictions on the coordinates.
-Linear forms with supports numbered 30, 31 and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
-We obtain 883 supports whose forms belong to $S$, 3 of which require a restriction on the coordinates of forms.
-Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved.
+Linear forms with support 5 do not satisfy the necessary condition, but checking this requires three steps, whereas our algorithm checks only the first two steps (see above).
+It is clear that linear forms with supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1); that is, such forms do not belong to $S$.
+This follows immediately from Table 1, similar to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+In the remaining cases, it is necessary to write out the coordinates of the infinitesimal action explicitly.
+It turns out that the linear forms with supports numbered 25 and 33 lie in $S$ without any restrictions on their coordinates.
+Linear forms with supports numbered 30, 31, and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+Thus, we obtain 883 supports whose corresponding forms belong to $S$, three of which require an additional condition on the coordinates.
+This completes the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
 Example 2.5.
@@
 We will show that $\lambda\in S$.
-Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}$.
-Let us explicitly write down the coordinates we are interested in.
-We will denote the structural constants in the expression $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
+Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}^{*}$.
+Let us write down explicitly the coordinates of interest.
+We denote the structure constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
 Their values can be found in [KP23].
@@
 $\begin{split}\mu_{\alpha_{1}+\alpha_{2}}&=\mu_{\alpha_{2}+2\alpha_{3}}=\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=0;\\ \mu_{\alpha_{1}}&=-\lambda([x,e_{\alpha_{1}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{1}}])=-C_{\alpha_{2},\alpha_{1}}x_{\alpha_{2}}\lambda(e_{\alpha_{1}+\alpha_{2}})=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ &=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\lambda_{\alpha_{1}+\alpha_{2}}}\mu_{\alpha_{1}};\\ \mu_{\alpha_{3}}&=-\lambda([x,e_{\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{3}+\alpha_{4}}-2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{4}}&=-\lambda([x,e_{\alpha_{4}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])\\ &=-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\\ &=\dfrac{\lambda_{\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{2}+\alpha_{3}}-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{3}}.\end{split}$
-Let us also consider the support numbered 25.
-All roots, except $\alpha_{2}$, obviously satisfy condition (1).
-The coordinate $\mu_{\alpha_{2}}$ is equal to $x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}}$, so if the polynomial $x_{\alpha_{3}}$ can be expressed as a linear combination of $\{\mu_{\beta}\}_{\beta\succ\alpha_{2}}$ then $\alpha_{2}$ satisfies condition (1).
-That is exactly the case:
+Let us also consider support number 25.
+For all roots except $\alpha_{2}$, condition (1) obviously holds.
+The coordinate $\mu_{\alpha_{2}}$ equals $x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}}$.
+Therefore, if $x_{\alpha_{3}}$ can be expressed as a linear combination of $\{\mu_{\beta}\}_{\beta\succ\alpha_{2}}$, then $\alpha_{2}$ satisfies condition (1).
+This is indeed the case:
 $\begin{split}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+2\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}};\\ x_{\alpha_{3}}&=\dfrac{\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}-\mu_{\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}+\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{3\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
-Finally, let us consider forms with the supports 30, 31 and 32.
-Let the support contain roots $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$.
+Finally, consider forms with supports 30, 31, and 32.
+Suppose the support contains the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4}$.
 Let us check whether condition (1) holds for the root $\alpha_{2}$.
-Let us note that
+Note that
 $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+x_{\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{2}}&=-x_{\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}}.\end{split}$
-We see that if $\beta\succ\alpha_{2}$ and $x_{\alpha_{3}+\alpha_{4}}$ enters $\mu_{\beta}$ then $\beta=\alpha_{2}+\alpha_{3}$; similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ enters $\mu_{\beta}$ then $\beta=\alpha_{2}+2\alpha_{3}$.
-Hence, the coordinate $\mu_{\alpha_{2}}$ should be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$.
-It is easy to check that this is true if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
-Let us check whether condition (1) holds for the root $\alpha_{4}$. Note that
+We see that if $\beta\succ\alpha_{2}$ and the variable $x_{\alpha_{3}+\alpha_{4}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+\alpha_{3}$.
+Similarly, if $\beta\succ\alpha_{2}$ and $x_{\alpha_{1}}$ appears in $\mu_{\beta}$, then $\beta=\alpha_{2}+2\alpha_{3}$.
+Hence, the coordinate $\mu_{\alpha_{2}}$ must be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$.
+It is easy to verify that this holds if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+Now, let us check condition (1) for the root $\alpha_{4}$. Note that
 $\begin{split}\mu_{\alpha_{1}}&=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}}-x_{\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}.\end{split}$
-Arguing as above we obtain that the coordinate $\mu_{\alpha_{4}}$ should be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$.
-It turns out that this is true if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+By an argument analogous to the one above, we find that the coordinate $\mu_{\alpha_{4}}$ must be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$.
+It turns out that this condition is also equivalent to $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
 3. Classification of orbits
-In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, which will complete a classification of coadjoint orbits for type $F_{4}$.
-For the next stage, we will need a concept of a singular root.
+In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, thereby completing the classification of coadjoint orbits for type $F_{4}$.
+For the next stage, we will need the concept of a singular root.
 Definition 3.1. Let $\alpha,\gamma\in\Phi^{+}$. The root $\alpha$ is called singular to $\gamma$ if $\gamma-\alpha\in\Phi^{+}$. The set of all roots singular to $\gamma$ is denoted by $S_{\gamma}$.
-It is obvious that the singular roots from $S_{\gamma}$ are divided into pairs of the form $\alpha+\beta=\gamma$.
-Next, unless stated otherwise, everything will take place over the closure of the field $\overline{\mathbb{F}}$. We introduce two $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$-modules:
+It is obvious that the singular roots in $S_{\gamma}$ are divided into pairs of the form $\alpha+\beta=\gamma$.
+From now on, unless stated otherwise, everything will take place over the algebraic closure $\overline{\mathbb{F}}$. We introduce two $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$-modules:
 $\begin{split}M&=\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]\otimes_{\mathbb{F}}\mathfrak{n},\\ M^{*}&=\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]\otimes_{\mathbb{F}}\mathfrak{n}^{*}.\end{split}$
-Here $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ denotes the ring of polynomials over $\overline{\mathbb{F}}$ in independent variables in $x_{\alpha},\alpha\in\Phi^{+}$.
-Next, we define the pairing $M^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ by putting
+Here $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ denotes the ring of polynomials over $\overline{\mathbb{F}}$ in the independent variables $x_{\alpha},\alpha\in\Phi^{+}$.
+We define a pairing $M^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ by
 $\langle f\otimes\lambda,g\otimes x\rangle\mapsto\lambda(x)fg.$
-Everywhere below we will write $\lambda(x)$ instead of $\langle\lambda,x\rangle$ for $\lambda\in M^{*},x\in M$.
-Notice that $\mathfrak{n}$ is a $\mathbb{F}$-subspace in $M$, and $\mathfrak{n}^{*}$ is a $\mathbb{F}$-subspace in $M^{*}$.
-Therefore, two other pairings $\mathfrak{n}^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $M^{*}\times\mathfrak{n}\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ are automatically defined.
-We introduce a Lie algebra structure on $M$ in an obvious way: for $f,g\in\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $x,y\in\mathfrak{n}$, then
+In what follows, we will write $\lambda(x)$ instead of $\langle\lambda,x\rangle$ for $\lambda\in M^{*},x\in M$.
+Note that $\mathfrak{n}$ is a $\mathbb{F}$-subspace of $M$, and $\mathfrak{n}^{*}$ is a $\mathbb{F}$-subspace of $M^{*}$.
+Therefore, two other pairings, $\mathfrak{n}^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $M^{*}\times\mathfrak{n}\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$, are automatically induced.
+We introduce a Lie algebra structure on $M$ in the obvious way: for $f,g\in\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $x,y\in\mathfrak{n}$, set
 $f\otimes x,g\otimes y]=fg\otimes[x,y].$
-Since $\mathfrak{n}$ can be identified with upper triangular matrices with coefficients from $\mathbb{F}$, we can identify $M$ with upper triangular matrices with coefficients from $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$.
-Therefore, we can define $\exp(x)$ for $x\in M$.
-Since the characteristic of $\mathbb{F}$ is sufficiently large, $\exp(x)$ is well defined for all $x\in M$.
+Since $\mathfrak{n}$ can be identified with upper-triangular matrices with entries in $\mathbb{F}$, we can identify $M$ with upper-triangular matrices whose entries lie in $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$.
+Consequently, we can define $\exp(x)$ for $x\in M$.
+Since the characteristic of $\mathbb{F}$ is sufficiently large, $\exp(x)$ is well-defined for all $x\in M$.
 Let $N_{M}$ be the group consisting of the matrices $\{\exp(x),x\in M\}$.
-The actions $N_{M}\curvearrowright M$ and $N_{M}\curvearrowright M^{*}$ are naturally defined.
-The explicit description of the action $N_{M}\curvearrowright M^{*}$ is as follows: if $x\in M,\lambda\in M^{*}$, then
+The actions $N_{M}\curvearrowright M$ and $N_{M}\curvearrowright M^{*}$ are defined naturally.
+An explicit description of the action $N_{M}\curvearrowright M^{*}$ is as follows: for $x\in M$ and $\lambda\in M^{*}$,
 $\exp(x).\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda\left(\sum\limits_{i=0}^{\infty}\dfrac{(-1)^{i}}{i!}\underbrace{[x,[x,\dots,[x,e_{\alpha}]\dots]]}_{\text{$i$ commutators}}\right)\otimes e_{\alpha}^{*}.$
-Denote the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$.
-For forms $\lambda\in\mathfrak{n}^{*}$, we will consider the elements $\exp(\chi).\lambda\in M^{*}$.
-Let us designate $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$.
-Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ for certain $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
-Now we formulate one of the most important theorems of this work. Let $y_{\beta},\beta\in\Phi^{+}$, be independent variables.
+Denote by $\chi$ the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$.
+For a form $\lambda\in\mathfrak{n}^{*}$, we will consider the element $\exp(\chi).\lambda\in M^{*}$.
+Let us also set $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$.
+Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ with $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
+Now we formulate one of the most important theorems of this work. Let $y_{\beta},\beta\in\Phi^{+}$ be independent variables.
 Theorem 3.2.
-Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, $\gamma\in\Phi^{+}$ satisfies condition (1).
+Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1).
 Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
@@
 Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
-We will show that for $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds.
-Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}$.
-Let us explicitly write down the coordinates we are interested in.
-As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we will denote the structural constants in the expression $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
+We will show that for this $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds.
+Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}^{*}$.
+Let us write down explicitly the coordinates of interest.
+As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we denote the structural constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
 Their values can be found in [KP23].
@@
 $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=0;\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}}&=\lambda(e_{\alpha_{2}})+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{2}}]])=\lambda_{\alpha_{2}}+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}}]])\\ &=\lambda_{\alpha_{2}}+\dfrac{1}{2}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}C_{\alpha_{3},\alpha_{2}}x_{\alpha_{3}}^{2}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}^{2}}{4\mu_{\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{3}+\alpha_{4}}])\\ &=\lambda_{\alpha_{4}}-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
-In fact, the validity of formula (2) in this example could have been understood without explicitly writing out the coordinates.
-It is enough to look at the Table 1.
-Due to the fact that $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, $x_{\alpha_{3}}$ can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
-Since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
-Further, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$.
-Since $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
-Now let us move on to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-First, we will prove it in a particular case, for forms with a support consisting of one root.
-Obviously, any such a form lies in $S$.
+In fact, the validity of formula (2) in this example could have been understood without writing out the coordinates explicitly.
+It is enough to examine Table 1.
+Since $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, the variable $x_{\alpha_{3}}$ can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
+Consequently, since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ depends only on $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
+Furthermore, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$.
+As $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
+Now let us proceed to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+First, we prove it in the special case of forms whose support consists of a single root.
+Any such form obviously belongs to $S$.
 Lemma 3.4.
@@
 Denote $D=\textup{supp}(\lambda)=\{\delta\}$.
-Let us note that in this case, the equality (1) only depends on $D$ and not on $\lambda$, and holds for those $\gamma$ that are not singular to $\delta$.
+Note that in this case, condition (1) depends only on $D$, not on the specific $\lambda$, and it holds for those $\gamma$ that are not singular to $\delta$.
 Let $\mu=\exp(\chi).\lambda$.
 We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}.
-Note that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$, so for these cases, the condition is obviously satisfied.
+Observe that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$.
+For these two cases, the condition is obviously satisfied.
 It remains to prove it for $\gamma\prec\delta$.
@@
 $\begin{split}\mu(e_{\beta})=&-\lambda([x,e_{\beta}])+\dfrac{1}{2!}\lambda([x,[x,e_{\beta}]])-\dfrac{1}{3!}\lambda([x,[x,[x,e_{\beta}]]])+\ldots\\ =&-\lambda([x_{\delta-\beta}e_{\delta-\beta},e_{\beta}])+F(x)=-\lambda([e_{\delta-\beta},e_{\beta}])x_{\delta-\beta}+F(x)=C\lambda(e_{\delta})x_{\delta-\beta}+F(x),\end{split}$
-where $C\neq 0$ and $F(x)$ is a polynomial in $x_{\alpha}$, where $\alpha<\delta-\beta$. Thus,
+where $C\neq 0$ and $F(x)$ is a polynomial in variables $x_{\alpha}$ with $\alpha<\delta-\beta$. Consequently,
 $x_{\delta-\beta}=\dfrac{\mu(e_{\beta})-F(x)}{C\lambda(e_{\delta})}$ (3)
 Let $\beta_{1},\beta_{2},\ldots,\beta_{2k}$ be the singular roots to $\delta$, listed in decreasing order with respect to lexicographical order.
-They are paired such that $\beta_{i}+\beta_{2k-i+1}=\delta$.
-Let $\gamma\notin S_{\delta}$ with $\delta\succ\gamma$.
-Consider $\mu(e_{\gamma})$.
-Suppose that there is no rational function $H$ over $\overline{\mathbb{F}}$ of independent variables $t_{\beta},\beta\succ\gamma$ such that $\mu(e_{\gamma})=H(\mu(e_{\beta}),\beta\succ\gamma)$.
-The coordinate $\mu(e_{\gamma})$ is a polynomial in $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$.
-It does not contain $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$.
-Let $\beta_{2k-j+1}$ be the largest root from $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ is in $\mu(e_{\gamma})$.
-We express $x_{\beta_{2k-j+1}}$ using formula (3) through $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}.
-Thus, now $\mu(e_{\gamma})$ is a polynomial that does not include $x_{\beta_{2k-j+1}}$.
-We perform a similar operation for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$.
-We note that at each step, $x_{\beta_{i}}$, which was initially absent and has already been removed, cannot reappear, as we are proceeding in decreasing order of singular roots.
-As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}(t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$.
-Note that all $\beta_{i}\succ\gamma,i=1,\ldots,j$.
-According to our assumption, $\mu(e_{\gamma})$ cannot be expressed in terms of the subsequent coordinates, which means that this polynomial is not identically equal to some number and contains at least one variable $x_{\alpha}$.
-Then, there exists an open Zariski subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is not a constant.
-Fix any such set $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$.
-Choose $\varepsilon\notin S_{\delta}$ and $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$ such that the polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is not constant.
-The field is algebraically closed, therefore for any $c_{0}\in\overline{\mathbb{F}}$, there exists such $c_{\varepsilon}\in\overline{\mathbb{F}}$ that $g(c_{\varepsilon})=c_{0}$.
-We fix $c_{0}$ and $c_{\varepsilon}$.
-We can choose any $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$.
-By substituting $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into the formula (3), we will sequentially assign values to $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ so that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$.
-For these values, $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
-Thus, there exists an open Zariski subset $V\subseteq\overline{\mathbb{F}}^{2k}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ there exist $\overline{\mu}\in\overline{N}.\lambda$ for which the coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values.
+They come in pairs such that $\beta_{i}+\beta_{2k-i+1}=\delta$.
+Now take $\gamma\notin S_{\delta}$ with $\delta\succ\gamma$ and consider $\mu(e_{\gamma})$.
+Suppose, for contradiction, that there is no rational function $H$ over $\overline{\mathbb{F}}$ in the independent variables $t_{\beta},\beta\succ\gamma$ such that $\mu(e_{\gamma})=H(\mu(e_{\beta}),\beta\succ\gamma)$.
+The coordinate $\mu(e_{\gamma})$ is a polynomial in the $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$.
+It does not contain the variable $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$.
+Let $\beta_{2k-j+1}$ be the largest root in $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ appears in $\mu(e_{\gamma})$.
+Using formula (3), we express $x_{\beta_{2k-j+1}}$ in terms of $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}.
+Thus, $\mu(e_{\gamma})$ becomes a polynomial that no longer involves $x_{\beta_{2k-j+1}}$.
+We perform a similar elimination successively for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$.
+Note that at each step, a variable $x_{\beta_{i}}$ that was initially absent or already removed cannot reappear, because we proceed in decreasing order of the singular roots.
+As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}[t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta}]$.
+Observe that $\beta_{i}\succ\gamma$ for all $i=1,\ldots,j$.
+By our assumption, $\mu(e_{\gamma})$ cannot be expressed solely in terms of the subsequent coordinates.
+This implies that $G$ is not identically constant and actually depends on at least one variable $x_{\alpha}$.
+Then there exists a non-empty Zariski-open subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the specialized polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is non-constant.
+Fix such a tuple $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$.
+Choose an index $\varepsilon\notin S_{\delta}$ and constants $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$, so that the one-variable polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is non-constant.
+The field is algebraically closed, for any chosen $c_{0}\in\overline{\mathbb{F}}$ there exists a $c_{\varepsilon}\in\overline{\mathbb{F}}$ such that $g(c_{\varepsilon})=c_{0}$.
+Fix $c_{0}$ and the corresponding $c_{\varepsilon}$.
+Now, choose arbitrary values $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$.
+Substituting the values $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into formula (3), we can determine values for $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ sequentially, in such a way that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$.
+For this assignment, we obtain $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
+Therefore, there exists a non-empty Zariski-open subset $V\subseteq\overline{\mathbb{F}}^{2k+1}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ , we can find an element $\overline{\mu}\in\overline{N}.\lambda$ whose coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values.
 The orbit $\overline{N}.\lambda$ is an algebraic variety.
-From our construction, it follows that there is a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, therefore $\dim\overline{N}.\lambda>2k$.
-But $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$.
-A variety of dimension $2k$ cannot contain $\overline{\mathbb{F}}^{2k+1}$.
-This is a contradiction.
+Our construction yields a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, which implies $\dim\overline{N}.\lambda\geq 2k+1$.
+However, $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$.
+This contradiction completes the proof.
 $\square$
-We introduce more notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
+We introduce further notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
 $T_{\alpha,\beta}(x):=(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\in\overline{\mathbb{F}}[x_{\gamma},\gamma\in\Phi^{+}]$. It is clear that if $\lambda\in\mathfrak{n}^{*},\alpha\in\Phi^{+}$, then
@@
 Let $\lambda\in S,~D=\textup{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1).
-Starting from $D$ and $\gamma$, we will construct the set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
+Starting from $D$ and $\gamma$, we construct a set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition, it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
 Example 3.5.
-Table 1 helps to visualize how the polynomials $T_{\alpha,\beta}(x)$ look like.
-It is enough to look at the cell in the row numbered $\alpha$ and the column numbered $\beta$.
-For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, we denote the non-zero constant $C_{\gamma,\delta}$ by letting $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$.
+Table 1 helps visualize the structure of the polynomials $T_{\alpha,\beta}(x)$.
+One simply examines the cell in row $\alpha$ and column $\beta$.
+For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, denote the corresponding non-zero structure constant by $C_{\gamma,\delta}$, so that $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$.
 Each monomial in $T_{\alpha,\beta}(x)$ has the form
 $(-1)^{n+1}\dfrac{1}{n!}C_{\gamma_{1},\alpha}C_{\gamma_{2},\alpha+\gamma_{1}}\ldots C_{\gamma_{n},\alpha+\gamma_{1}+\ldots+\gamma_{n-1}}x_{\gamma_{1}}x_{\gamma_{2}}\ldots x_{\gamma_{n}},$
-where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 1 to $n$, $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$.
-Hence, it is enough to subsequently consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots.
-For instance, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
-Note that if $\alpha\nleq\beta$ then $T_{\alpha,\beta}(x)=0$, while $T_{\alpha,\alpha}(x)=1$.
-In all other cases, $T_{\alpha,\beta}(x)$ has non-zero degree.
-Let us call a pair of roots $(\alpha,\beta)$ nontrivial if $\alpha<\beta$ (i.e., $T_{\alpha,\beta}(x)\neq const$).
+where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 11 to $n$, and $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$.
+Hence, it suffices to consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots.
+For example, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
+Note that if $\alpha\nleq\beta$, then $T_{\alpha,\beta}(x)=0$, while $T_{\alpha,\alpha}(x)=1$.
+In all other cases, $T_{\alpha,\beta}(x)$ has positive degree.
+We call a pair $(\alpha,\beta)$ nontrivial if $\alpha<\beta$ (i.e., $T_{\alpha,\beta}(x)\neq const$).
 Obviously, each trivial pair belongs to all $\zeta^{\prime}_{D,\gamma}$.
-Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ again.
+Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
 There $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
-The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ do not contain nontrivial pairs.
+The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ contain no nontrivial pairs.
 Indeed, there are no $\beta$ in $D$ such that $\beta>\alpha_{2}+2\alpha_{3}$ or $\beta>\alpha_{2}+\alpha_{3}+\alpha_{4}$.
@@
 $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})=2x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}(\exp(\chi).\lambda)(e_{\alpha_{2}+2\alpha_{3}})$
-and $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)=-x_{\alpha_{3}}^{2}$, the pair $(\alpha_{2},\alpha_{2}+2\alpha_{3})$ enters $\zeta^{\prime}_{D,\alpha_{2}}$.
-This pair also belongs to $\zeta^{\prime}_{D,\alpha_{3}+\alpha_{4}}$ and $\zeta^{\prime}_{D,\alpha_{4}}$, but $\zeta^{\prime}_{D,\alpha_{4}}$ contains one more nontrivial pair, namely, $(\alpha_{4},\alpha_{3}+\alpha_{4})$, because $T_{\alpha_{4},\alpha_{3}+\alpha_{4}}(x)=x_{\alpha_{3}}$.
-The remaining pairs of the form $(\gamma,\alpha)$, where $\alpha\in D$ and $\gamma$ satisfies condition (1), are trivial, therefore, they lie in all $\zeta^{\prime}_{D,\gamma}$.
-The lemma below highlights why we consider exactly these pairs.
+and $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)=-x_{\alpha_{3}}^{2}$, the pair $(\alpha_{2},\alpha_{2}+2\alpha_{3})$ belongs to $\zeta^{\prime}_{D,\alpha_{2}}$.
+This pair also lies in $\zeta^{\prime}_{D,\alpha_{3}+\alpha_{4}}$ and $\zeta^{\prime}_{D,\alpha_{4}}$.
+Moreover, $\zeta^{\prime}_{D,\alpha_{4}}$ contains one additional nontrivial pair, namely, $(\alpha_{4},\alpha_{3}+\alpha_{4})$, because $T_{\alpha_{4},\alpha_{3}+\alpha_{4}}(x)=x_{\alpha_{3}}$.
+The remaining pairs of the form $(\gamma,\alpha)$ with $\alpha\in D$ and $\gamma$ satisfies condition (1) are trivial and therefore belong to all $\zeta^{\prime}_{D,\gamma}$.
+The lemma below explains why we focus precisely on these pairs.
 Lemma 3.6.
-Let $\lambda\in S,D=\textup{supp}(\lambda)$, $\gamma\in\Phi^{+}$ satisfy condition (1).
-Let for any $\alpha\in D$ the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$ and formula (2) is true for all $\gamma^{\prime}\succ\gamma$ satisfying condition (1).
-Then formula (2) is also true for $\lambda$, $D$ and $\gamma$.
+Let $\lambda\in S,D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1).
+Suppose that for every $\alpha\in D$, the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$, and formula (2) holds for all $\gamma^{\prime}\succ\gamma$ that satisfy condition (1).
+Then formula (2) also holds for $\lambda$, $D$ and $\gamma$.
 Proof. Consider
 $(\exp(\chi).\lambda)(e_{\gamma})=\sum\limits_{\alpha\in D}\lambda(e_{\alpha})T_{\gamma,\alpha}(x)=\lambda(e_{\gamma})+\sum\limits_{\alpha\in D,\alpha\succ\gamma}\lambda(e_{\alpha})T_{\gamma,\alpha}(x).$
-Let $\alpha\in D$, $\alpha\succ\gamma$.
-We have $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, hence $T_{\gamma,\alpha}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$.
-It follows from the construction of $S$ that $\alpha$ satisfies condition (1).
-Hence, by our assumption, $\lambda(e_{\alpha})$ can be also expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
+Let $\alpha\in D$ with $\alpha\succ\gamma$.
+Since $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$.
+By the definition of $S$, the root $\alpha$ satisfies condition (1).
+Therefore, by the inductive hypothesis, $\lambda(e\alpha)$ can also be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
 $\lambda(e_{\alpha})=(\exp(\chi).\lambda)(e_{\alpha})-F_{D,\alpha}((\exp(\chi).\lambda)(e_{\beta}),\beta\succ\alpha).$
-Finally, $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, so condition (2) is satisfied, as required.
+Consequently, the difference $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ is a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, which means condition (2) is satisfied.
 $\square$
-Finally, we are now ready for the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+We are now ready to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
 Proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-If $D=\{\alpha\}$ then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-In fact, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$.
-Furthermore, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$.
-Finally, for $\gamma\prec\alpha$ satisfying condition (1), we see that $T_{\gamma,\alpha}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
-Now, we will prove formula (2) step by step for all $\lambda\in S$ by subsequent consideration of supports from the list.
-(Note that most of the supports can be considered similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
-  • Let the largest (with respect to the lexicographic order) root in the support be less than $\alpha_{2}+\alpha_{3}+\alpha_{4}$.
-In these case, Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for all supports and all $\gamma\in\Phi^{+}$ satisfying condition (1), so the prove is complete.
-(In fact, $T_{\gamma,\alpha}(x)$ for all $\gamma\in\Phi^{+}$ satisfying condition (1) are equal up to multiplication by a certain nonzero scalar to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0 or 1.)
-  • Let the largest root in the support be equal to $\alpha_{2}+\alpha_{3}+\alpha_{4}$.
+If $D=\{\alpha\}$, then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+Indeed, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$.
+Moreover, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$.
+Finally, for $\gamma\prec\alpha$ satisfying condition (1), the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
+Now we will prove formula (2) step by step for all $\lambda\in S$ by considering the supports from the list one after another.
+(Note that most supports can be treated similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$; see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
+  • Suppose the largest root (with respect to the lexicographic order) in the support is smaller than $\alpha_{2}+\alpha_{3}+\alpha_{4}$.
+For all such supports and for every $\gamma\in\Phi^{+}$ satisfying condition (1), Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, so the proof is complete.
+(Indeed, for all relevant $\gamma$, the polynomials $T_{\gamma,\alpha}(x)$ are, up to a non-zero scalar factor, equal to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0, or 11.)
+  • Suppose the largest root in the support equals $\alpha_{2}+\alpha_{3}+\alpha_{4}$.
 There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$.
-Due to the fact that the condition of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is satisfied for supports consisting of one root, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ or not (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$).
-The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, hence there is noting to prove in this case.
-  • Next, let the largest root in the support be equal to $\alpha_{2}+2\alpha_{3}$.
-Here, the first non-trivial case arises, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ can not be proved for all such supports just by subsistent applying of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-The support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$.
-This happened due to the fact that in the table cells at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ the same root $\alpha_{2}+\alpha_{3}$ is written.
+Since Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ covers the case of one-element supports, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$).
+The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, so nothing remains to prove in this case.
+  • Next, suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}$.
+Here the first nontrivial case appears, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be proved for all supports of this type by a straightforward application of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+Such a support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$.
+The complication arises because the table entries at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ both involve the same root $\alpha_{2}+\alpha_{3}$.
 However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$.
-Hence, the following situation can occur: $\alpha_{4}$ satisfies condition (1), but we can not apply Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-Therefore, here we have to compute the coordinates of $\mu$ explicitly:
+Consequently, a situation may occur where $\alpha_{4}$ satisfies condition (1) but Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be applied directly.
+Therefore, in this case we must compute the coordinates of $\mu$ explicitly:
     $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}=&~\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=&~\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}}=&-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-\\ &-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{3}+\alpha_{4}}=&-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&-x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{3}}=&-\lambda([x,e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{3}}]])=-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{2}}e_{\alpha_{2}},[x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{4}}e_{\alpha_{4}},e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x_{\alpha_{4}}e_{\alpha_{4}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}}])\\ =&-C_{\alpha_{2}+\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\dfrac{1}{2}C_{\alpha_{2},\alpha_{3}}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\\ &+\dfrac{1}{2}(C_{\alpha_{4},\alpha_{3}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}+C_{\alpha_{2},\alpha_{3}}C_{\alpha_{4},\alpha_{2}+\alpha_{3}})x_{\alpha_{2}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&~2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}+x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}-x_{\alpha_{2}}x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{4}}=&~\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{4}}])+\frac{1}{2}\lambda([x,[x,e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})+\\ &+\dfrac{1}{2}C_{\alpha_{3},\alpha_{4}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\\ &+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\mu_{\alpha_{3}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
-    This shows that formula (2) holds for root $\alpha_{4}$.
-  • Now, let the largest root in the support be equal either to $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or to $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$.
-In these cases, everything is also almost trivial due to Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-The only non-trivial case is when the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$.
-At first glance, it seems that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$.
-In fact, this is not the case.
-From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ is expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
-Therefore, $x_{\alpha_{3}}$ can be found from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$, thus $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ contains only $x_{\alpha_{3}}$.
-  • If the largest root in the support is equal to $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the non-trivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$.
-This case follows immediately from the formulas in the Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
-  • If the largest root in the support is equal to $\alpha_{1}+\alpha_{2}+\alpha_{3}$, then the all non-trivial cases can be reduced to the case $D=\{\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4},\alpha_{2}\}$.
-Let us write down the coordinates we are interested in:
+    This shows that formula (2) holds for the root $\alpha_{4}$.
+  • Now suppose the largest root in the support equals either $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$.
+In these cases, the statement also follows almost trivially from Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+The only nontrivial situation occurs for the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$.
+At first glance, it might seem that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$.
+However, this is not the case.
+From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+Consequently, $x_{\alpha_{3}}$ can be recovered from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$.
+Since $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ depends only on $x_{\alpha_{3}}$.
+  • If the largest root in the support equals $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the nontrivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$.
+This case follows immediately from the formulas in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ together with the analysis of the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
+  • If the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}$, then all nontrivial cases reduce to $D=\{\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4},\alpha_{2}\}$.
+Let us write down the relevant coordinates:
     $\begin{split}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}&=\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}},\\ \mu_{\alpha_{1}+\alpha_{2}}&=-x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}},\\ \mu_{\alpha_{1}}&=-x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}},\\ \mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}&=\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+2\alpha_{3}}&=-x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=-\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}},\\ \mu_{\alpha_{2}+\alpha_{3}}&=x_{\alpha_{1}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}}+x_{\alpha_{3}+\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}-\dfrac{3}{2}x_{\alpha_{3}}x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}}&=\lambda_{\alpha_{2}}-x_{\alpha_{1}}x_{\alpha_{3}}\lambda_{\alpha_{1}+\alpha_{2}+\alpha_{3}}-x_{\alpha_{3}+\alpha_{4}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{3}}^{2}x_{\alpha_{4}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}\\ &=\lambda_{\alpha_{2}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}}+\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}}^{2}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}^{2}},\\ \mu_{\alpha_{3}+\alpha_{4}}&=-x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\dfrac{\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}}.\end{split}$
-    We conclude that for roots $\alpha_{2}$ and $\alpha_{3}+\alpha_{4}$ formula (2) holds.
-  • Next, if the largest root in the support is equal to $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those with roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$.
+    We conclude that for the roots $\alpha_{2}$ and $\alpha_{3}+\alpha_{4}$, formula (2) holds.
+  • Next, if the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those containing the roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$.
 One can easily show that
     $\mu_{\alpha_{2}}=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{1}+\alpha_{2}}^{2}\mu_{\alpha_{2}+2\alpha_{3}+2\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}.$
-    First, consider the case when $\lambda_{\alpha_{2}+2\alpha_{3}}=0$.
-Clearly, the root $\alpha_{2}$ satisfies formula (2).
-Since $\alpha_{2}+2\alpha_{3}\notin D$, $D$ may contain either $\alpha_{3}$ or $\alpha_{2}+\alpha_{3}$.
+    First, consider the case $\lambda_{\alpha_{2}+2\alpha_{3}}=0$.
+Then clearly the root $\alpha_{2}$ satisfies formula (2).
+Since $\alpha_{2}+2\alpha_{3}\notin D$, the support $D$ may contain either $\alpha_{3}$ or $\alpha_{2}+\alpha_{3}$.
 But
     $\begin{split}\mu_{\alpha_{3}}&=\lambda_{\alpha_{3}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}+\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}}+2\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+2\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}},\\ \mu_{\alpha_{2}+\alpha_{3}}&=\lambda_{\alpha_{2}+\alpha_{3}}+\dfrac{-\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}\mu_{\alpha_{1}+\alpha_{2}}+\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}}-\\ &-2\dfrac{\mu_{\alpha_{2}+2\alpha_{3}+2\alpha_{4}}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}}}{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}.\end{split}$
-    One can see that in both cases formula (2) is satisfied.
-Second, if $\lambda_{\alpha_{2}+2\alpha_{3}}\neq 0$ formula (2) for $\alpha_{2}$ follows from the fact that $\lambda_{\alpha_{2}+2\alpha_{3}}$ can be found from $\mu_{\alpha_{2}+2\alpha_{3}}$, while $x_{\alpha_{3}}$ can be found from $\mu_{\alpha_{2}+\alpha_{3}}$.
-  • The largest root in the support is equal to $\alpha_{1}+\alpha_{2}+2\alpha_{3}$.
-Here everything essentially boils down to the supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$, $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.(All other cases can be reduced to this two supports arguing similarly to the above.)In the first case, we have the following equations on coordinates of $\mu$:
+    One can see that formula (2) is satisfied in both situations.
+Second, if $\lambda_{\alpha_{2}+2\alpha_{3}}\neq 0$, then formula (2) for $\alpha_{2}$ follows from the fact that $\lambda_{\alpha_{2}+2\alpha_{3}}$ can be obtained from $\mu_{\alpha_{2}+2\alpha_{3}}$, while $x_{\alpha_{3}}$ can be obtained from $\mu_{\alpha_{2}+\alpha_{3}}$.
+  • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}$.
+Here the analysis essentially reduces to the two supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.(All other cases can be reduced to these two by arguments similar to those above.)For the first support, we have the following equations for the coordinates of $\mu$:
     $\begin{split}\mu_{\alpha_{2}+\alpha_{3}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+2\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{2}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+2\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{4}}&=\dfrac{-\mu_{\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
-    For the second support, we have the following equations on coordinates of $\mu$:
+    For the second support, the equations for the coordinates of $\mu$ are:"
     $\begin{split}\mu_{\alpha_{2}}&=-\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+\alpha_{3}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}^{2}\mu_{\alpha_{2}+2\alpha_{3}}}{4\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}^{2}};\\ \mu_{\alpha_{4}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-2\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
-  • The largest root in the support is equal to $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$.
-Despite the fact that there are a lot of such supports, they are all amenable to simple analysis.
-Each such a support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{1}+\alpha_{2}+\alpha_{3}$.
-Therefore, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$ and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as a rational function in $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as a polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$).
-Besides, $x_{\alpha_{1}}$ can be expressed as a rational function in $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$.
-It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ enters a monomial in a coordinate of $\mu_{\alpha}$ then this monomial can be expressed as a rational function in coordinates $\mu_{\beta}$ with $\beta\succ\alpha$.
-Moreover, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$.
-Finally, those monomials that can not expressed in such a way satisfy formula (2) due to the analysis above.
-A similar analysis can also be done with supports where the largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
-Further steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ works or the equations for the coordinates can be easily explicitly computed.
+  • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$.
+Although there are many such supports, they are all amenable to a simple analysis.
+Each such support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, and $\alpha_{1}+\alpha_{2}+\alpha_{3}$.
+Consequently, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$, and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as rational function of $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$).
+Moreover, $x_{\alpha_{1}}$ can be expressed as a rational function of $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$.
+It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ appears in a monomial in a coordinate $\mu_{\alpha}$, then that monomial can be expressed as a rational function of the coordinates $\mu_{\beta}$ with $\beta\succ\alpha$.
+Furthermore, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$.
+Finally, those monomials that cannot be expressed in this way satisfy formula (2) by the analysis above.
+A similar analysis applies to supports whose largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
+The remaining steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, or the equations for the coordinates can be written out explicitly in a straightforward manner.
 Thus, Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved.
 $\square$
-Finally, let us formulate and prove the last theorem, which allows us to obtain the main result.
-Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
+Finally, we formulate and prove the last theorem, which leads to the main result.
+Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ with $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
 Proof.
 Assume the contrary.
-Let $\lambda\neq\lambda^{\prime}$ and at the same time $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$.
+Let $\lambda\neq\lambda^{\prime}$ but $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$.
 Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$.
-Then $\gamma$ must be in the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$.
-Since both of these forms belong to $S$, the root $\gamma$ satisfies condition (1).
-Let $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$.
-It follows from Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that there exist a rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
+Then $\gamma$ belongs to the support of at least one of the linear forms $\lambda$ or $\lambda^{\prime}$.
+Since both forms lie in $S$, the root $\gamma$ satisfies condition (1).
+Set $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$.
+By Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, there exist rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
 $\mu(e_{\gamma})=\lambda(e_{\gamma})+F_{D,\gamma}(\mu(e_{\beta}),\beta\succ\gamma)$
@@
 Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$.
-Hence, we may choose our functions $F_{D,\gamma}$ for various $D$ and $\gamma$ so that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$.
-Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary form from their common orbit, $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$.
+Therefore, we may choose our functions $F_{D,\gamma}$ for different $D$ and $\gamma$ in such a way that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$.
+Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary element of their common orbit, where $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$.
 Then
@@
 […]
-Thus, distinct forms in $S$ define have distinct orbits.
-Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$ then $N.\lambda\neq N.\lambda^{\prime}$.
+Thus, distinct forms in $S$ have distinct orbits.
+Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$, then $N.\lambda\neq N.\lambda^{\prime}$.
 It remains to prove that all coadjoint orbits are obtained in this way.
-Consider the orbits over the field $\mathbb{F}_{q}$.
-It is easy to calculate $|S|$.
-Let $D$ be a support from the list, $|D|=k$.
+Consider the orbits over the finite field $\mathbb{F}_{q}$.
+One can easily compute $|S|$.
+Let $D$ be a support from the list, with $|D|=k$.
 There are $(q-1)^{k}$ linear forms with this support.
-If they all lie in $S$, then this support defines $(q-1)^{k}$ forms from $S$.
-If there is one restriction on the scalars, then the support defines $(q-1)^{k-1}$ forms from $S$, since the restriction means that one coordinate is expressed through others.
+If they all lie in $S$, then this support contributes $(q-1)^{k}$ forms to $S$.
+If there is one restriction on the scalars, then the support yields $(q-1)^{k-1}$ forms in $S$, because the restriction expresses one coordinate in terms of the others.
 For example,
 $\lambda_{\alpha_{1}+\alpha_{2}}=\dfrac{\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}{\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{2}}$
-for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, one can easily calculate that
+for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, a straightforward calculation gives
 $\begin{split}|S|&=1+24(q-1)+140(q-1)^{2}+288(q-1)^{3}+256(q-1)^{4}\\ &+124(q-1)^{5}+40(q-1)^{6}+9(q-1)^{7}+(q-1)^{8}.\end{split}$
-The sum of the coefficients is exactly 883, the number of supports in the list plus three supports with restrictions.
-Since each form from $S$ defines its own orbit, the set $S$ defines $|S|$ orbits.
-But in work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is calculated using group-theoretic methods.
-And it is exactly equal to $|S|$.
-Hence, over the field $\mathbb{F}_{q}$ there are no other orbits.
-The main result now is completely proved.
-Similarly, other root systems $\Phi$ can be considered.
-Let us fix a linear lexicographic order $\succ$ on the positive roots of $\Phi^{+}$.
-Let $\lambda\in\mathfrak{n}^{*}$, $\gamma\in\Phi^{+}$.
-Consider the following matrices.
+The sum of the coefficients is exactly 883, which is the number of supports in the list plus three supports with restrictions.
+Since each form in $S$ determines its own orbit, the set $S$ accounts for $|S|$ orbits.
+However, in the work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is computed by group-theoretic methods, and it coincides exactly with $|S|$.
+Hence, over $\mathbb{F}_{q}$ there are no other orbits.
+This completes the proof of the main result.
+The same approach can be applied to other root systems $\Phi$.
+Fix a total lexicographic order $\succ$ on the positive roots $\Phi^{+}$.
+For $\lambda\in\mathfrak{n}^{*}$ and $\gamma\in\Phi^{+}$, define the matrices
 $\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
-Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: any $\gamma\in\textup{supp}(\lambda)$ satisfies condition (1). We can formulate a hypothesis.
-Conjecture 3.8. For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
+Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ satisfying the following condition: every $\gamma\in\textup{supp}(\lambda)$ fulfills condition (1). We can now state a conjecture.
+Conjecture 3.8. For every coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on that orbit.
 Matvey A.

### v5

@@
 It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called the coadjoint action.
-Such orbits play a key role in A.A.
-Kirillov’s orbit method.
+Such orbits play a key role in $A.$ Kirillov’s orbit method.
 In this work, we classify the orbits of this action in the case of $\Phi=F_{4}$ in terms of supports of canonical forms.
@@
 The support of a form $\lambda\in\mathfrak{n}^{*}$ is the set of positive roots $\alpha\in\Phi^{+}$ such that $\lambda(e_{\alpha})\neq 0$.
-Keywords: coadjoint orbit, the orbit method, root system.
-AMS subject classification: 17B08, 17B10, 17B22, 17B25, 17B30, 17B35.
-1. Introduction and the main result
+keywords:
+Coadjoint orbit, the orbit method, root system.
+[…]
+[…]
+1 Introduction and the main result
 Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra of $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space of $\mathfrak{n}$.
 The group NN is also a maximal unipotent subgroup of the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup of $G$.
-The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$.
+The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g\cdot\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$.
 According to the orbit method introduced by A.A.
@@
 Almost all coadjoint orbits that have been studied well so far are so-called orbits associated with rook placements.
-They are discussed in detail in the works [An95], [AN06], [Ko12], [Ko13], [Pa08], [IPa09], [Ig09], [Ig11], [Ig12], [IS21] and [IS23].
+They are discussed in detail in the works [1, 2, 11, 12, 13, 8, 4, 5, 7, 9, 6].
 Explicitly, the coadjoint action is given by:
-$\begin{split}(\exp(x).\lambda)(y)=\lambda(\exp(-\mathrm{ad}_{x})(y))=\lambda(y)-\lambda([x,y])+\dfrac{1}{2!}\lambda([x,[x,y]])-\ldots,\\ x,y\in\mathfrak{n},~\lambda\in\mathfrak{n}^{*}.\end{split}$
+$\begin{split}(\exp(x)\cdot\lambda)(y)=\lambda(\exp(-\mathrm{ad}_{x})(y))=\lambda(y)-\lambda([x,y])+\dfrac{1}{2!}\lambda([x,[x,y]])-\ldots,\\ x,y\in\mathfrak{n},~\lambda\in\mathfrak{n}^{*}.\end{split}$
 Let $\{e_{\alpha},\alpha\in\Phi^{+}\}$ be a basis of $\mathfrak{n}$ consisting of root vectors, and let $\{e_{\alpha}^{*},\alpha\in\Phi^{+}\}$ be the dual basis in $\mathfrak{n}^{*}$.
 The support of a linear form $\lambda\in\mathfrak{n}^{*}$ is defined as the subset of positive roots given by
-$\textup{supp}(\lambda)=\{\alpha\in\Phi^{+}\colon\lambda(e_{\alpha})\neq 0\}.$
+$\operatorname{supp}(\lambda)=\{\alpha\in\Phi^{+}\colon\lambda(e_{\alpha})\neq 0\}.$
 Our goal is to describe the coadjoint orbits for the case of $\Phi=F_{4}$.
@@
 A list of supports will be provided with the following properties.
-  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ whose supports belong to the given list (even if the supports coincide), the orbits $N.\lambda_{1}$ and $N.\lambda_{2}$ are different.
-  • For any linear form $\lambda\in\mathfrak{n}^{*}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
+  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ whose supports belong to the given list (even if the supports coincide), the orbits $N\cdot\lambda_{1}$ and $N\cdot\lambda_{2}$ are different.
+  • For any linear form $\lambda\in\mathfrak{n}^{*}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N\cdot\lambda=N\cdot\lambda^{\prime}$.
 In other words, each coadjoint orbit contains a unique linear form whose support belongs to the given list.
@@
 Let $\lambda\in\mathfrak{n}^{*}$, $\gamma\in\Phi^{+}$. Consider the following matrices.
-$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
+$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}\cdot\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
 The columns of the matrices are numbered by all positive roots, and the rows are numbered by certain positive roots.
 The matrix $A_{\lambda,\gamma}$ is the matrix $B_{\lambda,\gamma}$ with one more row.
-Further, let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: for any $\gamma\in\textup{supp}(\lambda)$,
-$\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}.$ (1)
+Further, let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: for any $\gamma\in\operatorname{supp}(\lambda)$,
+$\operatorname{rk}A_{\lambda,\gamma}=\operatorname{rk}B_{\lambda,\gamma}.$ (1)
 Finally, we can formulate the main result of this work.
-Theorem 1.1. For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
-This theorem generalizes the result of the work [IS23].
+Theorem $1.$
+For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
+This theorem generalizes the result of the work [6].
 This article discusses the so-called orbits associated with rook placements.
@@
 We say that the coadjoint orbit $\Omega_{D,\xi}$ of the linear form $f_{D,\xi}$ is associated with the rook placement $D$ and the map $\xi$. Thus, $f_{D,\xi}$ is a linear form with support $D$.
-In [IS23] it is proved that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$.
+In [6] it is proved that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$.
 Moreover, if $D$ is a non-singular rook placement, then $\lambda=f_{D,\xi}$ belongs to $S$.
@@
 Therefore, distinct linear forms with the same support $D$ lie in different coadjoint orbits.
-A result similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for the root system $\Phi=G_{2}$ and field $\mathbb{C}$.
-In the same work [IS23], the equations defining the basic subvarieties $\mathcal{O}_{D,\xi}$ are given.
+A result similar to Theorem 1.1 holds for the root system $\Phi=G_{2}$ and field $\mathbb{C}$.
+In the same work [6], the equations defining the basic subvarieties $\mathcal{O}_{D,\xi}$ are given.
 By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
 $\mathcal{O}_{D,\xi}=\sum_{\alpha\in D}\Omega_{\{\alpha\},\xi_{\alpha}},$
-where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. That work also proves AndrГ©’s stratification for $G_{2}$:
+where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. That work also proves André’s stratification for $G_{2}$:
 $\mathfrak{n}^{*}=\bigsqcup_{D,\xi}\mathcal{O}_{D,\xi},$
 where the union is taken over all non-singular rook placements $D$ and all maps $\xi\colon D\to\mathbb{C}^{\times}$.
-In this case $\Phi^{+}=\{\alpha,\beta,\alpha+\beta,2\alpha+\beta,3\alpha+\beta,3\alpha+2\beta\}$.
-Let us introduce the notation of structural constants as in the work [IS23]: $e_{\alpha},e_{\beta}]=c_{1}\cdot e_{\alpha+\beta}$, $e_{\alpha},e_{\alpha+\beta}]=c_{2}\cdot e_{2\alpha+\beta}$, $e_{\alpha},e_{2\alpha+\beta}]=c_{3}\cdot e_{3\alpha+\beta}$, $e_{3\alpha+\beta},e_{\beta}]=c_{4}\cdot e_{3\alpha+2\beta}$, $e_{\alpha+\beta},e_{2\alpha+\beta}]=c_{5}\cdot e_{3\alpha+2\beta}$.
+In this case $\Phi^{+}=\{\alpha,\beta,\alpha+\beta,2\alpha+\beta,3\alpha+\beta,3\alpha+2\beta\}$. Let us introduce the notation of structural constants as in the work [6]:
+$e_{\alpha},e_{\beta}]=c_{1}\cdot e_{\alpha+\beta},\quad[e_{\alpha},e_{\alpha+\beta}]=c_{2}\cdot e_{2\alpha+\beta},\quad[e_{\alpha},e_{2\alpha+\beta}]=c_{3}\cdot e_{3\alpha+\beta},$
+$[e_{3\alpha+\beta},e_{\beta}]=c_{4}\cdot e_{3\alpha+2\beta}\quad\text{and}\quad[e_{\alpha+\beta},e_{2\alpha+\beta}]=c_{5}\cdot e_{3\alpha+2\beta}.$
 For almost every rook placement $D$, we have $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$.
 The only exception is the case $D=\{\alpha+\beta,3\alpha+\beta\}$.
-In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon{\beta,\alpha+\beta,3\alpha+\beta}\to\mathbb{C}^{\times}$ is such that its restriction to $D$ coincides with $\xi$.
+In this case $\mathcal{O}_{D,\xi}$ is the union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon{\beta,\alpha+\beta,3\alpha+\beta}\to\mathbb{C}^{\times}$ is such that its restriction to $D$ coincides with $\xi$.
 Let us write down the equations defining the orbits for $D={\alpha+\beta,3\alpha+\beta}$.
@@
 In the case $\Phi=G_{2}$ the set $S$ consists of linear forms with supports from this list:
-$\begin{split}\varnothing,\{\alpha\},\{\beta\},\{\alpha,\beta\},\{\alpha+\beta\},\{2\alpha+\beta\},\{2\alpha+\beta,\beta\},\{3\alpha+\beta\},\\ \{3\alpha+\beta,\alpha+\beta\},\{3\alpha+\beta,\alpha+\beta,\beta\},\{3\alpha+\beta,\beta\},\{3\alpha+2\beta\},\{3\alpha+2\beta,\alpha\}.\end{split}$
-The orbits of these forms correspond precisely to all the orbits whose equations are provided in [IS23] and in the present work.
-Funding. The work is supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075–15–2022–287).
-Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during the writing of this paper.
-2. Description of the set of canonical forms
-In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and to demonstrate its use.
+$\begin{split}\emptyset,\{\alpha\},\{\beta\},\{\alpha,\beta\},\{\alpha+\beta\},\{2\alpha+\beta\},\{2\alpha+\beta,\beta\},\{3\alpha+\beta\},\\ \{3\alpha+\beta,\alpha+\beta\},\{3\alpha+\beta,\alpha+\beta,\beta\},\{3\alpha+\beta,\beta\},\{3\alpha+2\beta\},\{3\alpha+2\beta,\alpha\}.\end{split}$
+The orbits of these forms correspond precisely to all the orbits whose equations are provided in [6] and in the present work.
+2 Description of the set of canonical forms
+In this section, we will obtain an explicit description of the set $S$ to prove Theorem 1.1 and to demonstrate its use.
 At first glance, it might seem that this set does not admit a nice description.
@@
 A list of 880 supports is presented in the Appendix.
-Proposition 2.1. The set $S$ consists of all linear forms with supports from the list and forms with one of the three supports
-[…]
+Proposition 2.1.
+The set $S$ consists of all linear forms with supports from the list and forms with one of the three supports
+$\begin{split}&\{\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}\},\\ &\{\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}\},\\ &\{\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}\},\end{split}$
 and of the forms with one of the three supports listed below, provided their coordinates satisfy the equation
@@
 Table 1
+[…]
+[…]
 $\ee\backslash e^{*}$ $\alpha_{4}$ $\alpha_{3}$ $\alpha_{3}+\alpha_{4}$ $\alpha_{2}$ $\alpha_{2}+\alpha_{3}$ $\alpha_{2}+\alpha_{3}+\alpha_{4}$ $\alpha_{2}+2\alpha_{3}$ $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$
@@
 […]
+$\alpha_{3}+\alpha_{4}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{2}$ $cellcolor[HTML]{ffa500}\alpha_{2}+\alpha_{3} \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{2}+\alpha_{3}+\alpha_{4}$
 […]
+$\alpha_{2}+\alpha_{3}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{4}$ $cellcolor[HTML]{ffa500}\alpha_{3} \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{3}+\alpha_{4}$ $\alpha_{3}+2\alpha_{4}$
 […]
-[…]
-[…]
-[…]
-[…]
+$\alpha_{2}+2\alpha_{3}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{4}$ $2\alpha_{4}$
+$\alpha_{2}+2\alpha_{3}+\alpha_{4}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{4}$
 […]
@@
 First, let us reformulate condition (1).
-The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: for $x,y\in\mathfrak{n}$ and $\lambda\in\mathfrak{n}^{*}$, define $(x.\lambda)(y):=-\lambda([x,y])$.
+The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: for $x,y\in\mathfrak{n}$ and $\lambda\in\mathfrak{n}^{*}$, define $(x\cdot\lambda)(y):=-\lambda([x,y])$.
 This action is called tangent.
 Let us expand $x\in\mathfrak{n}$ in the chosen basis:
-$x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}.$
-Let $\lambda\in\mathfrak{n}^{*}$, $\beta\in\Phi^{+}$. Consider the coordinates of $x.\lambda$ in the dual basis:
-$(x.\lambda)(e_{\beta})=-\lambda([\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},e_{\beta}])=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
+$x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}.$
+Let $\lambda\in\mathfrak{n}^{*}$, $\beta\in\Phi^{+}$. Consider the coordinates of $x\cdot\lambda$ in the dual basis:
+$(x\cdot\lambda)(e_{\beta})=-\lambda([\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},e_{\beta}])=-\sum_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
 We can view these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$.
-By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
-Clearly, $\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$).
-In other words, $(x.\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
-Example 2.2.
-Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
+By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x\cdot\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x\cdot\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
+Clearly, $\operatorname{rk}A_{\lambda,\gamma}=\operatorname{rk}B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$).
+In other words, $(x\cdot\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x\cdot\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
+Example $2.$
+Let $D=\operatorname{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
 We will show that such a $\lambda$ lies in $S$.
@@
 The highlighted blue columns correspond to the roots in $D$.
-The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, which is equivalent to $\alpha+\beta\in D$.
-It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row corresponding to $\beta$ in the highlighted column.
-For example, the combination $(x.\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
+The linear combination $(x\cdot\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, which is equivalent to $\alpha+\beta\in D$.
+It can be seen from the construction of the table that $(x\cdot\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row corresponding to $\beta$ in the highlighted column.
+For example, the combination $(x\cdot\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
 Now let us check condition (1) for all roots from $D$.
-As you can see from the table, $(x.\lambda)(e_{\alpha_{2}+2\alpha_{3}})=0$.
+As you can see from the table, $(x\cdot\lambda)(e_{\alpha_{2}+2\alpha_{3}})=0$.
 Naturally, this will always be the case for the maximal root in the support.
-Note that $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$.
-Further, $(x.\lambda)(e_{\alpha_{2}})=0$, because there are no orange cells in the blue columns in the row $\alpha_{2}$; $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})=0$ for similar reasons.
-Finally, $(x.\lambda)(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$, which means that it is a scalar multiple of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$.
-It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x.\lambda)(e_{\alpha_{4}})$.
-This is simply because $x_{\alpha_{3}}$ as a variable can be expressed in terms of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient.
+Note that $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$.
+Further, $(x\cdot\lambda)(e_{\alpha_{2}})=0$, because there are no orange cells in the blue columns in the row $\alpha_{2}$; $(x\cdot\lambda)(e_{\alpha_{3}+\alpha_{4}})=0$ for similar reasons.
+Finally, $(x\cdot\lambda)(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$, which means that it is a scalar multiple of $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$.
+It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x\cdot\lambda)(e_{\alpha_{4}})$.
+This is simply because $x_{\alpha_{3}}$ as a variable can be expressed in terms of $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x\cdot\lambda)(e_{\alpha_{2}})$, $(x\cdot\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x\cdot\lambda)(e_{\alpha_{3}})$, $(x\cdot\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient.
 The Sufficient condition below is based on this idea.
 Consider the tangent action $\mathfrak{n}\curvearrowright\mathfrak{n}^{*}$.
-Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$.
+Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$.
 Starting from the support $D$ and the root $\gamma$, we will construct a set $\zeta_{D,\gamma}$ of positive roots.
@@
 Suppose that for every root $\gamma\in D$, the following condition holds: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $\alpha\in\zeta_{D,\gamma}$.
-Then $\lambda\in S$.
-Proof.
+Then $\lambda\in S$
+Proof 2.4.
 Fix a root $\gamma\in D$.
-Consider the orbit of the infinitesimal action $\mathfrak{n}.\lambda$.
-Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, where $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$.
+Consider the orbit of the infinitesimal action $\mathfrak{n}\cdot\lambda$.
+Let $\mu\in\mathfrak{n}\cdot\lambda$, $\mu=x\cdot\lambda$, where $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$.
 Here $\mu$ is given, while the values $x_{\alpha}$ are unknown.
@@
 This can be rewritten as:
-$\lambda([e_{\delta},e_{\beta}])x_{\delta}=-\mu(e_{\beta})-\sum\limits_{\alpha\neq\delta}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
+$\lambda([e_{\delta},e_{\beta}])x_{\delta}=-\mu(e_{\beta})-\sum_{\alpha\neq\delta}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
 We claim that the right-hand side is a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
@@
 Since $\lambda([e_{\delta},e_{\beta}])\neq 0$, it follows that $x_{\delta}$ can also be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
-[…]
-Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ will also be satisfied for it.
-Therefore, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none do.
+Now consider $\mu(e_{\gamma})=-\sum_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\gamma}])x_{\alpha}$.
+By the hypothesis of the proposition, if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
+Consequently, the same holds for $\mu(e_{\gamma})$, which is equivalent to condition (1) for the root $\gamma$.
+The result follows.
+Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition 2.3 will also be satisfied for it.
+Therefore, for a given support $D$, either all forms with this support satisfy the conditions of Proposition 2.3, or none do.
 Hence, it suffices to check the conditions for a single form with a given support, for example, for the form with all coordinates equal to 11 at the basis vectors $e_{\alpha}^{*}$ with $\alpha\in D$.
 Proposition 2.5.
-(Necessary condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$, $\gamma\in D$.
-Suppose that there exist positive roots $\delta_{1},\delta_{2},\dots\delta_{k},\epsilon_{1},\epsilon_{2},\dots\epsilon_{k}$ (not necessarily distinct) satisfying the conditions:
+(Necessary condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$, $\gamma\in D$.
+Suppose that there exist positive roots $\delta_{1},\delta_{2},\dots\delta_{k},\varepsilon_{1},\varepsilon_{2},\dots\varepsilon_{k}$ (not necessarily distinct) satisfying the conditions:
   • 1.
@@
   • 2.
-    $\lambda([e_{\delta_{i}},e_{\beta}])\neq 0$ and $\beta\notin\zeta_{D,\gamma}\iff\beta=\epsilon_{i-1}$ or $\beta=\epsilon_{i}$ for all $i=\overline{1,k-1}$;
+    $\lambda([e_{\delta_{i}},e_{\beta}])\neq 0$ and $\beta\notin\zeta_{D,\gamma}\iff\beta=\varepsilon_{i-1}$ or $\beta=\varepsilon_{i}$ for all $i=\overline{1,k-1}$;
   • 3.
-    $\lambda([e_{\alpha},e_{\epsilon_{i}}])\neq 0$ and $\alpha\succeq\gamma\iff\alpha=\delta_{i}$ or $\alpha=\delta_{i+1}$ for all $i=\overline{1,k}$;
+    $\lambda([e_{\alpha},e_{\varepsilon_{i}}])\neq 0$ and $\alpha\succeq\gamma\iff\alpha=\delta_{i}$ or $\alpha=\delta_{i+1}$ for all $i=\overline{1,k}$;
   • 4.
-    $\lambda([e_{\delta_{k}},e_{\epsilon_{k}}])\neq 0$ and $\epsilon_{k}\notin\zeta_{D,\gamma}$.
+    $\lambda([e_{\delta_{k}},e_{\varepsilon_{k}}])\neq 0$ and $\varepsilon_{k}\notin\zeta_{D,\gamma}$.
 Then $\lambda\notin S$.
-Proof.
-It is enough to show that $(x.\lambda)(e_{\gamma})$ cannot be expressed as a linear combination of $\{(x.\lambda)(e_{\beta})\}_{\beta\succ\gamma}$.
-In other words, consider the orbit of the tangent action $\mathfrak{n}.\lambda$, and let $\mu\in\mathfrak{n}.\lambda$.
+Proof 2.6.
+It is enough to show that $(x\cdot\lambda)(e_{\gamma})$ cannot be expressed as a linear combination of $\{(x\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$.
+In other words, consider the orbit of the tangent action $\mathfrak{n}\cdot\lambda$, and let $\mu\in\mathfrak{n}\cdot\lambda$.
 It is enough to check that if all values $\mu(e_{\beta})$ are known for $\beta\succ\gamma$, then the value of $\mu(e_{\gamma})$ can be arbitrary.
-We will construct $\mu^{\prime}\in\mathfrak{n}.\lambda$ such that $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is an arbitrary given value.
-Let $\mu^{\prime}=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$.
-It follows from the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$.
-We define all other values $x_{\alpha}$, except $x_{\epsilon_{i}}$, to be the same as for $\mu$.
-We will determine the values $x_{\epsilon_{i}}$ by induction on $i=\overline{1,k-1}$.
+We will construct $\mu^{\prime}\in\mathfrak{n}\cdot\lambda$ such that $\mu^{\prime}(e_{\beta})=\mu(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu^{\prime}(e_{\gamma})=A$, where AA is an arbitrary given value.
+Let $\mu^{\prime}=x\cdot\lambda$, $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$.
+It follows from the proof of Proposition 2.3 that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ is uniquely defined by the values $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$.
+We define all other values $x_{\alpha}$, except $x_{\varepsilon_{i}}$, to be the same as for $\mu$.
+We will determine the values $x_{\varepsilon_{i}}$ by induction on $i=\overline{1,k-1}$.
 Base case.
-Let $i=1$, $A=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$.
-Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$, then either $\alpha=\epsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$.
-Hence, in the right-hand side, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ except possibly $x_{\epsilon_{1}}$, are already defined.
-Since $\lambda([e_{\delta_{1}},e_{\epsilon_{1}}])\neq 0$, the value of $x_{\epsilon_{1}}$ is uniquely determined as well.
-After this, the only expression $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that may still depend on $x_{\epsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$.
-If $k=1$, then nothing else depends on $x_{\epsilon_{1}}$.
+Let $i=1$, $A=\sum_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{1}},e_{\alpha}])x_{\alpha}$.
+Clearly, if $\lambda([e_{\delta_{1}},e_{\alpha}])\neq 0$, then either $\alpha=\varepsilon_{1}$ or $\alpha\in\zeta_{D,\gamma}$.
+Hence, in the right-hand side, all values $x_{\alpha}$ with nonzero $\lambda([e_{\delta_{1}},e_{\alpha}])$ except possibly $x_{\varepsilon_{1}}$, are already defined.
+Since $\lambda([e_{\delta_{1}},e_{\varepsilon_{1}}])\neq 0$, the value of $x_{\varepsilon_{1}}$ is uniquely determined as well.
+After this, the only expression $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ that may still depend on $x_{\varepsilon_{1}}$ is $\mu^{\prime}(e_{\delta_{2}})$.
+If $k=1$, then nothing else depends on $x_{\varepsilon_{1}}$.
 Inductive step.
-Suppose the value $x_{\epsilon_{i-1}}$ has been found.
-Consider $\mu(e_{\delta_{i}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$.
-On the right-hand side, only $x_{\epsilon_{i-1}},x_{\epsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$, can appear with non-zero coefficients.
-The value $x_{\epsilon_{i-1}}$ was determined in the previous step.
-Thus, on the right-hand side, only $x_{\epsilon_{i}}$ may be undefined, and it definitely appears.
-From this equation, the value $x_{\epsilon_{1}}$ is uniquely determined.
-The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that may still depend on $x_{\epsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
+Suppose $x_{\varepsilon_{i-1}}$ has been found.
+Consider $\mu(e_{\delta_{i}})=\sum_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{i}},e_{\alpha}])x_{\alpha}$.
+On the right-hand side, only $x_{\varepsilon_{i-1}},x_{\varepsilon_{i}}$ and $x_{\alpha}$, for which $\alpha\in\zeta_{D,\gamma}$, can appear with non-zero coefficients.
+The value $x_{\varepsilon_{i-1}}$ was determined in the previous step.
+Thus, on the right-hand side, only $x_{\varepsilon_{i}}$ may be undefined, and it definitely appears.
+From this equation, the value $x_{\varepsilon_{1}}$ is uniquely determined.
+The only $\mu^{\prime}(e_{\beta}),\beta\succ\gamma$ that may still depend on $x_{\varepsilon_{i}}$ is $\mu^{\prime}(e_{\delta_{i+1}})$.
 Case $i=k$.
-Consider $\mu(e_{\delta_{k}})=\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$.
-The right-hand side definitely contains $x_{\epsilon_{k}}$ and possibly $x_{\epsilon_{k-1}}$.
+Consider $\mu(e_{\delta_{k}})=\sum_{\alpha\in\Phi^{+}}\lambda([e_{\delta_{k}},e_{\alpha}])x_{\alpha}$.
+The right-hand side definitely contains $x_{\varepsilon_{k}}$ and possibly $x_{\varepsilon_{k-1}}$.
 The value $x_{\varepsilon_{k-1}}$ was determined in the previous step (or does not appear if $k=1$).
 All other $x_{\alpha}$ are also defined.
-Therefore, the value $x_{\epsilon_{k}}$ is uniquely determined.
-Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depends on $x_{\epsilon_{k}}$, and its value no longer affects anything else.
-Thus, we have computed all coefficients $x_{\alpha}$ and constructed the required $\mu^{\prime}$. $\square$
-Also note that the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ do not depend on the specific coordinates of $\lambda$, but only on its support.
+Therefore, the value $x_{\varepsilon_{k}}$ is uniquely determined.
+Moreover, no other $\mu^{\prime}(e_{\beta}),\beta\succeq\gamma$ depends on $x_{\varepsilon_{k}}$, and its value no longer affects anything else.
+Thus, we have computed all coefficients $x_{\alpha}$ and constructed the required $\mu^{\prime}$.
+Also note that the conditions of Proposition 2.5 do not depend on the specific coordinates of $\lambda$, but only on its support.
 Therefore, it is sufficient to check these conditions for just one form with a given support; for example, we may assume that all its nonzero coordinates are equal to 11.
-The conditions of Propositions Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are formulated so that they can be checked algorithmically.
-Below we present algorithms for checking the necessary and sufficient conditions for a support to belong to the set SS."
+The conditions of Propositions 2.3 and 2.5 are formulated so that they can be checked algorithmically.
+Below we present algorithms for checking the necessary and sufficient conditions for a support to belong to the set $S$.
 Algorithm
@@
 The enumeration proceeds according to the following rules.
-First, the support $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$} is checked against the necessary (respectively, sufficient) condition.
+First, the support of the element $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$} is checked against the necessary (respectively, sufficient) condition.
 Next, suppose we have checked a support $D$ against the necessary (respectively, sufficient) condition.
@@
 This yields the set $\zeta_{D,\gamma}$.
-Checking the sufficient condition for the support $D$ is straightforward.
-We iterate over the row indexed by $\gamma$ and verify the condition from Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+Checking the sufficient condition for the support $D$ is straightforward. We iterate over the row indexed by $\gamma$ and verify the condition from Proposition 2.3.
 The necessary condition is checked as follows.
@@
 It is clear that linear forms with supports numbered 2, 6, 12, 14, 16, 18, 19, 20 and 21 do not satisfy condition (1); that is, such forms do not belong to $S$.
-This follows immediately from Table 1, similar to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-In the remaining cases, it is necessary to write out the coordinates of the infinitesimal action explicitly.
+This follows immediately from Table 1, similar to Example $2.$ In the remaining cases, it is necessary to write out the coordinates of the infinitesimal action explicitly.
 It turns out that the linear forms with supports numbered 25 and 33 lie in $S$ without any restrictions on their coordinates.
-Linear forms with supports numbered 30, 31, and 32 lie in $S$ if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
-Thus, we obtain 883 supports whose corresponding forms belong to $S$, three of which require an additional condition on the coordinates.
-This completes the proof of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+Linear forms with supports numbered 30, 31, and 32 lie in $S$ if and only if
+$\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+Thus, we obtain 883 supports whose corresponding forms belong to $S$, three of which require an additional condition on the coordinates. This completes the proof of Proposition 2.1.
 Example 2.7.
@@
 We will show that $\lambda\in S$.
-Let $\mu\in\mathfrak{n}.\lambda,~\mu=x.\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}^{*}$.
+Let $\mu\in\mathfrak{n}\cdot\lambda$, $\mu=x\cdot\lambda,~x=\sum\nolimits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},~\lambda=\sum\nolimits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\nolimits_{\alpha\in\Phi^{+}}\mu_{\alpha}e_{\alpha}^{*}$.
 Let us write down explicitly the coordinates of interest.
 We denote the structure constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
-Their values can be found in [KP23].
+Their values can be found in [10].
 Then
-$\begin{split}\mu_{\alpha_{1}+\alpha_{2}}&=\mu_{\alpha_{2}+2\alpha_{3}}=\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=0;\\ \mu_{\alpha_{1}}&=-\lambda([x,e_{\alpha_{1}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{1}}])=-C_{\alpha_{2},\alpha_{1}}x_{\alpha_{2}}\lambda(e_{\alpha_{1}+\alpha_{2}})=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ &=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\lambda_{\alpha_{1}+\alpha_{2}}}\mu_{\alpha_{1}};\\ \mu_{\alpha_{3}}&=-\lambda([x,e_{\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=-x_{\alpha_{4}}\lambda_{\alpha_{3}+\alpha_{4}}-2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{4}}&=-\lambda([x,e_{\alpha_{4}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])\\ &=-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\\ &=\dfrac{\lambda_{\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{2}+\alpha_{3}}-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{3}}.\end{split}$
+$\begin{split}\mu_{\alpha_{1}+\alpha_{2}}&=\mu_{\alpha_{2}+2\alpha_{3}}=\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=0;\\ \mu_{\alpha_{1}}&=-\lambda([x,e_{\alpha_{1}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{1}}])=-C_{\alpha_{2},\alpha_{1}}x_{\alpha_{2}}\lambda(e_{\alpha_{1}+\alpha_{2}})=-x_{\alpha_{2}}\lambda_{\alpha_{1}+\alpha_{2}};\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=\\ &=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\\ &=-x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ &=x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{\lambda_{\alpha_{1}+\alpha_{2}}}\mu_{\alpha_{1}};\\ \mu_{\alpha_{3}}&=-\lambda([x,e_{\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])\\ &=-C_{\alpha_{4},\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\\ &-x_{\alpha_{4}}\lambda_{\alpha_{3}+\alpha_{4}}-2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{4}}&=-\lambda([x,e_{\alpha_{4}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])\\ &=-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\\ &=x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\dfrac{\lambda_{\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{2}+\alpha_{3}}-\dfrac{\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}}{2\lambda_{\alpha_{2}+2\alpha_{3}}}\mu_{\alpha_{3}}.\end{split}$
 Let us also consider support number 25.
@@
 Hence, the coordinate $\mu_{\alpha_{2}}$ must be a linear combination of $\mu_{\alpha_{2}+2\alpha_{3}}$ and $\mu_{\alpha_{2}+\alpha_{3}}$.
-It is easy to verify that this holds if and only if $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+It is easy to verify that this holds if and only if
+$\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
 Now, let us check condition (1) for the root $\alpha_{4}$. Note that
@@
 By an argument analogous to the one above, we find that the coordinate $\mu_{\alpha_{4}}$ must be a linear combination of $\mu_{\alpha_{1}}$ and $\mu_{\alpha_{3}+\alpha_{4}}$.
-It turns out that this condition is also equivalent to $\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
-3. Classification of orbits
-In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, thereby completing the classification of coadjoint orbits for type $F_{4}$.
-For the next stage, we will need the concept of a singular root.
-Definition 3.1. Let $\alpha,\gamma\in\Phi^{+}$. The root $\alpha$ is called singular to $\gamma$ if $\gamma-\alpha\in\Phi^{+}$. The set of all roots singular to $\gamma$ is denoted by $S_{\gamma}$.
+It turns out that this condition is also equivalent to
+$\lambda_{\alpha_{1}+\alpha_{2}}\lambda^{2}_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda^{2}_{\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
+3 Classification of orbits
+In this section, we will prove Theorem 1.1, thereby completing the classification of coadjoint orbits for type $F_{4}$. For the next stage, we will need the concept of a singular root.
+Definition 3.1.
+Let $\alpha,\gamma\in\Phi^{+}$. The root $\alpha$ is called singular to $\gamma$ if $\gamma-\alpha\in\Phi^{+}$. The set of all roots singular to $\gamma$ is denoted by $S_{\gamma}$.
 It is obvious that the singular roots in $S_{\gamma}$ are divided into pairs of the form $\alpha+\beta=\gamma$.
@@
 An explicit description of the action $N_{M}\curvearrowright M^{*}$ is as follows: for $x\in M$ and $\lambda\in M^{*}$,
-$\exp(x).\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda\left(\sum\limits_{i=0}^{\infty}\dfrac{(-1)^{i}}{i!}\underbrace{[x,[x,\dots,[x,e_{\alpha}]\dots]]}_{\text{$i$ commutators}}\right)\otimes e_{\alpha}^{*}.$
-Denote by $\chi$ the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$.
-For a form $\lambda\in\mathfrak{n}^{*}$, we will consider the element $\exp(\chi).\lambda\in M^{*}$.
+$\exp(x)\cdot\lambda=\sum_{\alpha\in\Phi^{+}}\lambda\left(\sum_{i=0}^{\infty}\dfrac{(-1)^{i}}{i!}\underbrace{[x,[x,\dots,[x,e_{\alpha}]\dots]]}_{\text{$i$ commutators}}\right)\otimes e_{\alpha}^{*}.$
+Denote by $\chi$ the element $\chi=\sum_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$.
+For a form $\lambda\in\mathfrak{n}^{*}$, we will consider the element $\exp(\chi)\cdot\lambda\in M^{*}$.
 Let us also set $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$.
-Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ with $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
+Note that if $\widetilde{x}=\sum_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ with $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x})\cdot\lambda=(\exp(\chi)\cdot\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
 Now we formulate one of the most important theorems of this work. Let $y_{\beta},\beta\in\Phi^{+}$ be independent variables.
 Theorem 3.2.
-Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1).
+Let $\lambda\in S$, $D=\operatorname{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1).
 Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
 $\mu(e_{\gamma})=\lambda(e_{\gamma})+F_{D,\gamma}(\mu(e_{\beta}),\beta\succ\gamma),$ (2)
-where $\mu=\exp(\chi).\lambda\in M^{*}$.
-Example 3.3.
-Consider the same support as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
+where $\mu=\exp(\chi)\cdot\lambda\in M^{*}$.
+Example $3.$
+Consider the same support as in Example $2.$ Let
+$D=\operatorname{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
 We will show that for this $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds.
-Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}^{*}$.
+Let $\mu=\exp(\chi)\cdot\lambda,~\lambda=\sum_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}^{*}$.
 Let us write down explicitly the coordinates of interest.
-As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we denote the structural constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
-Their values can be found in [KP23].
+As in Example 2.7, we denote the structural constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$.
+Their values can be found in [10].
 We have
-$\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=0;\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}}&=\lambda(e_{\alpha_{2}})+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{2}}]])=\lambda_{\alpha_{2}}+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}}]])\\ &=\lambda_{\alpha_{2}}+\dfrac{1}{2}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}C_{\alpha_{3},\alpha_{2}}x_{\alpha_{3}}^{2}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}^{2}}{4\mu_{\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{3}+\alpha_{4}}])\\ &=\lambda_{\alpha_{4}}-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
+$\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}&=\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}&=0;\\ \mu_{\alpha_{2}+\alpha_{3}}&=-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\\ &=2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}};\\ \mu_{\alpha_{2}}&=\lambda(e_{\alpha_{2}})+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{2}}]])=\lambda_{\alpha_{2}}+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}}]])=\\ &=\lambda_{\alpha_{2}}+\dfrac{1}{2}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}C_{\alpha_{3},\alpha_{2}}x_{\alpha_{3}}^{2}\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}}-x_{\alpha_{3}}^{2}\lambda_{\alpha_{2}+2\alpha_{3}}=\\ &=\lambda_{\alpha_{2}}-\dfrac{\mu_{\alpha_{2}+2\alpha_{3}}^{2}}{4\mu_{\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{3}+\alpha_{4}}&=\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{3}+\alpha_{4}};\\ \mu_{\alpha_{4}}&=\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{3}+\alpha_{4}}])=\\ &=\lambda_{\alpha_{4}}-C_{\alpha_{3},\alpha_{4}}x_{\alpha_{3}}\lambda(e_{\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{3}}\lambda_{\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
 In fact, the validity of formula (2) in this example could have been understood without writing out the coordinates explicitly.
 It is enough to examine Table 1.
-Since $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, the variable $x_{\alpha_{3}}$ can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
-Consequently, since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ depends only on $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
+Since $\mu(e_{\alpha_{2}+\alpha_{3}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, and since $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$, the variable $x_{\alpha_{3}}$ can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
+Moreover, since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ depends only on $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$.
 Furthermore, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$.
 As $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
-Now let us proceed to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-First, we prove it in the special case of forms whose support consists of a single root.
-Any such form obviously belongs to $S$.
+Now let us proceed to the proof of Theorem 3.2. First, we prove it in the special case of forms whose support consists of a single root. Any such form obviously belongs to $S$.
 Lemma 3.4.
-Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for any $\lambda\in\mathfrak{n}^{*}$ with $|\textup{supp}(\lambda)|=1$.
-Proof.
-Denote $D=\textup{supp}(\lambda)=\{\delta\}$.
+Theorem 3.2 holds for any $\lambda\in\mathfrak{n}^{*}$ with $|\operatorname{supp}(\lambda)|=1$.
+Proof 3.5.
+Denote $D=\operatorname{supp}(\lambda)=\{\delta\}$.
 Note that in this case, condition (1) depends only on $D$, not on the specific $\lambda$, and it holds for those $\gamma$ that are not singular to $\delta$.
-Let $\mu=\exp(\chi).\lambda$.
+Let $\mu=\exp(\chi)\cdot\lambda$.
 We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}.
@@
 Note that at each step, a variable $x_{\beta_{i}}$ that was initially absent or already removed cannot reappear, because we proceed in decreasing order of the singular roots.
-As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}[t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta}]$.
-Observe that $\beta_{i}\succ\gamma$ for all $i=1,\ldots,j$.
+As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e.,
+$\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$
+for some polynomial $G\in\overline{\mathbb{F}}[t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta}]$. Observe that $\beta_{i}\succ\gamma$ for all $i=1,\ldots,j$.
 By our assumption, $\mu(e_{\gamma})$ cannot be expressed solely in terms of the subsequent coordinates.
 This implies that $G$ is not identically constant and actually depends on at least one variable $x_{\alpha}$.
-Then there exists a non-empty Zariski-open subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the specialized polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is non-constant.
+Then there exists a non-empty Zariski open subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the specialized polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is non-constant.
 Fix such a tuple $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$.
-Choose an index $\varepsilon\notin S_{\delta}$ and constants $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$, so that the one-variable polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is non-constant.
+Choose an index $\varepsilon\notin S_{\delta}$ and constants $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$, so that the one-variable polynomial
+$g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$
+is non-constant.
 The field is algebraically closed, for any chosen $c_{0}\in\overline{\mathbb{F}}$ there exists a $c_{\varepsilon}\in\overline{\mathbb{F}}$ such that $g(c_{\varepsilon})=c_{0}$.
@@
 Now, choose arbitrary values $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$.
-Substituting the values $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into formula (3), we can determine values for $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ sequentially, in such a way that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$.
-For this assignment, we obtain $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
-Therefore, there exists a non-empty Zariski-open subset $V\subseteq\overline{\mathbb{F}}^{2k+1}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ , we can find an element $\overline{\mu}\in\overline{N}.\lambda$ whose coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values.
-The orbit $\overline{N}.\lambda$ is an algebraic variety.
-Our construction yields a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, which implies $\dim\overline{N}.\lambda\geq 2k+1$.
-However, $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$.
+Substituting the values $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into formula (3), we can determine values for $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ sequentially, in such a way that $\mu(e_{\beta_{1}})=c_{1}$, $\mu(e_{\beta_{2}})=c_{2}$, $\ldots$ and $\mu(e_{\beta_{2k}})=c_{2k}$.
+For this assignment, we obtain
+$\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}.$
+Therefore, there exists a non-empty Zariski open subset $V\subseteq\overline{\mathbb{F}}^{2k+1}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$, we can find an element $\overline{\mu}\in\overline{N}\cdot\lambda$ such that $\overline{\mu}(e_{\gamma})=c_{1}$, $\overline{\mu}(e_{\beta_{1}})=c_{2}$, $\ldots$, $\overline{\mu}(e_{\beta_{2k}})c_{2k}$.
+The orbit $\overline{N}\cdot\lambda$ is an algebraic variety.
+Our construction yields a surjection from $\overline{N}\cdot\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, which implies $\dim\overline{N}\cdot\lambda\geq 2k+1$.
+However,
+$\dim\overline{N}\cdot\lambda=\dim\overline{\mathfrak{n}}\cdot\lambda=\dim\mathfrak{n}\cdot\lambda=|S_{\delta}|=2k.$
 This contradiction completes the proof.
-$\square$
-We introduce further notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
-$T_{\alpha,\beta}(x):=(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\in\overline{\mathbb{F}}[x_{\gamma},\gamma\in\Phi^{+}]$. It is clear that if $\lambda\in\mathfrak{n}^{*},\alpha\in\Phi^{+}$, then
-$\begin{split}(\exp(\chi).\lambda)(e_{\alpha})=\left(\exp(\chi).\left(\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})e_{\beta}^{*}\right)\right)(e_{\alpha})&=\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})(\exp(\chi).e_{\beta}^{*})(e_{\alpha})\\ &=\sum\limits_{\beta\in\textup{supp}(\lambda)}\lambda(e_{\beta})T_{\alpha,\beta}(x).\end{split}$
+We introduce further notation.
+Let $\alpha,\beta\in\Phi^{+}$.
+Denote $T_{\alpha,\beta}(x):=(\exp(\chi).e_{\beta}^{*})(e_{\alpha})$ in $\overline{\mathbb{F}}[x_{\gamma},\gamma\in\Phi^{+}]$.
+It is clear that if $\lambda\in\mathfrak{n}^{*},\alpha\in\Phi^{+}$, then
+$(\exp(\chi)\cdot\lambda)(e_{\alpha})$ $=\left(\exp(\chi)\cdot\left(\sum_{\beta\in\operatorname{supp}(\lambda)}\lambda(e_{\beta})e_{\beta}^{*}\right)\right)(e_{\alpha})$
+$=\sum_{\beta\in\operatorname{supp}(\lambda)}\lambda(e_{\beta})(\exp(\chi).e_{\beta}^{*})(e_{\alpha})$
+$=\sum_{\beta\in\operatorname{supp}(\lambda)}\lambda(e_{\beta})T_{\alpha,\beta}(x).$
 Now consider the following construction (cf.
-the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$).
-Let $\lambda\in S,~D=\textup{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1).
-Starting from $D$ and $\gamma$, we construct a set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition, it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
+the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition 2.3).
+Let $\lambda\in S,~D=\operatorname{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1).
+Starting from $D$ and $\gamma$, we construct a set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition, it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
 Example 3.6.
@@
 Hence, it suffices to consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots.
-For example, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
+For example,
+$T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
 Note that if $\alpha\nleq\beta$, then $T_{\alpha,\beta}(x)=0$, while $T_{\alpha,\alpha}(x)=1$.
@@
 Obviously, each trivial pair belongs to all $\zeta^{\prime}_{D,\gamma}$.
-Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-There $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
+Let us refer to Example $3.$ There $D=\operatorname{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$.
 The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ contain no nontrivial pairs.
@@
 Since
-$(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})=2x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}(\exp(\chi).\lambda)(e_{\alpha_{2}+2\alpha_{3}})$
+$(\exp(\chi)\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})=2x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=2x_{\alpha_{3}}(\exp(\chi)\cdot\lambda)(e_{\alpha_{2}+2\alpha_{3}})$
 and $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)=-x_{\alpha_{3}}^{2}$, the pair $(\alpha_{2},\alpha_{2}+2\alpha_{3})$ belongs to $\zeta^{\prime}_{D,\alpha_{2}}$.
@@
 Lemma 3.7.
-Let $\lambda\in S,D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1).
+Let $\lambda\in S,D=\operatorname{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1).
 Suppose that for every $\alpha\in D$, the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$, and formula (2) holds for all $\gamma^{\prime}\succ\gamma$ that satisfy condition (1).
 Then formula (2) also holds for $\lambda$, $D$ and $\gamma$.
-Proof. Consider
-$(\exp(\chi).\lambda)(e_{\gamma})=\sum\limits_{\alpha\in D}\lambda(e_{\alpha})T_{\gamma,\alpha}(x)=\lambda(e_{\gamma})+\sum\limits_{\alpha\in D,\alpha\succ\gamma}\lambda(e_{\alpha})T_{\gamma,\alpha}(x).$
+Proof 3.8.
+Consider
+$(\exp(\chi)\cdot\lambda)(e_{\gamma})=\sum_{\alpha\in D}\lambda(e_{\alpha})T_{\gamma,\alpha}(x)=\lambda(e_{\gamma})+\sum_{\alpha\in D,\alpha\succ\gamma}\lambda(e_{\alpha})T_{\gamma,\alpha}(x).$
 Let $\alpha\in D$ with $\alpha\succ\gamma$.
-Since $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$.
+Since $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$.
 By the definition of $S$, the root $\alpha$ satisfies condition (1).
-Therefore, by the inductive hypothesis, $\lambda(e\alpha)$ can also be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
-$\lambda(e_{\alpha})=(\exp(\chi).\lambda)(e_{\alpha})-F_{D,\alpha}((\exp(\chi).\lambda)(e_{\beta}),\beta\succ\alpha).$
-Consequently, the difference $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ is a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, which means condition (2) is satisfied.
-$\square$
-We are now ready to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-Proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-If $D=\{\alpha\}$, then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-Indeed, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$.
+Therefore, by the inductive hypothesis, $\lambda(e\alpha)$ can also be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
+$\lambda(e_{\alpha})=(\exp(\chi)\cdot\lambda)(e_{\alpha})-F_{D,\alpha}((\exp(\chi)\cdot\lambda)(e_{\beta}),\beta\succ\alpha).$
+Hence, the difference $(\exp(\chi)\cdot\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ is a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, which means condition (2) is satisfied.
+We are now ready to prove Theorem 3.2.
+Proof of Theorem 3.2.
+If $D=\{\alpha\}$, then Theorem 3.2 holds by Lemma 3.4.
+Indeed, in this case $(\exp(\chi)\cdot\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$.
 Moreover, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$.
-Finally, for $\gamma\prec\alpha$ satisfying condition (1), the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
+Finally, for $\gamma\prec\alpha$ satisfying condition (1), the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
 Now we will prove formula (2) step by step for all $\lambda\in S$ by considering the supports from the list one after another.
-(Note that most supports can be treated similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$; see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
+(Note that most supports can be treated similarly to Example 3.3; see below.) Recall that
+$\lambda=\sum_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi)\cdot\lambda\in M^{*}$.
   • Suppose the largest root (with respect to the lexicographic order) in the support is smaller than $\alpha_{2}+\alpha_{3}+\alpha_{4}$.
-For all such supports and for every $\gamma\in\Phi^{+}$ satisfying condition (1), Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, so the proof is complete.
+For all such supports and for every $\gamma\in\Phi^{+}$ satisfying condition (1), Lemma 3.7 applies, so the proof is complete.
 (Indeed, for all relevant $\gamma$, the polynomials $T_{\gamma,\alpha}(x)$ are, up to a non-zero scalar factor, equal to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0, or 11.)
@@
 There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$.
-Since Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ covers the case of one-element supports, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$).
-The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, so nothing remains to prove in this case.
+Since Lemma 3.4 covers the case of one-element supports, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$).
+The conditions of Lemma 3.7 are also satisfied, so nothing remains to prove in this case.
   • Next, suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}$.
-Here the first nontrivial case appears, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be proved for all supports of this type by a straightforward application of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
+Here the first nontrivial case appears, i.e., Theorem 3.2 cannot be proved for all supports of this type by a straightforward application of Lemma 3.7.
 Such a support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$.
@@
 However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$.
-Consequently, a situation may occur where $\alpha_{4}$ satisfies condition (1) but Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be applied directly.
+Consequently, a situation may occur where $\alpha_{4}$ satisfies condition (1) but Lemma 3.7 cannot be applied directly.
 Therefore, in this case we must compute the coordinates of $\mu$ explicitly:
-    $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}=&~\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=&~\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}}=&-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-\\ &-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{3}+\alpha_{4}}=&-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&-x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{3}}=&-\lambda([x,e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{3}}]])=-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{2}}e_{\alpha_{2}},[x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{4}}e_{\alpha_{4}},e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x_{\alpha_{4}}e_{\alpha_{4}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}}])\\ =&-C_{\alpha_{2}+\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\dfrac{1}{2}C_{\alpha_{2},\alpha_{3}}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\\ &+\dfrac{1}{2}(C_{\alpha_{4},\alpha_{3}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}+C_{\alpha_{2},\alpha_{3}}C_{\alpha_{4},\alpha_{2}+\alpha_{3}})x_{\alpha_{2}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&~2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}+x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}-x_{\alpha_{2}}x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{4}}=&~\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{4}}])+\frac{1}{2}\lambda([x,[x,e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})+\\ &+\dfrac{1}{2}C_{\alpha_{3},\alpha_{4}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\\ &+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\mu_{\alpha_{3}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
+    $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}=&~\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=&~\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}}=&-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=\\ =&-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-\\ &-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{3}+\alpha_{4}}=&-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&-x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{3}}=&-\lambda([x,e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{3}}]])=-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{2}}e_{\alpha_{2}},[x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{4}}e_{\alpha_{4}},e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x_{\alpha_{4}}e_{\alpha_{4}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}}])\\ =&-C_{\alpha_{2}+\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\dfrac{1}{2}C_{\alpha_{2},\alpha_{3}}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\\ &+\dfrac{1}{2}(C_{\alpha_{4},\alpha_{3}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}+C_{\alpha_{2},\alpha_{3}}C_{\alpha_{4},\alpha_{2}+\alpha_{3}})x_{\alpha_{2}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&~2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}+x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}-x_{\alpha_{2}}x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{4}}=&~\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{4}}])+\frac{1}{2}\lambda([x,[x,e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})+\\ &+\dfrac{1}{2}C_{\alpha_{3},\alpha_{4}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\\ &+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\mu_{\alpha_{3}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
     This shows that formula (2) holds for the root $\alpha_{4}$.
-  • Now suppose the largest root in the support equals either $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$.
-In these cases, the statement also follows almost trivially from Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
-The only nontrivial situation occurs for the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$.
-At first glance, it might seem that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$.
+  • Now suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$.
+In these cases, the statement also follows almost trivially from Lemma  3.4 and Lemma 3.7.
+The only nontrivial situation occurs for the support
+    $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$.
+    At first glance, it might seem that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$.
 However, this is not the case.
-From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$.
-Consequently, $x_{\alpha_{3}}$ can be recovered from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$.
+From Lemma 3.4 we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ can be expressed as a rational function of
+    $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}.$
+    Consequently, $x_{\alpha_{3}}$ can be recovered from $(\exp(\chi)\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$.
 Since $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ depends only on $x_{\alpha_{3}}$.
   • If the largest root in the support equals $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the nontrivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$.
-This case follows immediately from the formulas in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ together with the analysis of the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
+This case follows immediately from the formulas in Example 2.7 together with the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
   • If the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}$, then all nontrivial cases reduce to $D=\{\alpha_{1}+\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4},\alpha_{2}\}$.
@@
     We conclude that for the roots $\alpha_{2}$ and $\alpha_{3}+\alpha_{4}$, formula (2) holds.
-  • Next, if the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those containing the roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$.
+  • Next, if the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those containing the roots
+    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\quad\alpha_{2}+2\alpha_{3}+2\alpha_{4}\quad\text{and}\quad\alpha_{2}+2\alpha_{3}+\alpha_{4}.$
     One can easily show that
@@
   • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}$.
-Here the analysis essentially reduces to the two supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.(All other cases can be reduced to these two by arguments similar to those above.)For the first support, we have the following equations for the coordinates of $\mu$:
+Here the analysis essentially reduces to the two supports $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{1}+\alpha_{2}+2\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}\}$.
+(All other cases can be reduced to these two by arguments similar to those above.) For the first support, we have the following equations for the coordinates of $\mu$:
     $\begin{split}\mu_{\alpha_{2}+\alpha_{3}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+2\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{2}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{2}+2\alpha_{3}}}{\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}};\\ \mu_{\alpha_{4}}&=\dfrac{-\mu_{\alpha_{3}}\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{1}+\alpha_{2}}\mu_{\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
-    For the second support, the equations for the coordinates of $\mu$ are:"
+    For the second support, the equations for the coordinates of $\mu$ are:
     $\begin{split}\mu_{\alpha_{2}}&=-\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{2}+\alpha_{3}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}+\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}^{2}\mu_{\alpha_{2}+2\alpha_{3}}}{4\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}^{2}};\\ \mu_{\alpha_{4}}&=\dfrac{\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-2\mu_{\alpha_{1}}\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}}{2\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}}.\end{split}$
@@
 Finally, those monomials that cannot be expressed in this way satisfy formula (2) by the analysis above.
-A similar analysis applies to supports whose largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
-The remaining steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, or the equations for the coordinates can be written out explicitly in a straightforward manner.
-Thus, Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved.
+A similar analysis applies to the cases where the largest root of the support is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
+The remaining steps can be analyzed similarly: either Lemma 3.7 applies, or the equations for the coordinates can be written out explicitly in a straightforward manner.
+Thus, Theorem 3.2 is proved.
 $\square$
 Finally, we formulate and prove the last theorem, which leads to the main result.
-Theorem 3.7. Let $\lambda,\lambda^{\prime}\in S$ with $\lambda\neq\lambda^{\prime}$. Then $\overline{N}.\lambda\neq\overline{N}.\lambda^{\prime}$.
-Proof.
+Theorem 3.9.
+Let $\lambda,\lambda^{\prime}\in S$ with $\lambda\neq\lambda^{\prime}$. Then $\overline{N}\cdot\lambda\neq\overline{N}\cdot\lambda^{\prime}$.
+Proof 3.10.
 Assume the contrary.
-Let $\lambda\neq\lambda^{\prime}$ but $\overline{N}.\lambda=\overline{N}.\lambda^{\prime}$.
+Let $\lambda\neq\lambda^{\prime}$ but $\overline{N}\cdot\lambda=\overline{N}\cdot\lambda^{\prime}$.
 Let $\gamma$ be the largest positive root such that $\lambda(e_{\gamma})\neq\lambda^{\prime}(e_{\gamma})$.
@@
 Since both forms lie in $S$, the root $\gamma$ satisfies condition (1).
-Set $\mu=\exp(\chi).\lambda,~\mu^{\prime}=\exp(\chi).\lambda^{\prime}$.
-By Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, there exist rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
+Set $\mu=\exp(\chi)\cdot\lambda,~\mu^{\prime}=\exp(\chi)\cdot\lambda^{\prime}$.
+By Theorem 3.2, there exist rational functions $F_{D,\gamma}$ and $F_{D^{\prime},\gamma}$ such that
 $\mu(e_{\gamma})=\lambda(e_{\gamma})+F_{D,\gamma}(\mu(e_{\beta}),\beta\succ\gamma)$
@@
 $\mu^{\prime}(e_{\gamma})=\lambda^{\prime}(e_{\gamma})+F_{D^{\prime},\gamma}(\mu^{\prime}(e_{\beta}),\beta\succ\gamma).$
-Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ and $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$.
+Note that if $\lambda(e_{\beta})=\lambda^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$, then $\mu(e_{\gamma})-\lambda(e_{\gamma})=\mu^{\prime}(e_{\gamma})-\lambda^{\prime}(e_{\gamma})$ and $\mu(e_{\beta})=\mu^{\prime}(e_{\beta})$ for all $\beta\succ\gamma$ .
 Therefore, we may choose our functions $F_{D,\gamma}$ for different $D$ and $\gamma$ in such a way that $F_{D,\gamma}=F_{D^{\prime},\gamma}$ whenever $\{\beta\in D\mid\beta\succ\gamma\}=\{\beta\in D^{\prime}\mid\beta\succ\gamma\}$.
-Let $\widetilde{\mu}=\exp(\widetilde{x}).\lambda$ be an arbitrary element of their common orbit, where $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$.
+Let $\widetilde{\mu}=\exp(\widetilde{x})\cdot\lambda$ be an arbitrary element of their common orbit, where $\widetilde{x}=\sum_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$.
 Then
@@
 $\widetilde{\mu}(e_{\gamma})=\lambda^{\prime}(e_{\gamma})+F_{D,\gamma}(\widetilde{\mu}(e_{\beta}),\beta\succ\gamma).$
-[…]
+Note that $F_{D,\gamma}(\widetilde{\mu}(e_{\beta}),\beta\succ\gamma)$ does not become infinite because $\widetilde{\mu}(e_{\gamma})$, $\lambda(e_{\gamma})$, and $\lambda^{\prime}(e_{\gamma})$ are well-defined numbers.
+Hence, $\lambda(e_{\gamma})=\lambda^{\prime}(e_{\gamma})$.
+This is a contradiction.
 Thus, distinct forms in $S$ have distinct orbits.
-Note that $N.\lambda\subseteq\overline{N}.\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$, then $N.\lambda\neq N.\lambda^{\prime}$.
+Note that $N\cdot\lambda\subseteq\overline{N}\cdot\lambda$, so if $\lambda,\lambda^{\prime}\in S$ and $\lambda\neq\lambda^{\prime}$, then $N\cdot\lambda\neq N\cdot\lambda^{\prime}$.
 It remains to prove that all coadjoint orbits are obtained in this way.
@@
 $\lambda_{\alpha_{1}+\alpha_{2}}=\dfrac{\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}^{2}}{\lambda_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{2}}$
-for the supports 30, 31, 32, see page Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Thus, a straightforward calculation gives
+for the supports 30, 31, 32, see page 2. Thus, a straightforward calculation gives
 $\begin{split}|S|&=1+24(q-1)+140(q-1)^{2}+288(q-1)^{3}+256(q-1)^{4}\\ &+124(q-1)^{5}+40(q-1)^{6}+9(q-1)^{7}+(q-1)^{8}.\end{split}$
@@
 Since each form in $S$ determines its own orbit, the set $S$ accounts for $|S|$ orbits.
-However, in the work [GMR16], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is computed by group-theoretic methods, and it coincides exactly with $|S|$.
+However, in the work [3], the number of coadjoint orbits for $F_{4}$ over a finite field $\mathbb{F}_{q}$ is computed by group-theoretic methods, and it coincides exactly with $|S|$.
 Hence, over $\mathbb{F}_{q}$ there are no other orbits.
@@
 For $\lambda\in\mathfrak{n}^{*}$ and $\gamma\in\Phi^{+}$, define the matrices
-$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
-Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ satisfying the following condition: every $\gamma\in\textup{supp}(\lambda)$ fulfills condition (1). We can now state a conjecture.
-Conjecture 3.8. For every coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on that orbit.
-Matvey A.
-Surkov: Samara National Research University, Ak.
-Pavlova 1, Samara, 443011, Russia; Saint Petersburg University, 7/9 Universitetskaya nab., St.
-Petersburg, 199034, Russia E-mail adress: surkovmatveya@yandex.ru
+$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}\cdot\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
+Let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ satisfying the following condition: every $\gamma\in\operatorname{supp}(\lambda)$ fulfills condition (1). We can now state a conjecture.
+Conjecture 3.11.
+For every coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on that orbit.
+Acknowledgments
+This work was supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no.
+075–15–2022–287).
+I thank M.
+V.
+Ignatev and A.
+N.
+Panov for helpful comments and support during the writing of this paper.
 Appendix
@@
   875. 875.
-    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},~\alpha_{2},~\alpha_{3},~\alpha_{4}$
+    $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4},\alpha_{2},\alpha_{3},\alpha_{4}$
   876. 876.
@@
         a[2][1]=3;
-$Ψ$
         m[23][10]=-1;
@@
         m[2][1]=1;
-$Ψ$
         b = new String[25];
-        b[1]="a_4"; b[2]="a_3"; b[3]="a_3+a_4"; b[4]="a_2"; b[5]="a_2+a_3"; b[6]="a_2+a_3+a_4"; b[7]="a_2+2a_3";
-        b[8]="a_2+2a_3+a_4"; b[9]="a_2+2a_3+2a_4"; b[10]="a_1"; b[11]="a_1+a_2"; b[12]="a_1+a_2+a_3";
-        b[13]="a_1+a_2+a_3+a_4"; b[14]="a_1+a_2+2a_3"; b[15]="a_1+a_2+2a_3+a_4"; b[16]="a_1+a_2+2a_3+2a_4";
-        b[17]="a_1+2a_2+2a_3"; b[18]="a_1+2a_2+2a_3+a_4"; b[19]="a_1+2a_2+2a_3+2a_4"; b[20]="a_1+2a_2+3a_3+a_4";
-        b[21]="a_1+2a_2+3a_3+2a_4"; b[22]="a_1+2a_2+4a_3+2a_4"; b[23]="a_1+3a_2+4a_3+2a_4"; b[24]="2a_1+3a_2+4a_3+2a_4";
+        b[1]="a_4"; b[2]="a_3"; b[3]="a_3+a_4"; b[4]="a_2"; b[5]="a_2+a_3";
+        b[6]="a_2+a_3+a_4"; b[7]="a_2+2a_3"; b[8]="a_2+2a_3+a_4"; b[9]="a_2+2a_3+2a_4";
+        b[10]="a_1"; b[11]="a_1+a_2"; b[12]="a_1+a_2+a_3"; b[13]="a_1+a_2+a_3+a_4";
+        b[14]="a_1+a_2+2a_3"; b[15]="a_1+a_2+2a_3+a_4"; b[16]="a_1+a_2+2a_3+2a_4";
+        b[17]="a_1+2a_2+2a_3"; b[18]="a_1+2a_2+2a_3+a_4"; b[19]="a_1+2a_2+2a_3+2a_4";
+        b[20]="a_1+2a_2+3a_3+a_4"; b[21]="a_1+2a_2+3a_3+2a_4"; b[22]="a_1+2a_2+4a_3+2a_4";
+        b[23]="a_1+3a_2+4a_3+2a_4"; b[24]="2a_1+3a_2+4a_3+2a_4";
         c = new double[25][25];
@@
         in.close();
-$Ψ$
         while(true) {
@@
                 for(int j=1; j<=24; j++) {
-                if(roots.contains(a[i][j])||roots.contains(a[j][i])) { c[i][j]=(m[i][j]-m[j][i])*lambda[a[i][j]+a[j][i]];}
+                if(roots.contains(a[i][j])||roots.contains(a[j][i]))
+                { c[i][j]=(m[i][j]-m[j][i])*lambda[a[i][j]+a[j][i]];}
                 else { c[i][j]=0; }}}
@@
         System.out.print(n); }
-    int x; int[][]a; int[][]m; String[]b; double[][]c; double[]lambda; ArrayList<Integer> roots; ArrayList<Integer> depend;
-    boolean q; boolean p; int dep; int last; boolean rank; int n; int count; int rowold; int colold; int rownew; int colnew;
+    int x; int[][]a; int[][]m; String[]b; double[][]c; double[]lambda;
+    ArrayList<Integer> roots; ArrayList<Integer> depend;
+    boolean q; boolean p; int dep; int last; boolean rank; int n; int count; int rowold; int colold;
+    int rownew; int colnew;